A Complete Bibliography of Publications in the Bulletin of Mathematical Biophysics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.01

Title word cross-reference

(M, R) [804, 1102]. 3 [1083]. ^{+} [900, 572]. ^{−} [572]. ^{o} [678]. _{2} [556, 559, 844]. _{2}^{+} [572]. α [682]. (M, ℜ) [842]. N [435, 531, 65, 637, 1009, 972]. Q [450]. T [452].

-body [1083, 972]. -perceptron [682]. -person [435]. -pool [1009]. -systems [804, 842, 1102].

1964 [719]. 1965 [789].

29th [790].

êtres [1040].

Abhängigkeit [893]. ability [441]. abnormal [993]. absorption [556, 559, 36, 711]. Abstract [979, 805, 584, 598, 627, 852, 690, 703, 803, 823, 603, 644, 661]. abstraction [145]. abstracts [458]. Academic [795, 764]. accessible [861]. accommodation [126, 432]. accommodative [68]. accompanying [844]. account [741, 660]. Accounting [381]. accumulation [330]. achieved [433]. acid [679, 814, 480, 828]. acids [1032, 968, 1001]. acoustic [879]. across [655, 747, 1013, 1080, 1147, 1033, 549, 1118]. action [666, 672, 1135, 153, 273]. actions [189]. activated [1080]. activation [1029]. Active [1143, 493, 229, 357, 655, 768, 537, 442, 469, 576, 36, 1018, 951, 549]. activities [779, 939, 1051, 1087]. activity [48, 54, 59, 65, 343, 1009, 103, 1121, 657, 486]. actomyosin [311, 349]. acts [293]. acuity [462]. adaptation [1080, 632, 652, 765]. added [735]. Addendum [192]. Addition [267]. adequate [570, 646, 670]. adjacent [261]. adjoining [412, 440]. adjustment [320, 432]. Adsorption [707, 717, 81, 889, 683, 693, 728, 833]. aerosol [515]. aesthetic [67]. aesthetics [79, 88, 121, 127, 133]. affected [859]. affective [105]. affinities [1037]. after [974, 82]. age [959, 1017, 346, 947]. age-dependent [1017, 346]. aggregates [284, 23, 33, 28, 39]. aggregation [64]. Aggressiveness [1081]. aging [1010, 592]. aid [532]. air [237, 244, 1152]. air-borne [237, 244]. airborne [643]. aléatoire [1045]. algebra [814, 266]. algebraic [803]. alkaline [245]. allometric [732, 406]. allometry [1063]. along [742]. alpha [761]. alterations [575, 644]. alternate [74, 907]. alternating [517, 223, 415]. alternation [428]. altruistic [571]. alveolar [1099]. ameboid [111, 182]. America [790]. American [579]. amines [677]. amoeboid [5, 99]. amphetamine [907]. analogous [747, 950]. analyseurs [675]. Analysis [455, 731, 767, 399, 531, 781, 815, 861, 1099, 686, 934, 883, 955, 998, 704, 1130, 954, 513, 669, 741, 278, 362, 904, 1008, 565, 587, 605, 183, 971, 84, 69, 348, 838, 240, 277, 678, 853, 865, 199, 878, 593, 606, 1070, 1086, 858]. Analytical [875, 1051, 1146, 1025]. analyzed [274]. analyzer [675]. and/or [766]. aneurysm [822]. Angles [856]. animal [246, 269, 327, 227, 232, 1159, 222]. animals [181, 95, 1027, 136]. anisotropic [886]. anisotropy [1046]. Announcement [430, 482, 956, 1148, 1161, 75, 749]. Announcements [600]. anodal [98]. anti [857]. anti-multiplier [857]. antidiuresis [841]. Anxieties [148]. aorta [715, 741, 751, 1043]. aortic [669, 741, 731, 982, 530, 592, 1008, 878]. aperiodic [725]. apparatus [364]. apparent [357, 768]. appears [863]. Application [1101, 1015, 61, 391, 418, 39, 660, 513, 844, 655, 1080, 1005, 57, 221, 393, 633, 1050, 819, 762, 443, 661, 1073]. Applications [961, 1017, 266, 612, 818, 1117, 149, 166, 607, 639, 642, 123]. applied [698, 766, 1012, 223, 662, 300]. appreciable [351]. approach [798, 1021, 605, 965, 1135, 191, 179, 434, 934, 160, 227, 232, 22, 74, 92, 93, 137, 206, 226, 257, 334, 454, 488, 520, 505, 696, 913, 973, 986, 546, 188, 1025]. approaches [496]. Approximate [331, 313, 325, 335, 251, 204, 336]. Approximation [624, 316, 380, 500, 1069]. APS [852]. arational [953]. area [463]. arrangement [134]. arrangements [302]. arterial [530, 612, 886, 887, 896, 917, 932, 1099, 884, 965, 305, 324, 379, 691, 944]. artificial [82]. arts [230]. aspects [1109, 542, 238, 239, 551, 713, 276, 196, 420, 32, 345, 519, 740, 800, 595, 803]. assemblies [621, 636, 563]. associated [464]. associations [514]. assuming [522]. assumption [995, 862, 1036, 358]. Asymmetric [28, 57]. Asynchrony [164]. athletes [893]. atropine [907]. attacks [1122]. audible [259, 446]. audiogyral [280]. auditory [178, 67]. automata [809]. automation [525, 619]. Automobile [626, 524, 519, 548, 547, 561, 645, 712, 684, 848, 849, 926, 1038, 1044]. autopolyploids [216, 228]. auxin [229]. available [1024]. average [1067]. avoidance [687, 686, 689, 960]. Awards [287, 310, 388]. axial [86, 985]. axiomatic [946]. axon [307]. axons [975].

back [872]. bacteria [341]. bacterial [441, 303]. balance [697]. balanced [906, 938]. ballistocardiography [472]. balloon [1023]. bar [907, 1027]. based [655, 243]. Baserga [723]. bases [423]. Basic [361, 1108, 843, 844, 862, 886]. basis [230, 657, 706]. be [168]. becomes [863]. beds [1155]. bees [648]. beetle [396, 404]. before [863]. behavior [384, 425, 893, 445, 621, 636, 390, 993, 342, 1146, 1054, 277, 293, 219, 224, 264, 317, 444, 460, 533, 596, 601, 759, 772, 831, 913, 952, 953, 988, 1026, 659, 1136, 386, 774, 1055, 944]. Behavioral [805]. beings [1040]. Beitrag [479]. beliefs [272, 255]. Bernstein [30, 41]. between [89, 552, 24, 161, 404, 407, 1078, 1125, 1045, 938, 176, 52, 132, 298, 290, 583, 817, 799, 894, 953, 350, 656, 43]. bias [344, 258, 358, 359, 368, 490]. biased [473]. bibliographical [424]. bidirectional [807]. bilateral [251]. Bimolecular [891, 892, 836]. binding [746]. binocular [183]. bio [1031]. bio-sociology [1031]. biochemical [924, 1021, 403, 1030, 1126, 1160]. Biochemistry [116]. biographical [718]. Biological [558, 569, 1010, 801, 732, 422, 570, 259, 376, 392, 447, 448, 446, 450, 979, 655, 1080, 1147, 805, 961, 796, 1033, 214, 217, 238, 239, 461, 514, 377, 970, 1146, 1054, 819, 818, 62, 112, 22, 205, 224, 402, 601, 868, 845, 972, 973, 1018, 987, 494, 507, 575, 662, 644, 690, 703, 716, 803, 634, 529, 528, 539, 914, 642, 458, 2, 58, 78, 492]. biologies [603]. biologist [1111]. Biology [1089, 324, 1076, 482, 580, 590, 599, 609, 620, 664, 483, 730, 381, 823, 268, 305, 564, 572, 714, 94, 92, 93, 219, 254, 298, 373, 382, 387, 394, 408, 397, 426, 421, 436, 431, 467, 468, 481, 496, 504, 520, 505, 548, 534, 554, 584, 598, 627, 645, 667, 712, 684, 816, 848, 849, 895, 952, 926, 942, 988, 986, 1069, 1038, 1044, 864, 607, 635, 661, 593, 606, 423, 452, 660]. biomass [959]. Biomathematics [640]. biomechanics [594]. biomedical [676, 699]. biometry [580, 609, 650]. Bionics [802, 1020]. Biophysical [665, 149, 166]. Biophysics [1114, 141, 111, 53, 45, 333, 372, 10, 15, 1, 29, 44, 49, 101, 304, 428, 526, 5, 11, 17, 22, 23, 33, 40, 51, 74, 79, 67, 88, 100, 121, 127, 133, 145, 144, 131, 132, 147, 154, 148, 168, 167, 226, 524, 547, 561, 184, 150, 872]. biosociology [272, 444, 255]. biosynthesis [898]. biotopological [798]. biotopology [898, 497]. bipolar [652]. birth [483, 819]. bisexual [1117]. Bistable [840]. Blair [242]. block [1079]. Blood [343, 324, 722, 556, 559, 612, 1008, 886, 887, 896, 917, 932, 202, 965, 328, 305, 118, 487, 434, 881, 124, 204, 883, 131, 132, 177, 677, 885, 557, 122, 123, 140, 197, 1086, 1025]. blood-brain-cerebrospinal [883]. blood-flow [557]. blood-glucose [722]. blood-tissue [434, 124, 204, 122, 123, 140]. bodies [4, 20]. body [1083, 744, 551, 972, 1118, 119]. bolus [1025]. bone [1113]. Book [710, 709, 63]. borne [237, 244]. boundary [707, 717, 405, 439]. bounded [1125]. boundedness [647]. brain [783, 883, 1100, 1027, 885]. brain-like [783]. brain-operated [1027]. branched [372]. branches [156, 159, 856, 125, 128]. branching [977, 856, 1149, 343, 1107, 793, 808, 1017, 1117]. breakdown [450]. breath [1145]. bromide [61]. bronchial [856]. buffer [828]. bulk [919, 976]. Bulletin [872]. bumping [852]. burns [610].

C [1041]. calcium [1154]. Calculation [515, 1029, 1061, 612, 378]. calculus [104, 157, 175, 103]. called [9]. Can [193, 431, 307]. canal [858]. cancer [191, 220, 322, 137, 206]. canonical [628]. capacities [746]. capacity [289, 323]. capillaries [89, 1025]. capillary [320, 328, 366, 299, 354]. capsomeres [757]. carbon [278]. carboxylase [766]. cardiac [295, 360, 1143, 1157]. cardio [1034]. cardio-pulmonary [1034]. cardiovascular [726, 333, 880, 646, 670, 498, 611, 758]. caries [828, 829, 830]. Carlo [1062]. Carolina [640]. carotid [758]. carrier [330, 469]. carriers [1029]. case [768, 49, 251, 360, 337, 837, 212, 548, 972, 755, 336]. Casson [980, 1050]. cat [1023, 858]. Catalytic [892]. catatonia [363]. cause [193, 210]. cavities [23]. Cell [318, 876, 1116, 747, 889, 621, 636, 905, 974, 826, 605, 86, 1022, 1, 84, 69, 96, 827, 1078, 5, 3, 7, 18, 22, 27, 47, 100, 193, 210, 301, 150, 706, 113, 1071, 1067, 42, 563, 8, 19, 864]. cell-cycle [1116]. cell-growth [826]. celled [778, 997]. cells [259, 376, 392, 447, 448, 446, 450, 961, 71, 837, 838, 480, 788, 1068, 652, 50, 73, 77, 4, 20]. Cellular [922, 923, 86, 900, 243, 45, 108, 23, 33, 28, 39, 90, 784, 1067]. center [903, 1153]. centers [675]. central [141, 10, 9, 147, 154, 1134, 188]. centres [675]. cerebral [487]. cerebrospinal [883]. certain [581, 276, 880, 794, 597, 1026]. Chain [149, 166, 1137, 265, 1156]. chains [483, 1030, 797]. chambered [611]. change [912, 596]. changes [268, 428, 953, 575]. changing [320]. channels [263]. characteristics [564, 572, 246, 52, 485]. characterization [1095, 1070]. charge [1029, 1147, 899, 624]. charged [683, 693, 707, 717, 728, 833]. charges [50]. chemical [489, 523, 550, 1083, 843, 844, 414, 1135, 295, 360, 405, 412, 433, 439, 440, 941, 595, 642]. chemiperistaltic [1013]. chemotherapy [542, 551]. chez [1040]. Chicago [287, 310, 388]. chloride [36, 61, 119]. chloride-bromide-deuterium [61]. chloroplast [365]. choice [425]. Chorioretinal [1055, 1077]. chromosome [838]. chromosome-number [838]. chromosomes [42]. chronic [1110]. Church [635]. Cilia [1140, 875]. cinetica [629]. circadian [908]. circles [628]. circuit [54, 59, 36, 61, 80, 1122]. circuits [65]. Circulation [324, 305, 407, 881, 131, 132, 691, 738, 777, 124]. circulatory [343, 372, 725]. circumferential [873, 847]. clamp [1119]. class [793, 808, 1026]. clicks [256]. climatic [257]. Clinical [1128, 612]. clocks [908]. closed [1137, 782, 834]. closely [262]. cloud [422]. CO [559]. coagulation [202]. codes [679]. coding [753, 503, 518, 543, 589]. coefficient [235, 261]. coefficients [380, 393, 500, 282, 355]. coenzyme [229]. coenzyme-like [229]. cold [247]. collateral [738]. collection [1062]. College [640]. collide [378, 288]. collision [385]. colloidal [81, 64, 13, 395]. colonial [70]. colonies [484]. colony [130]. Color [692, 304, 428, 526, 678, 602, 526]. color-matching [678]. colors [428]. combinations [74, 67]. combinatorial [394, 476]. combined [858]. Comments [1004, 1041, 313, 325, 335, 955, 954, 518, 543, 720, 842]. Committee [457]. common [65, 522]. Communication [291, 893, 383, 263, 266, 356, 222]. Comparaison [820]. Comparison [1023, 689, 988, 475, 1139, 496, 14, 820]. comparisons [934]. compartment [974, 531, 815, 782, 834, 350, 1120]. Compartmental [724, 902]. compartmentalization [721]. compartments [1137, 680, 773, 1120]. compatibility [471]. competition [659, 404]. competitive [727]. competitors [935]. complete [962, 59, 946]. completion [383]. complex [905, 746, 549]. complexed [900]. complexes [891, 836]. complexity [928, 931, 945, 957]. component [1051]. components [614]. composite [1130]. Compound [437]. compressibility [376, 447]. computation [592, 292]. Computed [879]. Computer [991, 929, 676, 699, 830]. Concentration [1059, 56, 577, 85, 754, 405, 439, 500]. concentrations [331]. concentric [1074]. concept [443, 635, 984]. concepts [57, 138]. Concerning [866, 6, 955, 954, 1004, 1041, 56, 89, 844, 469, 536]. condition [327]. conditioned [248, 960]. conditioning [44, 49, 101, 194, 574]. conditions [877, 1030, 432, 557]. Conduction [975, 655, 766, 187, 195, 213, 30, 41, 978, 26, 34]. conductivity [901, 964]. conductor [491, 586]. conductors [502]. cone [692]. Conference [608, 1088]. confocal [1006]. conformations [968]. congestive [1093]. Congress [527, 663, 700, 1114, 1128]. connectedness [703]. connective [1028]. Connectivity [260, 992, 292]. Consciousness [277]. consequence [104]. consequences [295, 30, 41, 406, 627, 987]. Conservation [75, 959]. conservative [455, 399]. consideration [770]. Considerations [453, 897, 56, 89, 320, 896, 365, 330, 514, 478, 398, 40, 158, 290, 426, 927, 1019, 283, 807, 829]. constancy [526]. constant [523, 435, 194, 1096]. constants [1158, 767, 746, 1124]. constriction [69, 47, 8, 19]. Contagion [374, 332]. content [414, 931, 945, 437, 427, 642]. continuous [435, 1027, 733]. contours [462]. contractility [5, 810]. Contribution [216, 228, 143, 332, 374, 1107, 241, 293, 473, 561, 1082, 925, 15, 29, 353, 508, 121, 133, 373, 481, 548, 571, 952, 988, 150, 479, 1107]. Contributions [268, 333, 324, 372, 564, 10, 442, 469, 576, 23, 33, 28, 127, 426, 460, 534, 601, 1053, 1060, 1081, 163, 180, 1099, 305, 27, 79, 154, 547]. contributors [749]. Control [716, 810, 1004, 924, 1021, 916, 754, 824, 486, 222]. controlled [1106, 674]. convection [1145]. Convective [25, 1016]. convergence [756]. converging [792, 267]. Cooperative [728, 833]. cornea [1016, 948]. corneal [1091]. coronary [1002]. correction [1047, 436, 543]. Correlation [783]. correlations [52]. cortical [761]. counter [284]. countercurrent [400]. counters [1155]. counts [337, 336]. couple [1029]. coupled [766, 210]. Course [651, 1041, 1000, 1075, 631, 893, 596, 1039, 354]. crime [832, 824]. criteria [730, 493]. critical [463, 339]. croissance [918]. Cross [448, 392]. Cross-section [448, 392]. crown [156, 159]. Crystalline [209]. cultural [254, 296]. cultures [827]. Cunningham [31]. current [1119, 530, 637, 194, 502, 1143, 223, 415, 300, 491, 586, 615]. currents [1139]. curvature [1041, 1000, 1075]. curve [669, 577, 251, 535, 712]. curves [450, 741, 731, 1129, 1062, 1063, 315, 1058, 284]. Cybernetic [1020]. Cybernetics [222, 771]. Cycle [200, 1116, 567, 611]. cycles [198, 275]. Cyclic [1011, 1057]. cylinder [641]. cylinders [969, 711]. cylindrical [936]. cytochemical [433]. cytochrome [766, 1029]. Cytokinetics [890].

d’activités [1087]. damage [265]. dans [675]. D’Arcy [623]. data [530, 1101, 1023, 1014, 1009]. day [502]. dead [1047, 902, 522]. death [483]. Decay [826, 855, 998]. December [719]. Decision [425, 1111]. decomposition [1141]. deep [665]. defect [807]. Définition [1045, 537, 1045]. Definitions [493, 1090, 48]. Deformation [39, 82, 33]. degree [1079, 433]. delay [839]. Delayed [81, 813]. delta [449]. demyelinated [975]. density [1119, 959, 855, 532, 270, 659, 307]. dental [828, 829, 830]. deoxyribonucleic [968]. dependence [1080, 893, 839]. dependent [500, 674, 1017, 346]. deplacant [1045]. deposition [515]. deprivation [1150]. depth [55]. Derivation [846, 649, 588, 827, 910, 62, 535, 623, 350]. derived [1080, 385, 1023]. description [165, 174, 340, 881, 339]. Design [794, 1005, 714, 762, 570, 618, 646, 670, 623, 364]. desoxyribonucleic [480]. detecting [679]. Determination [746, 1124, 270, 282, 355, 1116, 1158, 592, 1036, 767, 1022, 380, 1015, 249, 1037]. determinations [782, 834, 261]. determined [139, 601]. determining [181]. deuterium [61]. deux [1138, 1045]. Development [735, 715, 751, 925, 484, 373, 943, 1039, 843, 830]. developments [254, 296]. diagnosis [1111]. Diagnostic [971, 1109]. diagram [129]. diameters [856]. diastole [669, 1144]. diastolic [612]. difference [138, 331, 378, 146]. Different [437, 484, 1078, 52]. differential [862, 861, 939, 754, 70, 541, 647, 146]. differentiating [199]. differentiation [778, 605]. diffraction [888]. diffusing [288]. Diffusion [64, 234, 1125, 405, 412, 433, 439, 440, 271, 13, 81, 844, 1158, 1010, 855, 438, 1059, 243, 313, 325, 335, 312, 940, 1097, 1098, 919, 976, 745, 71, 400, 742, 743, 66, 316, 380, 643, 500, 624, 997, 1078, 261, 282, 285, 1152, 508, 218, 212, 211, 250, 1145, 355, 14, 16, 25, 151]. diffusion-reaction [940]. digest [424]. digestion [705]. digraphs [996, 931]. Dilation [532]. Dilation-producing [532]. dilute [395]. dilution [557, 197]. Dimensional [593, 606, 793, 808, 337, 969, 336]. dimensionality [987]. dinner [719]. dioxide [278]. dipole [1129, 1139, 961, 501]. discharge [319, 1057]. discharges [1133]. Discrimination [24, 247, 143, 21, 38, 15, 29, 44, 49, 101, 495, 236, 548, 626, 654, 55, 21]. discs [969]. discussion [1107, 593, 336]. disorder [398]. disorders [1132]. dispersal [509, 465, 510]. dispersion [1034, 838]. dispersionless [1068]. dissipation [887, 932]. dissociable [587]. dissolution [828]. distance [347, 1045, 258]. distances [38]. Distefano [1004]. distensibility [982, 592]. distensible [351, 132]. distension [1035, 1065, 1144]. distinct [262, 464]. distributed [692, 939, 337, 1151, 317, 336]. distribution [459, 744, 471, 886, 887, 896, 917, 932, 552, 911, 551, 60, 231, 383, 785, 415, 883, 938, 176, 262, 444, 969, 1071, 300, 341, 58, 733, 119]. distributions [200, 464]. disturbance [351]. diuresis [841]. diving [665]. division [318, 84, 69, 96, 3, 7, 18, 47, 100, 193, 210, 298, 301, 113, 42, 8, 19]. DNA [480, 837, 503, 518, 543, 589, 811, 888, 967]. DNA-protein [503]. d’objectivation [1087]. doctoral [287, 310, 388]. domain [921]. dominance [375, 246, 269, 327, 735]. Donnan [309, 329]. d’organisations [1087]. dosage [161]. Dose [839, 315]. dose-effect [315]. Dose-rate [839]. double [1129, 614, 501]. double-dipole-layer [1129]. double-layer [501]. drag [211]. drive [645]. driving [524, 519, 548, 547, 561, 626, 712, 684, 848, 849, 926, 1038, 1044]. droite [1045]. droplets [237, 244]. Drosophila [859]. drug [551, 1022, 162, 161, 478, 1037]. drugs [592, 576, 629]. drum [173]. drying [352, 364]. Dual [753]. due [83, 475]. d’un [1107]. d’une [918, 1045]. duration [517, 611]. during [669, 552, 343, 742, 84, 47, 1144, 42]. dyes [474]. Dynamic [1070, 91]. Dynamical [713, 1135, 1141]. Dynamics [135, 1104, 186, 1127, 129, 701, 819, 818, 422, 47, 286, 544, 498].

each [437, 858]. ear [170, 173]. Eccentric [586, 501]. eclosion [859]. ecological [1024]. ecology [470]. editor [915, 750]. Effect [1016, 911, 901, 964, 172, 1047, 1012, 315, 107, 479, 1022, 463, 574, 149, 415, 576, 257, 296, 1002, 1035, 1065, 246, 1018]. Effects [447, 884, 361, 1046, 738, 1144, 731, 384, 844, 453, 592, 886, 441, 801, 194, 970, 269, 160, 454, 907, 1027, 758, 998, 68, 376, 444]. efficiency [129, 469]. egg [82]. egoistic [571]. Eighth [1128]. Einthoven [962]. Elastic [994, 715, 741, 751, 530, 612, 981, 860, 279, 132, 193, 774, 775]. elasticity [530, 429]. elastomeric [371]. elations [148]. electric [761, 683, 693, 728, 833]. Electrical [637, 256, 50, 14, 73, 77]. electrocardiographic [562, 622, 787]. electrocardiography [502, 964, 806].

Electroencephalography [1128]. electrolyte [624]. electrolytes [71, 285, 1049, 14, 50]. electromagnetic [1015]. electromagnetophoresis [937]. electron [655, 766, 747, 1029, 1080, 474, 572, 874]. electronic [564]. electrophysiological [897]. electrophysiology [843, 844]. electroplethysmographic [487]. electroretinogram [652]. Elementary [656, 682, 283]. elements [562, 566, 756, 58, 918, 918]. ellipsoidal [4, 20]. Elongation [8, 19, 69, 193, 4, 20]. Elovich [1080, 1147]. embryogenesis [1063]. emotional [958]. emphasis [268]. empirically [431]. empty [712]. enamel [828, 829, 830]. Encephalitis [83]. enclosed [371]. endocrine [451, 1121]. endocrinological [516]. energetics [1157]. Energy [1042, 1029, 887, 932, 1050, 242, 469, 617, 1093, 113, 1074]. engineering [651, 1008]. entre [1045]. Entropy [928, 931, 945, 957, 733, 957]. entry [161]. environment [454, 575]. environmental [296]. environmentally [644]. Enzymatic [653, 655, 825]. Enzyme [357, 385, 384, 1021, 655, 862, 565, 587, 97, 933, 627, 1085]. enzymes [766, 209]. enzymic [1106]. epidemics [146]. epilepsies [1132]. epileptic [1122]. epimorphism [570, 868]. epsilon [449]. equation [1080, 1147, 861, 313, 325, 335, 340, 940, 826, 108, 622, 378, 929, 588, 827, 62, 787, 70, 529, 528, 668, 671, 784, 723, 146]. Equations [744, 947, 843, 844, 401, 862, 886, 917, 855, 649, 760, 781, 1059, 248, 385, 939, 1135, 403, 754, 405, 412, 439, 440, 406, 1049, 122, 140, 541, 585, 647, 681, 119]. equilibria [1003, 1019, 1090]. Equilibrium [877, 85, 309, 459, 900, 746, 329, 62, 829]. Equivalence [115, 26]. equivalent [614, 1143, 736]. Ernesto [1066]. Errata [1065]. Erratum [447, 448, 776, 873]. error [679]. error-detecting [679]. Errors [49]. Ertragsverlauf [577]. Erythrocyte [886, 887, 896, 917, 932, 249]. erythrocytes [673]. essais [1040]. establishability [823, 720]. estimating [1074]. Estimation [982, 779, 487, 1084]. estimations [1043]. étages [791]. étude [1107, 791]. Evaluation [1014, 857, 888]. event [233]. Events [538, 893]. evidence [900, 1029]. Evolution [753, 196, 734]. exact [743, 292, 337]. examinations [430]. exchange [366, 434, 1125, 142, 204, 1155, 290, 350, 299, 354, 113]. exchangers [637]. exchanges [124, 122, 123, 140]. excitable [1139]. excitation [311, 349, 12, 98, 115, 126, 268, 319, 708, 295, 360, 242, 252, 155, 978, 1092, 68]. exclusion [727]. excretional [130]. exercise [592]. exhibits [1122]. exists [405, 439]. expansion [884]. expansions [393]. expectations [665]. expenditure [469, 617, 1074]. experiment [1158, 61]. Experimental [470, 1101, 181, 368]. experimenter [674]. experiments [455, 1119, 399, 531, 781, 815, 861, 377, 674, 941, 336]. expiration [904]. explanation [83, 946]. exponential [855, 315, 625, 998]. expressed [754]. expression [82, 812]. expressions [204]. expulsion [994]. extension [376, 392, 447, 448, 450, 31]. extensions [499]. external [344, 350]. externally [1012]. extinct [863]. Extinction [977]. extra [900]. extra-cellular [900]. extracellular [1013]. extracorporal [881]. extraocular [485]. extremities [158]. eye [846, 1061, 521, 774]. eye-positioning [846].

F. [954]. facilitation [253]. facon [1045]. factor [871, 126, 295, 360, 164, 223, 951]. Factors [462, 186, 857, 904, 269, 1009, 142, 257, 296, 601]. failure [1093]. farmaci [629]. favorable [863]. features [1122]. Feedback [513, 1004, 916, 990, 1030, 1121, 810, 1070]. fellowship [449]. Fellowships [599, 590, 620, 664]. fertilized [812]. Fiber [1144, 48, 54, 59, 65, 1143]. fibers [1041, 60, 213, 415, 1000, 1035, 1056, 1075, 1131, 1065]. Fick [243]. field [649, 760, 683, 693, 728, 833, 622, 787, 678]. fields [801, 761, 85, 1015, 16, 25]. file [745]. filling [1094]. findings [147]. Finite [851, 743]. firing [855]. First [1151, 1079, 1084, 791]. fishes [555, 470, 511]. fit [1129]. fitness [1024]. fixation [673]. fixations [633]. fixed [624, 405, 439, 521]. flicker [463]. flour [396, 404]. flow [875, 966, 1016, 151, 843, 844, 792, 886, 887, 896, 917, 932, 980, 965, 328, 1123, 343, 625, 742, 487, 985, 407, 641, 279, 469, 132, 677, 885, 557, 197, 1086, 775, 25, 944, 1025, 932]. flowgraph [1070]. flowing [351, 350]. flows [1096]. fluid [980, 371, 1050, 279, 883, 36, 61, 132, 869, 774]. fluids [37]. flux [549]. flux-force [549]. fluxes [1033, 637]. FMN [572]. FMNH [572]. Foerster [784, 723]. Fokker [668, 671]. food [555]. force [549, 31]. Forced [769, 904]. forces [66, 999, 211, 1103, 870, 858, 14]. Forecasting [1110]. forget [866]. Form [102, 187, 195, 120, 125, 128, 23, 33, 28, 32, 95, 110, 180, 70, 628]. formal [34]. formalized [755]. formation [331, 841, 563, 830, 23]. forme [1107]. Forms [176, 45, 86, 178]. formula [251]. Formulation [972, 565, 587, 937, 498]. foundations [417]. Fourth [1114]. fraction [1116, 900]. fragment [1014]. fragmentation [814]. frameworks [1070]. free [814]. freedom [488]. Freese [589]. French [918, 1107, 1138, 613, 675, 705, 820, 791, 1045, 1087, 1040, 424]. frequencies [259, 446, 428]. frequency [921, 463, 517, 415, 998, 879, 1064]. frozen [352, 364]. function [64, 453, 161, 176, 181, 618, 299]. functions [1062, 1009, 938, 678, 724, 914, 78]. fundamental [70, 122, 140, 529, 528]. Further [514, 27, 79, 154, 547, 800, 842, 518, 543, 604]. fusion [517].

Galvanic [53]. game [435, 420, 274]. game-theoretical [420]. ganglion [225]. Garshin [850]. gas [1047, 1036, 1097, 1098, 124, 142, 785, 1152]. gases [556, 559, 582, 204]. gastrulation [39]. gate [469]. Gaussian [471]. gel [1059, 729]. gel-like [1059]. gels [193]. gene [627]. gene-one [627]. General [871, 963, 1109, 666, 672, 531, 918, 295, 62, 36, 61, 92, 93, 387, 397, 481, 845, 927, 37, 815, 166, 593]. générale [918]. generalization [429, 243, 386, 723, 31]. Generalizations [1147, 469]. generalized [766, 965, 1135, 1017, 647]. generating [724]. generation [480]. generator [614, 1143]. generators [502, 491, 586, 615]. genetic [1024, 546, 595, 657, 688]. genetics [793, 808, 1017, 1117, 696, 737]. geometric [862]. Geometrical [1131, 1041, 1000, 1075]. geometrization [421, 436, 467]. German [893, 479, 577, 638, 648]. Gestalt [6]. Glucose [567, 722, 1112, 983]. glycolytic [235]. Gompertz [1063]. gossip [274]. govern [142]. governed [889]. gradient [1099]. gradual [884]. Graduate [609, 650, 748, 568]. Granduate [580]. graph [1014, 928, 426, 496, 661]. graph-theoretical [496]. graphical [633, 957]. graphs [922, 906, 928, 931, 945, 957, 942, 639, 437, 427, 945]. Green [914, 78]. Grenzfall [638]. grosser [638]. ground [617]. Group [688, 993, 383, 215, 812]. groups [996, 389, 298, 597, 952, 988]. growing [947]. Growth [876, 102, 401, 441, 778, 826, 221, 108, 191, 220, 465, 510, 813, 918, 827, 130, 153, 938, 406, 11, 17, 303, 346, 864, 634, 150, 70, 1068, 585, 647, 681].

half [162]. half-life [162]. Hamiltonian [94]. Harold [789]. health [257, 1073]. heart [962, 295, 360, 1046, 691, 738, 1093, 1035, 1065, 1144, 475]. hearts [1094]. heat [744, 247]. hedonistic [290]. hemodialysis [545]. hemoglobin [844, 582]. Henri [987]. herbicidal [189]. Heredity [322]. heterarchy [139]. heterogeneous [1097, 1098, 919, 976, 1048]. hierarchical [291]. Hill [115]. histograms [837]. historical [373]. history [334, 345]. hits [262, 464]. Hodgkin [820]. Holgate [977]. homeostasis [1154]. Homeostatic [754, 419]. homogeneous [815]. honey [648]. honeycomb [648]. Honigbiene [648]. honor [719]. Hooke [429]. hormone [990]. hormones [453, 779]. horopter [35]. host [701]. host-pathogen [701]. human [846, 530, 592, 612, 989, 856, 993, 237, 244, 643, 1152, 1159, 1070, 1118, 774, 694, 475]. humidity [967]. Huxley [820]. hybrid [1070]. hydration [889]. Hydrodynamics [752, 1025]. hydrostatic [172]. hydroxyapatite [829]. Hyman [850]. hypertrophied [1094]. hypertrophy [1144]. Hypothalamic [990]. hypothesis [471, 887, 990, 385, 1050, 233, 627, 987, 589].

ideal [637]. ideas [103]. identification [921, 1058]. ideologies [286]. ignition [307, 281]. II [448, 392, 455, 1072, 523, 559, 672, 151, 174, 844, 409, 195, 887, 159, 173, 228, 636, 760, 217, 325, 940, 587, 1098, 21, 54, 551, 572, 693, 15, 1, 49, 175, 244, 404, 428, 269, 891, 360, 686, 685, 808, 931, 337, 412, 440, 125, 149, 469, 171, 208, 359, 28, 93, 132, 148, 224, 548, 596, 627, 799, 895, 927, 1018, 1081, 935, 507, 703, 604, 150, 123, 606, 539, 671, 784, 882, 652, 77, 949, 829]. III [450, 1105, 172, 416, 896, 238, 565, 605, 38, 59, 29, 101, 526, 327, 892, 124, 945, 433, 128, 576, 232, 368, 33, 894, 926, 1082, 1156, 803, 1052, 830]. illness [958, 1110]. illusion [280]. im [638]. imbedded [815]. Imitative [317, 342, 264, 444, 460, 533, 596, 601, 759, 772, 913, 386]. immanent [103]. immunotaxonomy [934]. impedance [58]. impermeable [96]. implication [402]. implications [1111, 841, 62, 941]. implicit [525]. Important [749]. impulse [1042]. In-vivo [1043]. inaccessible [815]. inactivation [97]. incidence [454]. income [444]. Incompletely [1108]. incompressible [371, 279, 37]. inconsistencies [502]. incorporating [813]. incorporation [480]. increase [1061]. increases [577, 1077]. increasing [300]. index [221, 162, 928]. indicator [1034, 557, 197]. indicator-dilution [557]. individual [1041, 488, 544, 1000, 1035, 1065, 1056, 1075, 1131]. individuals [866, 176, 169, 171, 168, 298, 290, 560]. induced [875, 107, 428, 839, 575, 644]. Induction [605, 565]. industrial [631]. inelastic [385]. inequality [393]. Inert [1097, 1098, 124, 142, 204]. inert-gas [124]. infinite [743, 931]. infinitely [638]. Influence [705, 1033, 465, 295, 360, 560, 705]. influences [660]. Information [630, 374, 658, 382, 737, 642, 631, 651, 893, 414, 347, 332, 289, 323, 931, 945, 1040, 358, 359, 368, 391, 418, 443, 264, 410, 553, 578, 274, 540, 604, 437, 427, 660, 492]. informations [1040]. Informationsverarbeitung [893]. inherent [246]. inheritance [877]. Inhibition [385, 462, 1030, 251]. initial [1119]. initially [936]. injectable [861]. injustice [1031]. Input [284, 225, 251, 755]. Input-output [225, 251]. inputs [267, 1057]. instantaneous [982, 1043]. Institute [630, 568]. insulin [1112, 983]. intact [1046]. Integral [781, 340, 529, 528, 1086, 339]. integral-equation [340]. Integro [861]. Integro-differential [861]. intense [1061]. interacting [461, 1141]. Interaction [198, 556, 559, 1112, 453, 478, 566, 168, 72]. Interactions [825, 179, 130, 169, 171, 571]. intercellular [869]. interchange [89]. interdisciplinary [764]. interest [538, 168]. interesting [468]. interface [1080]. interfaces [747, 901, 964]. interference [43]. intermediates [250]. intermittent [794]. internal [242]. International [527, 663, 700, 1114, 1089, 1128, 1005, 1088, 1076, 458]. Interpretation [1009, 315, 920, 179, 276, 342, 377, 463, 495, 256, 154, 1102, 706, 521, 685]. intersymbol [660]. interval [1026]. intervals [241]. intestinal [995, 180]. intra [900]. intra-vacuolar [900]. intraocular [1127]. intrapulmonary [785]. introduction [1157]. invariant [796]. invertase [273]. investigation [875, 862, 339]. Ion [1013, 766, 747, 1080, 552, 330, 637, 331]. iona [833]. Ionic [999, 637]. ionization [270]. ionizing [905, 974, 271]. ions [655, 438, 745, 683, 693, 707, 717, 728, 234]. irradiated [905, 837]. irradiation [441, 974, 270]. isohemagglutination [633]. isohemagglutinins [673]. isometric [129]. isotope [480]. isotropes [675]. isotropic [675]. issues [872]. Italian [629]. IV [917, 531, 239, 65, 814, 957, 1038, 55]. Ivlev [763].

joint [989]. Journal [579, 698]. journals [75]. judgments [597]. July [789].

Katchalsky [1049]. Kedem [1049]. Kernleiter [46]. kidney [1051]. kidney-ureter [1051]. kidneys [857]. kinds [484]. Kinematics [989]. Kinetic [730, 1158, 655, 767, 66, 1015, 1136]. Kinetics [545, 480, 1105, 1083, 370, 1012, 862, 314, 531, 202, 214, 217, 238, 239, 326, 385, 97, 713, 1135, 245, 933, 1078, 983, 1151, 273, 864, 629, 828]. knee [989]. know [178]. Kommunikation [893]. labeled [1116, 350]. labeling [838]. labor [298]. labyrinthine [164]. Lakhani [977]. laminar [279]. Landahl [236]. Landau [850]. Laplace [1023, 498]. large [638, 997]. latency [588]. law [243, 910, 1068, 31, 429]. laws [1135, 702]. layer [1129, 1074, 501]. layers [614]. lead [562, 962]. Leadership [1060]. leakage [900]. Learned [984]. learner [591]. learning [991, 302, 143, 795, 574, 687, 689, 674, 685, 443, 535, 645, 832, 960]. left [1041, 982, 1043, 1006, 981, 1046, 1093, 821, 822, 882, 1052, 1000, 1035, 1065, 1075, 1144, 1074]. Legendre [393]. lemmas [48]. length [129, 638, 611, 1035, 1065]. length-tension [129]. lengths [38, 1094]. lethargica [83]. Letter [750, 719]. level [617]. levels [770]. Lie [993]. Life [410, 553, 578, 480, 162, 109, 387, 397, 512, 867, 935, 1156]. light [859, 147, 1061]. like [783, 229, 1059]. likelihood [1084]. limit [638]. limitations [665]. limiting [665, 289, 323]. line [863, 1045]. Linear [919, 483, 326, 403, 1137, 674, 756, 754, 87, 91, 294, 1085, 529, 528, 539, 1136, 541, 68, 836]. linearly [300]. linegraph [1070]. linegraph-flowgraph [1070]. linguistics [660]. link [289, 323]. linkage [216, 228]. Linnaean [851]. liquid [792, 747, 351, 393, 350, 775]. liquid-solid [747]. Living [361, 1004, 982, 592, 916, 398, 1040]. locality [492]. localization [357, 433]. locomotion [135, 185]. Logic [527, 663, 700, 240]. Logical [445, 466, 1005, 778, 104, 103, 145, 152, 525]. logistic [509]. long [994]. Longitudinal [742, 300]. looking [621, 636]. Lotka [573]. lower [136, 770]. lung [902, 1023, 237, 124, 1152, 618, 1145]. lungs [522].

machine [222]. machines [805, 923, 690, 703, 803]. macromolecular [655]. macromolecules [1028, 753]. macroscopic [937]. magnetic [801]. maintenance [1024]. make [431]. malarial [106, 118]. mammalian [625]. mammals [185]. mammillary [746, 1124]. man [278, 362, 982]. mapping [798]. mappings [658]. March [789]. Markoff [483]. masking [854]. Mass [1135, 390, 351, 293, 831, 1071]. Mass-action [1135]. matching [678]. material [711, 1130]. materials [883]. maternal [970]. Mathematical [558, 569, 872, 1089, 715, 751, 556, 559, 141, 53, 363, 451, 413, 45, 183, 324, 1, 304, 428, 526, 374, 785, 983, 1155, 169, 171, 5, 11, 17, 100, 145, 219, 224, 272, 298, 524, 645, 712, 848, 869, 926, 972, 1018, 1076, 1038, 1044, 607, 955, 184, 662, 498, 516, 878, 807, 411, 784, 954, 580, 590, 599, 609, 620, 664, 550, 1109, 542, 165, 174, 111, 925, 278, 362, 904, 621, 636, 1095, 565, 587, 1023, 920, 12, 551, 268, 305, 333, 372, 191, 10, 15, 29, 44]. mathematical [49, 69, 101, 118, 396, 404, 566, 1079, 332, 658, 105, 985, 484, 502, 793, 808, 30, 41, 252, 353, 859, 1154, 240, 293, 308, 736, 22, 23, 33, 28, 40, 51, 74, 79, 67, 88, 92, 93, 121, 127, 133, 144, 137, 131, 132, 147, 154, 148, 177, 168, 167, 230, 226, 257, 254, 334, 373, 387, 397, 444, 454, 481, 488, 511, 519, 548, 547, 584, 598, 561, 627, 684, 849, 913, 952, 988, 1002, 1031, 1133, 255, 853, 903, 1153, 150, 350, 299, 163, 593, 606, 76, 780]. mathematical [840, 386, 521, 774, 694, 828, 829, 830]. mathematical-physical [502]. mathematics [662]. matrices [190, 499]. matrix [798, 581, 157, 175, 992, 266]. Maximum [1084, 692, 746, 113]. may [1029, 168]. McCulloch [247]. mean [827]. meaning [365]. Measurement [1155, 1029, 1033, 221, 365, 469, 197]. measurements [1116, 532]. measures [957]. mechanical [450, 186, 904, 156, 159, 696, 840, 858, 774]. mechanics [461, 514, 536, 4, 20]. mechanism [1004, 229, 357, 1021, 846, 187, 195, 419, 170, 173, 21, 38, 330, 400, 96, 231, 245, 280, 112, 469, 3, 7, 18, 152, 432, 729, 758, 1093, 55]. mechanisms [550, 924, 730, 138, 1126, 1160, 595, 978]. media [81, 64, 815, 1059, 13, 1113]. mediation [1127]. mediators [295, 360]. Medical [456, 1109, 1111]. medicine [482, 536]. medium [1048, 624, 1078, 261]. Meeting [790]. meetings [835]. members [735]. membrane [1129, 1119, 311, 349, 117, 151, 172, 1012, 413, 552, 1033, 745, 268, 96, 203, 331, 1125, 62, 999, 261, 300, 874, 321]. membrane-bounded [1125]. membranes [1139, 655, 1013, 438, 549]. memory [941]. mental [958, 144, 148, 1073]. message [1142]. metabolic [312, 565, 587, 1108, 250]. metabolism [320, 28, 567, 14, 50]. metabolites [89, 779]. metabolizing [165, 174, 340, 66, 201, 163, 339]. metamorphosis [1003]. metaphase [839]. method [592, 612, 316, 380, 378, 500, 624, 633, 405, 439, 1096, 1069, 540, 604, 292, 1086, 1087]. Methode [1087]. Methodology [527, 663, 700]. Methods [348, 1023, 502, 557, 1037]. metric [773]. microbial [1071]. microphysical [595]. microscopy [474]. microstructures [675]. midcentury [456]. middle [170, 173]. migration [985]. mineral [479]. Mineralischer [479]. minimal [887, 932, 1050, 594]. minority [597]. mitoses [1116]. mitosis [226]. mitotic [221, 369]. Mixed [374, 332, 1108, 567]. mixed-cycle [567]. mixing [1033, 725]. mobile [572, 564]. mobiles [1045]. mobilities [1015]. mobility [1029]. mode [161]. Model [722, 955, 847, 954, 715, 751, 1129, 1109, 1024, 924, 1106, 846, 916, 441, 902, 961, 855, 1101, 621, 636, 974, 565, 587, 761, 778, 745, 795, 330, 319, 1022, 396, 404, 574, 687, 697, 854, 1079, 813, 1107, 769, 985, 686, 880, 701, 948, 975, 613, 30, 41, 1151, 762, 908, 859, 1154, 695, 740, 800, 1133, 346, 960, 1057, 853, 611, 604, 998, 903, 1153, 704, 865, 878, 780, 1071, 386, 774, 522, 694, 1074, 1001, 830, 873]. model-system [565, 587]. modèle [1107, 613]. modeles [820]. modeling [965, 1111]. Models [930, 907, 1026, 1027, 513, 1139, 1011, 489, 523, 542, 871, 451, 413, 1023, 912, 566, 689, 674, 785, 820, 983, 1155, 359, 736, 283, 1085, 840, 784, 1073]. modes [1011]. modification [295, 360]. modified [934]. modifier [933]. Möglichkeiten [893]. Molecular [550, 1105, 864, 978, 40, 520, 554, 584, 598, 627]. molecule [811]. molecules [679, 378, 288]. Monnier [115]. Monocular [55]. monographs [506]. monoids [814]. Monte [1062]. morphogenesis [116]. morphogens [951]. mortality [668, 671]. motion [164, 821, 882, 1052, 37]. motivated [793, 808, 176, 290]. motivation [169, 171]. motivational [179]. motor [521]. movement [42]. movements [111, 45, 5, 99, 182]. moving [1045, 521]. mucous [875]. multi [815, 778, 793, 808, 1017]. multi-celled [778]. multi-compartment [815]. multi-dimensional [793, 808]. multi-type [1017]. multicellular [1013]. multicompartment [861, 1136]. multidimensional [819]. Multiflux [539]. Multiple [1003, 779, 1111, 633]. multiplication [267]. multiplier [857]. Multipole [491, 501, 806, 586, 615]. multistable [1072]. multitype [1117]. muscle [731, 900, 1080, 981, 429, 129, 1143, 282, 485, 416]. Muscular [129]. musculoskeletal [1070]. musical [74, 67]. mutant [616, 341]. mutation [701, 863]. mutations [1095]. mutual [560]. myelinated [704, 865]. myocardial [1041, 1000, 1035, 1065, 1056, 1075, 1131]. myocardium [810, 1130].

Nährstoffe [479]. Nährstoffkonzentration [577]. nasal [244]. National [790]. nature [52, 512, 987]. Nelson [820]. nematic [369]. nephron [625, 742]. Nerve [30, 41, 1129, 187, 195, 413, 48, 54, 59, 65, 60, 213, 380, 1042, 415, 155, 355, 704, 865, 225, 300, 874, 978, 1092, 34]. nerve-fiber [48, 54, 59, 65]. nerve-ganglion [225]. nerves [295, 360]. Nervous [1121, 141, 115, 126, 10, 104, 103, 139, 9, 147, 154, 1002, 1134, 199, 188, 486]. net [1011, 247, 687, 854, 306, 307]. Nets [258, 490, 649, 760, 104, 114, 157, 175, 297, 139, 762, 794, 200, 207, 208, 241, 281, 473, 184, 215, 266, 275, 260, 292, 628, 190]. network [291]. networks [445, 466, 621, 636, 48, 54, 59, 65, 87, 91, 391, 418, 356]. Neural [138, 1011, 247, 621, 636, 649, 760, 1127, 21, 38, 114, 157, 175, 566, 687, 854, 762, 794, 1087, 241, 277, 152, 432, 1122, 960, 184, 215, 283, 55, 628, 190, 462, 636]. neurobiophysical [179, 276, 342, 695, 800]. Neuroendocrine [1004, 916]. Neuron [955, 954, 855, 613, 820, 87, 80, 91, 1133, 853, 998]. neuronal [289, 323, 1057]. neuronales [1087]. neurone [613, 820]. neurons [267, 284, 72, 755, 984]. neurophysiological [147, 740]. Neurophysiology [1128, 513]. neutral [508]. never [866]. news [866]. Nicolas [1115, 718]. NMR [900]. Non [68, 455, 900, 435, 399, 815, 126, 637, 85, 469, 988, 1049, 1141, 455]. Non-accommodative [68]. non-carrier [469]. non-complexed [900]. non-conservative [399, 455]. non-constant-sum [435]. non-electrolytes [1049]. non-equilibrium [900]. non-homogeneous [815]. non-ideal [637]. non-interacting [1141]. Non-linear [68]. non-normal [126]. non-steady-state [455]. non-uniform [85]. non-violent [988]. nondimensional [796]. nonhomogeneity [1046]. noninstantaneous [1033]. Nonlinear [312, 1043, 401, 212, 891]. nonoriented [942]. nonperiodic [1100, 1121]. nonuniform [785]. nonuniformly [317]. normal [993, 126, 1027, 1094]. North [640]. Note [915, 57, 390, 98, 108, 389, 235, 231, 329, 342, 377, 403, 517, 970, 1137, 393, 838, 51, 94, 212, 394, 654, 303, 476, 541, 2, 321, 253, 509, 370, 111, 977, 980, 247, 248, 340, 35, 71, 328, 625, 682, 114, 203, 245, 495, 510, 566, 643, 251, 910, 134, 223, 736, 52, 177, 205, 206, 211, 236, 264, 263, 467, 497, 488, 512, 504, 617, 667, 771, 772, 817, 849, 943, 1002, 1100, 1132, 1121, 870, 543, 603, 657, 804, 536, 155, 812, 275, 427, 423, 452, 1068, 824, 485, 486]. notes [653]. notice [789]. notion [525]. nucleic [679, 1032, 814]. number [906, 49, 1009, 838, 415, 601]. numbers [796]. numerical [1158]. numero} [1112]. nutrient [577]. nutrients [479]. nutrition [470, 511]. nutritional [130].

O [556, 844]. Obituary [1115, 789]. objectivation [1087]. oblique [153]. observation [1077]. Observations [1157, 487, 80]. obstructed [857]. obtained [532, 319, 487]. occur [307]. occurs [412, 440]. offers [764]. offspring [911]. Oklahoma [748]. Old [579]. one [933, 295, 360, 337, 412, 440, 596, 627, 542]. one-dimensional [337]. one-factor [295, 360]. One-organ [542]. only [1053]. Onsager [243]. onset [828, 829, 830]. open [370, 1007, 294, 1085]. operated [1027]. Operations [790, 835]. oppressed [952, 988]. optic [187, 195]. Optimal [367, 409, 416, 912, 1149, 302, 714, 432, 623]. optimum [156, 159]. order [398, 1084, 1151]. order-disorder [398]. orders [852]. organ [542, 453, 130]. organic [474, 196, 23, 33, 28, 32, 110, 180]. organism [991, 191, 70]. Organismic [963, 1072, 1104, 868, 845, 927, 1159, 943, 1069, 1053, 1060, 1039, 1081, 1082, 1150]. organisms [778, 997, 353, 52, 786, 845]. organization [1028, 57, 788]. organizations [1087]. organizing [608, 658]. organs [658]. orientation [353]. oriented [383]. origin [761, 109, 512]. orthotropic [860]. Oscillating [1007]. oscillations [1030, 1137, 769, 1100, 1121]. oscillator [762]. osmotic [889, 210]. other [794, 1003]. Ouchterlony [934]. our [749, 987]. Out-of-print [872]. Outline [118, 157, 227, 232, 92, 93, 137, 334, 454, 845, 986, 255]. Outlines [338]. output [251, 176, 284, 1093, 225, 755]. oxidase [766, 1029]. oxidation [655, 1029, 1012]. oxidation-reduction [655]. oxide [61]. Oxygen [997, 1016, 56, 362, 1101, 1099]. oxygenation [881]. oxygenator [881].

P. [1041]. packed [262]. pairs [996]. Pan [579]. panmictic [877]. paper [977]. paradigm [907]. paradox [809, 525]. parallel [25]. parameter [1096]. parameters [1116, 796, 1058, 1015, 356, 1131]. parasites [106, 118]. Parasitism [435, 420]. part [625, 173, 120, 125, 128, 149, 166, 784]. partial [995, 359]. Particle [643]. particles [985, 337, 508, 336]. particular [268, 84, 396, 404, 643, 360]. particulate [766]. particulates [643]. parties [601]. partly [861]. parts [1051]. passages [244]. Passing [548]. patches [1024]. Path [674]. Path-dependent [674]. pathogen [701]. pathological [1133, 706]. pathways [155]. patient [592]. pattern [396, 404]. patterns [641, 1057, 888]. or [766]. perceived [38]. perception [187, 195, 692, 462, 178, 74, 67, 88, 121, 770, 55, 602, 909, 949]. perceptron [682]. perfect [591]. performance [665]. perfusion [995, 1097, 1098]. Periodic [72, 725, 775, 968, 363, 326, 403]. Periodicities [294, 381, 205, 201, 869, 972, 973, 1018, 2]. Periodicity [647, 112]. periods [241]. peripheral [741]. Peristaltic [966]. permeability [117, 151, 172, 1048, 268, 203, 380, 249, 261, 282, 355, 321]. perseveration [495]. persistence [344]. person [435]. perspectives [482]. pertaining [502]. perturbation [781]. perturbation-tracer [781]. perturbed [1096]. pH [844]. phage [441]. phase [369, 707, 717, 393, 1152, 395]. phases [725]. Phenomena [558, 569, 253, 413, 268, 366, 854, 999, 281, 144, 148, 1003, 72]. phenomenon [1010, 1079]. Philosophy [527, 663, 700]. phosphatases [245]. phosphorylation [1029]. Photochemistry [632, 765]. photoreception [338]. photosensitivity [136]. Physical [196, 715, 751, 472, 502, 164]. physicochemical [319]. Physicomathematical [32, 697, 110]. Physics [816, 895, 94, 553, 578, 986, 635]. Physiological [1041, 1000, 1075, 513, 612, 1058, 134, 142, 869, 896]. pigeons [1026]. pigments [899]. piston [170]. Pitts [247]. pituitary [419]. Planck [668, 671]. plane [392, 448]. planktophage [555]. plantes [1107]. plants [229, 1107, 95]. plasma [983]. plasticity [9]. plateaux [471]. plates [934]. PLM [1062]. PLM-curves [1062]. Poincaré [987]. Point [1088, 1150, 113, 437, 423]. point-bases [423]. Poiseuille [980]. Poiseuille-type [980]. polarity [22, 27, 73, 77]. polarons [1029]. political [601]. polymer [1014]. Polymorphism [1003, 1024]. polynomial [393]. polysomic [877]. Polytechnic [568]. pool [1009]. Population [465, 374, 459, 1024, 866, 471, 877, 911, 396, 404, 510, 332, 813, 918, 1017, 1117, 358, 359, 368, 346, 812, 659, 585, 499]. populations [509, 959, 970, 819, 818, 317, 303, 1071, 1067, 947]. pore [743]. pores [508]. positioning [846]. positive [294]. possibilities [893, 679]. Possibility [874]. possible [229, 57, 106, 369, 49, 269, 112, 226, 210, 402, 426, 432, 454, 488, 520, 695, 740, 832, 869, 953, 942, 941, 642, 395]. possibly [1003]. Post [590, 599, 620, 664, 287, 310, 388]. Post- [599, 620, 664]. Post-and [590]. post-doctoral [287, 310, 388]. Postdoctoral [449, 477]. postulate [655, 870]. Potential [415, 552, 855, 365, 914, 78]. potentials [637, 707, 717, 50, 475]. potentiation [478]. Pourquoi [1138]. powdered [828]. power [910, 852]. practical [844]. pre [251]. pre-inhibition [251]. precursors [779]. predator [1156]. predator-prey [1156]. predicates [786]. predictions [431]. Predoctoral [599, 590, 620, 664]. prejudices [272, 255]. preliminary [56, 48]. preload [1056]. premiers [791]. prescribed [945]. presence [151, 843, 844, 1158]. present [502]. present-day [502]. preserving [991]. pressing [907, 1026, 1027]. Pressure [641, 669, 741, 731, 172, 1043, 889, 1008, 1091, 305, 324, 371, 407, 132, 611, 878]. pressure-volume [1043]. pressures [612, 210, 1094]. pressurevolume [1023]. prey [1156]. primary [125, 128, 595, 657]. primordial [426]. principle [573, 714, 594, 646, 670, 727, 113, 94]. principles [387, 397, 481, 607]. print [872]. privileged [1087]. privilégiées [1087]. Probabilistic [1111, 434, 160, 227, 241, 232, 154, 206]. probabilities [977, 685, 818]. Probability [905, 464, 288, 698, 466, 855, 638, 378, 863, 262, 553, 578]. problem [1083, 170, 191, 276, 234, 87, 91, 137, 131, 132, 168, 206, 218, 257, 290, 394, 454, 468, 570, 667, 503, 518, 543, 589, 275, 476, 723]. problems [243, 316, 297, 434, 500, 819, 818, 793, 808, 32, 88, 546, 662, 1020, 1142, 225, 539, 914, 78]. procedure [767, 181, 1070]. procedures [433, 368]. Proceedings [662]. process [489, 1109, 1083, 747, 899, 565, 587, 881, 819, 1117, 265]. processe [166]. Processes [1088, 483, 977, 565, 587, 1111, 793, 808, 1017, 149]. processing [893, 658]. produced [210, 1103, 1057, 1061]. producing [532]. products [996, 271]. Professor [719]. profile [1016]. profiles [1144]. program [676, 699]. programs [764]. prolate [1006]. Prolegomena [286]. propagation [351, 371, 293, 878, 1113]. Properties [940, 715, 741, 751, 923, 981, 403, 992, 1078, 680, 773, 604, 386, 760]. prophase [837]. prophase-trapping [837]. Proposals [1021, 963]. proposed [917, 729]. propositions [656]. protein [898, 814, 518, 543, 589, 503]. proteins [889]. protoplasm [57]. protoplasmic [76]. protrusions [76]. proximal [625]. psychoanalytic [920]. psychoneuroses [454]. psychophysical [15, 29, 236, 548, 626, 654]. psychophysics [702]. psychoses [105, 167]. pulmonary [1047, 1034]. pulsatile [1123]. pulsations [90]. Pulse [860, 936, 878]. pulses [944]. punishment [832]. pyruvate [766].

quadruped [158]. quadrupedal [135]. qualitative [1104]. Quantitative [1139, 565, 587, 881, 605, 345, 740]. quantities [732]. Quantum [1092, 1021, 1106, 564, 572, 696, 546, 657, 536]. quantum-mechanical [696]. quantum-theoretic [1021, 546, 657]. quasi [500, 897]. quasi-stationarity [897]. quasi-steady-state [500]. quasiergodic [638]. quasiergodischen [638].

R. [723]. Radial [976]. Radiation [361, 974, 839, 265]. radiation-induced [839]. radiations [905, 160, 271]. radicals [564]. radioactive [629, 711]. radioattivi [629]. radiobiological [666, 672]. Radiobiology [361, 234]. radiogenetics [233]. radiotracers [782, 834]. ramification [1107]. Random [344, 509, 465, 510, 297, 383, 200, 207, 208, 281, 306, 473, 275, 260, 292, 307, 734]. Randomization [241]. randomized [755]. randomly [337, 1045, 336]. Rapoport [251]. Rashevsky [1115, 115, 718, 925, 6, 108, 719, 870, 386, 423, 452, 26]. rate [523, 1158, 990, 767, 1124, 82, 782, 834, 839, 1026]. rates [779, 653, 1155, 296, 373]. rather [900, 889]. ratings [67]. rational [953]. rationale [961, 472]. rats [1026]. ray [888]. Re [823, 720, 888]. Re-establishability [823, 720]. re-evaluation [888]. reabsorbing [641]. reacting [1059]. reaction [489, 550, 1124, 940, 653, 933, 1085]. reactions [489, 523, 843, 844, 924, 1106, 655, 1007, 326, 414, 653, 825, 1084, 405, 412, 433, 439, 440, 210, 521]. read [719]. readers [749]. realizability [635]. realization [636]. realizations [1102]. reappraisal [816]. receptor [1037]. receptors [164]. reciprocity [490]. recirculating [921]. recombination [234, 271]. recombinations [688]. Reconstruction [814]. recovery [149]. rectification [1129, 223]. reduce [1009]. reducing [365]. reduction [655, 1022]. reference [229, 326, 106, 84, 396, 404, 643, 124, 5, 67, 432, 832, 1100]. Reflections [305, 324]. reflex [248]. Reflexes [675, 675]. refractory [241]. regeneration [570]. region [412, 440]. Regional [748, 568, 515, 1155]. regression [1111]. regulation [722, 990, 22, 27]. reinforcement [1027, 43]. reinterpretation [147]. relapsing [363]. related [316, 797, 852, 1142, 840]. relating [365]. Relation [1094, 369, 572, 162, 735, 128, 635, 755, 132]. Relational [1103, 714, 534, 627, 1003, 1019, 1069, 870, 494, 507, 575, 603, 661]. relations [425, 375, 246, 269, 327, 407, 583, 817, 799, 894, 549, 656]. Relationship [552, 1043, 962, 161, 450]. relationships [938]. relative [1099, 732, 928, 967]. relativity [1031]. Relaxation [1037]. relevant [794]. Remark [468, 755, 942, 1039, 1150]. Remarks [955, 408, 511, 519, 554, 696, 684, 1031, 1134, 1049, 42]. removal [118, 237, 244, 643]. renal [453, 752]. renewal [706]. repetitive [1133]. replication [804, 842]. represent [245]. Representation [492, 550, 280, 786, 516, 688, 501, 615]. representations [979, 475, 491, 586]. represented [403]. reproducing [809, 525, 619]. reproduction [946, 497, 1142]. request [579]. Research [790, 506, 835, 456, 1073]. reserpine [1027]. reservoir [530, 612, 743]. resistance [741, 988]. resolution [809]. resonance [379]. Resonances [259, 446, 376, 392, 447, 448, 450]. resources [659]. Respiration [1048, 278, 362, 1]. respiratory [515, 556, 559, 237, 244, 643, 903, 1153]. Response [306, 53, 60, 161, 194, 685, 794, 1096, 1026, 485, 944]. responses [49, 813, 256]. rest [343]. Restricted [818]. resulting [855, 50]. results [862, 661]. retaining [453]. retinal [610]. Retrieval [630]. return [82]. reversibility [805, 644, 703]. reversible [953]. reversibly [1059]. Review [710, 709]. reviews [63]. Reward [832]. rhythm [761, 859]. riboflavin [564]. ribonucleic [1001]. Riccati [1135]. ring [812]. Riots [952, 988]. risks [276]. road [712]. Roberge [954]. Roginsky [1147]. role [84, 643, 242, 544, 869, 595]. root [1008]. roots [581]. Rosen [809, 696]. rotations [858]. rote [535]. rubber [936]. rumor [308]. Rushton [26].

S. [723]. saline [1091]. saltatory [213]. sampling [995]. sarcomere [1094]. satisfaction [176, 181]. saturation [124]. schedule [1027]. schedules [1022, 1026]. schizophrenia [800]. schizophrenias [695, 740]. scholarly [75]. Science [527, 663, 700, 651]. sciences [662, 411, 458]. Scientific [579]. Scientist [457]. scientists [676, 699]. scopolamine [907]. score [327]. se [1045]. search [387, 397, 481]. secondary [128]. secreted [779]. secretion [1004, 916]. secretory [779]. section [392, 448]. security [457]. segregation [771]. selection [970]. Self [619, 1142, 711, 608, 991, 809, 658, 946, 22, 27, 525, 812]. Self-absorption [711]. self-fertilized [812]. self-organizing [608, 658]. self-preserving [991]. self-regulation [22, 27]. Self-reproducing [619, 809, 525]. Self-reproduction [1142, 946]. semicircular [858]. semiconduction [1029]. sense [1105]. sensitive [270]. sensitivity [692, 160]. sensoriels [791]. sensory [791, 1150]. separate [522]. separated [24]. septal [807]. sequence [1014]. sequences [814, 51]. sequential [805, 923, 565, 690, 703, 803]. Sergeevich [763]. serial [1142]. Service [1040]. services [1040]. Session [748, 568]. set [550, 1105, 934, 754, 496, 505]. set-theoretical [496, 505]. sets [1105, 797, 616, 583, 817, 799, 868, 845, 894, 943, 927, 1069, 1053, 1060, 1039, 1081, 1082, 1159, 1150, 656]. several [475]. Sex [1159]. sexes [1138, 1053]. shape [1041, 1107, 712, 1000, 1075, 811]. shaped [33, 39]. shapes [84, 85, 150]. shear [1123]. sheath [415]. shell [33, 39]. shell-shaped [39]. Short [631]. showing [998]. shunting [807]. sickness [164]. sides [331]. sigma [449]. signed [906]. significance [1041, 1000, 1075]. similar [1003, 1132]. similarities [796]. similarity [138, 732]. Simple [739, 425, 54, 59, 686, 689, 256, 62, 762, 80, 284, 225]. Simplification [917]. Simplified [119, 975]. simulated [1051]. simulation [1062, 991, 1152, 830]. Simultaneous [401, 1145, 405, 412, 433, 439, 440, 261]. single [876, 692, 745, 997, 233, 293, 1145, 1120, 1068, 858]. single-celled [997]. single-compartment [1120]. singularity [475]. sinus [758]. sites [683, 693, 707, 717, 728, 833]. situation [687, 686, 689]. situations [302, 597]. size [107, 757]. skeletal [129]. sketch [718]. skin [53]. sleep [705]. slightly [1096]. small [655, 389]. smallest [307]. smooth [731]. social [302, 912, 389, 269, 930, 484, 219, 224, 263, 298, 544, 597, 601, 845, 952, 988, 411]. societies [246, 269, 327, 1159]. Society [790, 735, 560, 953]. socio [358, 359, 368]. socio-structural [358, 359, 368]. sociology [227, 232, 168, 387, 397, 667, 816, 895, 986, 1031]. sodium [453, 697]. sol [729]. sol-gel [729]. solar [1077]. solid [730, 747]. solubility [828, 829, 830]. solute [1033]. solution [1083, 844, 582, 313, 325, 335, 940, 743, 316, 929, 723]. Solutions [826, 1051, 1146, 647, 395]. solving [405, 439]. Some [1109, 542, 56, 89, 665, 186, 320, 243, 365, 551, 305, 330, 324, 1030, 653, 1126, 1160, 1009, 793, 808, 130, 398, 30, 41, 80, 406, 420, 40, 88, 158, 226, 301, 345, 408, 402, 426, 511, 519, 554, 696, 740, 941, 987, 1019, 1031, 1134, 1049, 518, 543, 661, 720, 1102, 283, 42, 499, 37, 395, 715, 731, 751, 612, 57, 302, 214, 217, 238, 239, 245, 297, 434, 819, 818, 62, 405, 439, 32, 144, 148, 167, 254, 454, 627, 800, 973, 1122, 680, 773, 184, 215, 516, 136, 642, 269, 295, 927]. Some [803, 829]. sommeil [705]. sources [242]. Southern [748, 568]. space [459, 1047, 1013, 902, 987, 522]. spaced [574]. spaces [522]. span [480]. spastic [83]. spatial [1028, 280]. spatially [317]. Special [49, 326, 124, 67, 1100]. specialists [631]. species [461, 404, 934, 867]. specific [779, 939, 728, 833, 1051, 1009]. specification [984]. specificity [1021]. Spects [361]. speech [909, 949]. sphere [1125, 31]. spheres [711]. spherical [186, 1078, 757, 586]. spheroids [1006]. spindle [1080, 369]. spontaneous [319, 1133, 486]. sporadic [1121]. Sportlern [893]. sporulation [106]. Spread [374, 358, 359, 368, 866, 389, 347, 332, 308, 332, 374]. Stability [923, 853, 955, 954, 1090, 163]. stage [509]. stages [791]. staining [433]. State [640, 748, 455, 1139, 370, 730, 552, 399, 214, 217, 238, 239, 48, 54, 59, 65, 389, 743, 933, 500, 1108, 969]. states [207, 208, 294, 437]. static [87]. stationarity [897]. Statistical [506, 58, 461, 514, 104, 348, 256, 188, 341, 1073]. Statistics [748, 568, 287, 310, 388, 477, 417]. Steady [743, 933, 207, 208, 455, 1139, 370, 552, 399, 214, 217, 238, 239, 48, 54, 59, 65, 500, 279, 1108, 294, 969]. Steady-state [743, 1139, 399, 48, 54, 59, 65, 969]. Stear [1004]. Steigen [577]. stenosis [884]. Stevens [910]. stimulation [708, 223, 415, 521]. stimuli [24, 517, 51, 654, 755, 43]. Stimulus [1096, 194, 241]. Stimulus-response [1096]. Stochastic [489, 523, 1088, 1152, 1085, 513, 1083, 674, 685, 701, 1084, 346, 706, 1120, 668, 671]. Stoke [31]. stop [742]. storing [540, 604]. straight [1045, 712]. strains [532]. Strandedness [967, 1032]. stratagem [814]. strategic [893]. Strategischen [893]. stream [995, 118]. Streaming [16]. strength [120, 125, 128]. stress [958, 1002, 1144]. stressed [936]. stresses [532, 318, 1046, 193]. striated [129]. stroma [1091]. Strong [703]. Structural [1091, 356, 358, 359, 368, 575]. Structure [1136, 384, 996, 425, 959, 369, 1149, 396, 404, 246, 269, 327, 735, 930, 957, 158, 215, 299, 947, 1001]. Structures [375, 783, 851, 1103, 852, 642]. struggle [867, 935, 1156]. studied [723]. Studies [384, 12, 572, 44, 49, 101, 407, 674, 685, 110, 444, 386, 722, 556, 559, 995, 948, 1108, 794, 800]. study [287, 310, 388, 477, 1112, 191, 1107, 353, 508, 604, 791]. sub [194, 68]. sub-threshold [194, 68]. subject [674]. submissiveness [1081]. subnormal [337, 336]. substance [1059, 1051, 350]. substances [229, 1155]. substrate [933]. subsystems [1141]. successifs [1040]. successive [1040]. successively [735]. succinoxidase [314]. Suggested [181, 265, 368]. Suggestion [520, 257, 913, 973, 895]. Suggestions [67, 167, 254, 1133, 226, 301]. sum [435]. Summer [748, 568]. supercategores [1072]. supercategories [963, 1104, 979]. support [441]. supports [156, 159]. surface [961, 84, 163, 475]. surfaces [747, 1147]. surrounding [1078, 1125]. survey [456]. survival [471, 905, 974, 911]. suspended [985]. sustained [1030]. swelling [1091, 318]. symbiosis [435, 420]. symbols [437]. symmetry [1036]. sympathomimetic [677]. symposia [662]. Symposium [558, 569, 802, 1089, 361, 1076, 1005]. synapse [65]. synchronization [826]. synchronized [1067]. synchronous [106]. synthesis [565, 587, 895]. Syracuse [764]. System [992, 141, 367, 409, 530, 612, 726, 916, 990, 905, 746, 1124, 939, 565, 587, 305, 333, 324, 379, 10, 643, 1051, 708, 880, 1034, 883, 859, 9, 147, 154, 646, 670, 1121, 1134, 1141, 567, 810, 706, 188, 1085, 1120, 541, 485, 492, 694]. Systèmes [1040]. systemic [530, 612]. Systems [1008, 361, 513, 455, 608, 1004, 963, 1072, 1104, 370, 384, 165, 174, 367, 409, 416, 979, 451, 805, 922, 923, 916, 1005, 1007, 921, 399, 531, 746, 815, 861, 214, 217, 238, 239, 326, 312, 340, 920, 343, 372, 637, 66, 377, 403, 658, 782, 834, 891, 892, 836, 1146, 1054, 1009, 701, 398, 1108, 405, 412, 439, 440, 576, 1096, 897, 1040, 294, 201, 753, 680, 773, 494, 507, 595, 575, 644, 690, 703, 716, 803, 804, 842, 950, 951, 1102, 516, 724, 163, 199, 225, 529, 528, 539, 1136, 784, 58, 339, 486, 944]. systems-mathematical [920]. systole [611]. systolic [612].

taking [741]. Target [453, 666, 672]. Target-organ [453]. targets [262, 464]. task [383, 535]. task-oriented [383]. Temperature [1080, 969, 1061, 1077]. Temporal [788, 396, 404, 1087, 51]. temporally [24]. temporelles [1087]. tension [1016, 129, 84, 163, 847, 873]. Tensor [996, 562]. termolecular [1083]. terms [1009, 786]. tertiary [1001]. test [194]. tests [1047, 1040]. their [818, 128, 149, 166, 695, 607, 661, 1102]. them [1103]. theorem [756, 423]. Theorems [734, 267, 402, 37]. theoretic [1021, 546, 657]. Theoretical [478, 706, 89, 924, 365, 625, 245, 934, 948, 406, 420, 496, 505, 941, 770, 199, 1157]. Theorie [479, 918]. Theories [558, 569, 115, 252, 26]. Theory [1119, 1012, 958, 796, 1127, 221, 60, 428, 374, 182, 90, 197, 352, 364, 529, 528, 539, 376, 392, 447, 448, 450, 651, 963, 550, 1105, 666, 672, 117, 151, 172, 798, 562, 925, 530, 612, 655, 766, 747, 900, 889, 363, 862, 726, 1028, 673, 216, 228, 649, 760, 1091, 390, 143, 920, 6, 12, 48, 54, 59, 65, 107, 126, 1014, 369, 479, 66, 118, 332, 393, 918, 105, 653, 295, 360, 633, 841, 674, 685, 124, 204, 30, 41, 166, 242, 285, 223, 442, 469, 576, 36, 61, 87, 91, 806]. theory [169, 171, 241, 293, 308, 391, 418, 443, 473, 28, 27, 67, 99, 110, 177, 236, 263, 301, 382, 410, 460, 488, 511, 553, 571, 578, 601, 869, 845, 913, 972, 973, 1018, 1002, 1053, 1060, 1081, 1082, 274, 494, 507, 575, 623, 639, 661, 716, 136, 299, 557, 163, 180, 188, 737, 122, 123, 140, 739, 76, 113, 1120, 668, 671, 660, 822, 610, 1092, 34, 46, 68, 338, 321, 73, 77, 499, 909, 949, 828, 829, 830, 489]. therapeutic [162]. there [1053]. thermal [1146, 1054, 1055, 694]. thermodynamic [900, 238, 239]. thermodynamics [370]. thesis [635]. thick [775]. thick-walled [775]. thin [186]. thin-walled [186]. thinking [145, 152, 411]. third [267]. Thompson [623]. Thoroughly [374, 332]. those [572, 412, 440]. Three [969, 987, 336, 935]. Three-dimensional [969, 336]. three-dimensionality [987]. threshold [253, 413, 268, 194, 1126, 1160, 756, 306, 998, 68]. thrombosis [1002]. thymidine [838]. thyroid [1004, 419, 916, 990]. thyroid-pituitary [419]. thyroxin [754]. Time [210, 425, 981, 161, 347, 306, 480, 354, 596]. time-varying [981]. times [60, 383, 1118]. Tissue [220, 677, 56, 89, 320, 1028, 1101, 761, 1097, 1098, 919, 976, 1048, 221, 434, 124, 204, 299, 199, 122, 123, 140]. tissues [134, 5, 480, 352, 364]. tones [74, 67]. tonometry [948]. Topological [414, 452, 394, 408, 426, 431, 468, 496, 504, 680, 423]. Topology [387, 397, 57, 993, 139, 410, 402]. total [535]. traccianti [629]. traced [1051]. tracer [455, 399, 531, 781, 815, 861, 939, 1058, 377, 1009, 469]. tracers [629]. tracking [712]. tract [515, 237, 244, 643, 180]. traineeship [580, 609]. traineeships [650]. Training [676, 699, 640]. Traitement [1040]. Transfer [1118, 899, 782, 834, 941, 1142, 1001]. transfer-rate [782, 834]. transform [1023, 498]. transformation [729, 452]. transformations [426, 623]. transfusions [739]. Transient [366, 1054, 659, 944, 509, 1146]. transition [818]. transitions [393, 395]. transitivity [358, 359]. Transmission [761, 311, 349, 213]. Transport [582, 229, 357, 655, 766, 747, 1013, 1080, 1147, 1101, 768, 537, 1058, 1137, 1140, 412, 440, 442, 469, 576, 1018, 951, 493, 549, 529, 528, 539]. transverse [400]. trapping [837]. treatment [861, 331, 484, 337, 695, 336, 1040]. Tree [156, 159, 856, 1149]. trees [120, 125, 128, 153]. tri [86]. tri-axial [86]. trials [49, 574]. triangle [962]. Trucco [1066]. truncated [1058]. trunk [156, 120]. trunks [156, 159, 380, 355]. tryptophan [572]. tube [792, 351, 371, 985, 994, 860, 279, 132, 775]. tubes [966, 328, 936]. tubule [752]. tumor [1022]. tunneling [874]. Turing [951]. Two [684, 1090, 1142, 1024, 871, 170, 974, 126, 331, 378, 1138, 782, 834, 689, 256, 1045, 223, 176, 169, 171, 267, 168, 298, 290, 654, 1053, 951, 611, 72, 288, 522, 867]. two-chambered [611]. two-factor [126, 223, 951]. Tyler [723]. type [1021, 980, 692, 581, 637, 1017, 469, 431].

ultracentrifuge [929]. ultrafiltration [545]. uncertainty [535]. undamped [1100]. underlie [621, 636]. unendlich [638]. unequal [522]. unicellular [938]. unidirectional [1137]. unified [986]. uniform [614, 85, 873, 847]. unimolecular [489, 523]. Union [579]. unit [1064]. unitaires [1087]. unitary [963, 1087]. units [203]. universals [178]. University [287, 310, 388, 748, 764]. unreliability [597]. uphill [617]. upon [194, 454]. uptake [204]. ureter [1051]. ureters [857]. urethra [780, 840, 873, 847]. urine [841]. use [676, 699, 592, 474, 782, 834, 948, 942, 498, 723, 499]. used [502]. uses [651]. using [530, 612, 862, 692, 1014]. uterine [322]. utilization [555].

V [781, 932, 1044]. vacuolar [900]. vacuum [352]. Validity [204]. value [307]. values [139]. variability [160]. variable [210, 1026, 163]. variance [337, 336]. Variation [611]. variational [795, 667]. variations [741]. various [428, 938, 359]. varying [981, 557]. vascular [367, 409, 1149, 1155]. vascularization [869]. velocities [860, 936]. velocity [371, 213]. venous [777]. ventilation [522]. ventricle [1041, 1006, 1093, 821, 882, 1052, 1000, 1035, 1065, 1075, 1144, 1074]. Ventricular [807, 982, 1043, 981, 1046, 879, 822]. Verhaltens [893]. verifiable [431]. vessel [350]. vessels [177]. VI [204]. vibrations [177]. Victor [763]. view [1150, 113]. VII [815]. VIII [861]. violent [988]. Virginia [568]. virus [757]. visco [715, 741, 751]. visco-elastic [715, 741, 751]. viscoelastic [1023]. viscosity [376, 447]. viscous [792, 351, 371, 279, 132, 775, 4, 20, 37]. viscus [186]. vision [183, 304, 428, 526, 391, 418]. visual [899, 462, 178, 79, 88, 121, 127, 133, 432, 632, 652, 765]. Visually [38]. vitro} [746]. vivants [1040]. vivo [731, 1043, 1155]. voltage [1119, 844, 1012]. voltage-clamp [1119]. Volterra [573]. volume [982, 1043, 502, 270, 1067, 491, 586]. volumes [872, 1096]. Voukydi [1041].

Wabenzelle [648]. Wahrscheinlichkeit [638]. walk [83, 344]. walking [617]. Wall [1123, 371, 878]. walled [186, 775]. walls [641, 177]. washout [1047, 1036, 1145]. wasps [484]. water [151, 453, 889, 1013, 1033, 203, 841, 119]. water-flow [151]. Wave [1113, 392, 448]. waves [783, 1013, 305, 324, 371, 1100]. wealth [231]. Weber [46]. weights [21]. Weissenberger [955]. Wenckebach [1079]. Wettläufen [893]. which [1029, 779, 1009, 142, 405, 412, 439, 440, 168, 1122, 307]. White [789]. who [866]. whole [556, 559]. wide [1105]. will [378, 288]. Windkessel [769]. winds [153]. Wirkung [479]. within [744, 1051, 883]. Without [1097, 628]. word [638]. world [835]. Wortes [638]. Wortlänge [638]. Wurmserg [633].

X [888]. X-ray [888]. XIV [662].

yield [577].

Zeldovich [1147]. zero [637]. zur [479].

References

Landahl:1939:MBC

[1] H. D. Landahl. Mathematical biophysics of cell respiration II. Bulletin of Mathematical Biophysics, 1(1):1–17, March 1939. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

      link.springer.com/article/10.1007/BF02478008.

Weinberg:1939:NBP

[2] Alvin M. Weinberg. Note on biological periodicities. Bulletin of Mathematical Biophysics, 1(1):19–22, March 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

      com/article/10.1007/BF02478009.

Rashevsky:1939:MCDa

[3] N. Rashevsky. The mechanism of cell division. Bulletin of Mathematical Biophysics, 1(1):23–30, March 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

      article/10.1007/BF02478010.

Young:1939:MVBa

[4] Gale Young. On the mechanics of viscous bodies and elongation in ellipsoidal cells. Bulletin of Mathematical Biophysics, 1(1):31–46, March 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478011.

Rashevsky:1939:MBC

[5] N. Rashevsky. Mathematical biophysics of the cell with reference to the contractility of tissues and amoeboid movements. Bulletin of Mathematical Biophysics, 1(1):47–62, March 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478012.

Householder:1939:CRT

[6] A. S. Householder. Concerning Rashevsky’s theory of the “Gestalt”. Bulletin of Mathematical Biophysics, 1(1):63–73, March 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478013.

Rashevsky:1939:MCDb

[7] N. Rashevsky. The mechanism of cell division. Bulletin of Mathematical Biophysics, 1(2):ii, June 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478105;   http://link.springer.com/
content/pdf/10.1007/BF02478105.pdf.

Young:1939:ECCa

[8] Gale Young. Elongation and constriction in cell division. Bulletin of Mathematical Biophysics, 1(2):75–91, June 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478106.

Rashevsky:1939:CPC

[9] N. Rashevsky. On the so called plasticity of the central nervous system. Bulletin of Mathematical Biophysics, 1(2):93–94, June 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478107.

Landahl:1939:CMBa

[10] H. D. Landahl. Contributions to the mathematical biophysics of the central nervous system. Bulletin of Mathematical Biophysics, 1(2): 95–118, June 1939. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478108.

Rashevsky:1939:MBGa

[11] N. Rashevsky. Mathematical biophysics of growth. Bulletin of Mathematical Biophysics, 1(3):119–127, September 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478180.

Householder:1939:SMT

[12] A. S. Householder. Studies in the mathematical theory of excitation. Bulletin of Mathematical Biophysics, 1(3):129–141, September 1939. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478181.

Reiner:1939:DCM

[13] John M. Reiner. Diffusion in colloidal media. Bulletin of Mathematical Biophysics, 1(3):143–149, September 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478182.

Williamson:1939:CED

[14] Robert R. Williamson. A comparison of electrical and diffusion forces in the metabolism of electrolytes. Bulletin of Mathematical Biophysics, 1(4):151–157, December 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02481791.

Landahl:1939:CMBb

[15] H. D. Landahl. A contribution to the mathematical biophysics of psychophysical discrimination II. Bulletin of Mathematical Biophysics, 1(4): 159–176, December 1939. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481792.

Young:1939:SDF

[16] Gale Young. Streaming in diffusion fields. Bulletin of Mathematical Biophysics, 1(4):177–184, December 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481793.

Rashevsky:1939:MBGb

[17] N. Rashevsky. Mathematical biophysics of growth. Bulletin of Mathematical Biophysics, 1(4):185, December 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481797;   http://link.springer.com/
content/pdf/10.1007/BF02481797.pdf.

Rashevsky:1939:MCDc

[18] N. Rashevsky. The mechanism of cell division. Bulletin of Mathematical Biophysics, 1(4):185, December 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481794;   http://link.springer.com/
content/pdf/10.1007/BF02481794.pdf.

Young:1939:ECCb

[19] Gale Young. Elongation and constriction in cell division. Bulletin of Mathematical Biophysics, 1(4):185, December 1939. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02481796;
http://link.springer.com/content/pdf/10.1007/BF02481796.pdf.

Young:1939:MVBb

[20] Gale Young. On the mechanics of viscous bodies and elongation in ellipsoidal cells. Bulletin of Mathematical Biophysics, 1(4):185, December 1939. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02481795;
http://link.springer.com/content/pdf/10.1007/BF02481795.pdf.

Householder:1940:NMDa

[21] A. S. Householder. A neural mechanism for discrimination: II. Discrimination of weights. Bulletin of Mathematical Biophysics, 2(1):1–13, March 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478027.

Rashevsky:1940:AMB

[22] N. Rashevsky. An approach to the mathematical biophysics of biological self-regulation and of cell polarity. Bulletin of Mathematical Biophysics, 2(1):15–25, March 1940. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478028.

Rashevsky:1940:CMBa

[23] N. Rashevsky. Contributions to the mathematical biophysics of organic form I. Formation of cavities in cellular aggregates. Bulletin of Mathematical Biophysics, 2(1):27–36, March 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478029.

Landahl:1940:DBT

[24] H. D. Landahl. Discrimination between temporally separated stimuli. Bulletin of Mathematical Biophysics, 2(1):37–47, March 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478030.

Young:1940:CDP

[25] Gale Young. Convective diffusion in parallel flow fields. Bulletin of Mathematical Biophysics, 2(1):49–59, March 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478031.

Weinberg:1940:ECT

[26] Alvin M. Weinberg. The equivalence of the conduction theories of Rashevsky and Rushton. Bulletin of Mathematical Biophysics, 2(2): 61–64, June 1940. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478169.

Rashevsky:1940:FCT

[27] N. Rashevsky. Further contributions to the theory of cell polarity and self-regulation. Bulletin of Mathematical Biophysics, 2(2):65–67, June 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478170.

Rashevsky:1940:CMT

[28] N. Rashevsky. Contributions to the mathematical theory of organic form. II asymmetric metabolism of cellular aggregates. Bulletin of Mathematical Biophysics, 2(2):69–72, June 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478171.

Landahl:1940:CMB

[29] H. D. Landahl. A contribution to the mathematical biophysics of psychophysical discrimination III. Bulletin of Mathematical Biophysics, 2 (2):73–87, June 1940. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478172.

Offner:1940:NCTa

[30] Franklin Offner, Alvin Weinberg, and Gale Young. Nerve conduction theory: Some mathematical consequences of Bernstein’s model. Bulletin of Mathematical Biophysics, 2(2):89–103, June 1940. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478173.

Young:1940:GCE

[31] Gale Young. A generalization of Cunningham’s extension of Stoke’s law for the force on a sphere. Bulletin of Mathematical Biophysics, 2(3): 105–108, September 1940. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478174.

Rashevsky:1940:PAS

[32] N. Rashevsky. Physicomathematical aspects of some problems of organic form. Bulletin of Mathematical Biophysics, 2(3):109–121, September 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478175.

Rashevsky:1940:CMBb

[33] N. Rashevsky. Contributions to the mathematical biophysics of organic form III. Deformation of shell shaped cellular aggregates. Bulletin of Mathematical Biophysics, 2(3):123–126, September 1940. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478176.

Weinberg:1940:FTN

[34] Alvin M. Weinberg. On the formal theory of nerve conduction. Bulletin of Mathematical Biophysics, 2(3):127–133, September 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478177.

Householder:1940:NH

[35] Alston S. Householder. A note on the horopter. Bulletin of Mathematical Biophysics, 2(3):135–140, September 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478178.

Peters:1940:GFC

[36] H. C. Peters. The general fluid circuit theory of active chloride absorption. Bulletin of Mathematical Biophysics, 2(3):141–143, September 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478179.

Young:1940:SGT

[37] Gale Young. Some general theorems on the motion of incompressible viscous fluids. Bulletin of Mathematical Biophysics, 2(4):145–155, December 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478266.

Householder:1940:NMDb

[38] Alston S. Householder. A neural mechanism for discrimination III: Visually perceived lengths and distances. Bulletin of Mathematical Biophysics, 2(4):157–167, December 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478267.

Rashevsky:1940:DSS

[39] N. Rashevsky. Deformation of shell-shaped cellular aggregates: Application to gastrulation. Bulletin of Mathematical Biophysics, 2(4): 169–175, December 1940. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478268.

Rashevsky:1940:SCM

[40] N. Rashevsky. Some considerations on mathematical molecular biophysics. Bulletin of Mathematical Biophysics, 2(4):177–181, December 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478269.

Offner:1940:NCTb

[41] F. Offner, A. Weinberg, and G. Young. Nerve conduction theory: Some mathematical consequences of Bernstein’s model. Bulletin of Mathematical Biophysics, 2(4):183, December 1940. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478270;   http://link.springer.com/
content/pdf/10.1007/BF02478270.pdf.

Vsky:1941:SRM

[42] N. Rash. Vsky. Some remarks on the movement of chromosomes during cell division. Bulletin of Mathematical Biophysics, 3(1):1–3, March 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478101.

Young:1941:RIB

[43] Gale Young. On reinforcement and interference between stimuli. Bulletin of Mathematical Biophysics, 3(1):5–12, March 1941. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478102.

Landahl:1941:SMBa

[44] H. D. Landahl. Studies in the mathematical biophysics of discrimination and conditioning. I. Bulletin of Mathematical Biophysics, 3 (1):13–26, March 1941. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478103.

Householder:1941:MBC

[45] Alston S. Householder. Mathematical biophysics of cellular forms and movements. Bulletin of Mathematical Biophysics, 3(1):27–38, March 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478104.

Weinberg:1941:WTK

[46] Alvin M. Weinberg. Weber’s theory of the Kernleiter. Bulletin of Mathematical Biophysics, 3(2):39–55, June 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478218.

Rashevsky:1941:DCC

[47] N. Rashevsky. The dynamics of cell constriction during division. Bulletin of Mathematical Biophysics, 3(2):57–62, June 1941. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478219.

Householder:1941:TSSa

[48] Alston S. Householder. A theory of steady-state activity in nerve-fiber networks: I. Definitions and preliminary lemmas. Bulletin of Mathematical Biophysics, 3(2):63–69, June 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478220.

Landahl:1941:SMBb

[49] H. D. Landahl. Studies in the mathematical biophysics of discrimination and conditioning II: Special case: Errors, trials, and number of possible responses. Bulletin of Mathematical Biophysics, 3(2):71–77, June 1941.

CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478221.

Williamson:1941:ECP

[50] Robert R. Williamson. Electrical charges and potentials in cells resulting from metabolism of electrolytes. Bulletin of Mathematical Biophysics, 3 (3):79–87, September 1941. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478164.

Rashevsky:1941:NMB

[51] N. Rashevsky. Note on the mathematical biophysics of temporal sequences of stimuli. Bulletin of Mathematical Biophysics, 3(3):89– 92, September 1941. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478165.

Rashevsky:1941:NNC

[52] N. Rashevsky. A note on the nature of correlations between different characteristics of organisms. Bulletin of Mathematical Biophysics, 3(3): 93–95, September 1941. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478166.

Coombs:1941:MBG

[53] Clyde H. Coombs. Mathematical biophysics of the Galvanic skin response. Bulletin of Mathematical Biophysics, 3(3):97–103, September 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478167.

Householder:1941:TSSb

[54] Alston S. Householder. A theory of steady-state activity in nervefiber networks II: The simple circuit. Bulletin of Mathematical Biophysics, 3(3):105–112, September 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478168.

Stanton:1941:NMD

[55] Henry E. Stanton. A neural mechanism for discrimination IV: Monocular depth perception. Bulletin of Mathematical Biophysics, 3(4):113– 120, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477929.

Bloch:1941:SPC

[56] Ingram Bloch. Some preliminary considerations concerning concentration of oxygen in tissue. Bulletin of Mathematical Biophysics, 3(4): 121–126, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477930.

Gause:1941:NPA

[57] G. F. Gause. Note on a possible application of some concepts of topology to asymmetric organization of protoplasm. Bulletin of Mathematical Biophysics, 3(4):127–128, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477931.

Weinberg:1941:SDI

[58] A. M. Weinberg and A. S. Householder. Statistical distribution of impedance elements in biological systems. Bulletin of Mathematical Biophysics, 3(4):129–135, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477932.

Householder:1941:TSSc

[59] Alston S. Householder. A theory of steady-state activity in nerve-fiber networks III: The simple circuit in complete activity. Bulletin of Mathematical Biophysics, 3(4):137–140, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477933.

Landahl:1941:TDR

[60] H. D. Landahl. Theory of the distribution of response times in nerve fibers. Bulletin of Mathematical Biophysics, 3(4):141–147, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477934.

Peters:1941:AGF

[61] H. C. Peters. Application of the general fluid circuit theory to the chloride-bromide-deuterium oxide experiment. Bulletin of Mathematical Biophysics, 3(4):149–152, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477935.

Morales:1941:GME

[62] Manuel F. Morales and Nathan W. Shock. The general membrane equilibrium equation: Its simple derivation and some of its biological implications. Bulletin of Mathematical Biophysics, 3(4):153–160, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477936.

Weinberg:1941:BR

[63] Alvin M. Weinberg. Book reviews. Bulletin of Mathematical Biophysics, 3(4):161–162, December 1941. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477937.

Branson:1942:DFA

[64] Herman Branson. Diffusion as a function of aggregation in colloidal media. Bulletin of Mathematical Biophysics, 4(1):1–6, March 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477349.

Householder:1942:TSS

[65] Alston S. Householder. A theory of steady-state activity in nerve-fiber networks: IV. N circuits with a common synapse. Bulletin of Mathematical Biophysics, 4(1):7–14, March 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477350.

Landahl:1942:KTD

[66] H. D. Landahl. A kinetic theory of diffusion forces in metabolizing systems. Bulletin of Mathematical Biophysics, 4(1):15–26, March 1942. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477351.

Rashevsky:1942:SMB

[67] N. Rashevsky. Suggestions for a mathematical biophysics of auditory perception with special reference to the theory of aesthetic ratings of combinations of musical tones. Bulletin of Mathematical Biophysics, 4 (1):27–32, March 1942. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477352.

Weinberg:1942:NLE

[68] Alvin M. Weinberg. Non-linear excitation theory: Non-accommodative, sub-threshold effects. Bulletin of Mathematical Biophysics, 4(1):33– 44, March 1942. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477353.

Landahl:1942:MAE

[69] H. D. Landahl. A mathematical analysis of elongation and constriction in cell division. Bulletin of Mathematical Biophysics, 4(2):45–62, June 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477309.

Shock:1942:FFD

[70] Nathan W. Shock and Manuel F. Morales. A fundamental form for the differential equation of colonial and organism growth. Bulletin of Mathematical Biophysics, 4(2):63–71, June 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477310.

Householder:1942:NDE

[71] Alston S. Householder and Robert R. Williamson. A note on the diffusion of electrolytes in cells. Bulletin of Mathematical Biophysics, 4 (2):73–76, June 1942. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477311.

Sacher:1942:PPI

[72] George Sacher. Periodic phenomena in the interaction of two neurons. Bulletin of Mathematical Biophysics, 4(2):77–81, June 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477312.

Williamson:1942:TEPa

[73] Robert R. Williamson and Ingram Bloch. A theory of electrical polarity in cells. Bulletin of Mathematical Biophysics, 4(2):83–88, June 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477313.

Rashevsky:1942:AAM

[74] N. Rashevsky. An alternate approach to the mathematical biophysics of perception of combinations of musical tones. Bulletin of Mathematical Biophysics, 4(2):89–90, June 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477314.

Anonymous:1942:CSJ

[75] Anonymous. Conservation of scholarly journals an announcement. Bulletin of Mathematical Biophysics, 4(2):91, June 1942. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477315.

Stanton:1942:MTP

[76] Henry E. Stanton. A mathematical theory of protoplasmic protrusions. I. Bulletin of Mathematical Biophysics, 4(3):93–100, September 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477938.

Williamson:1942:TEPb

[77] Robert R. Williamson. A theory of electrical polarity in cells: II. Bulletin of Mathematical Biophysics, 4(3):101–105, September 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477939.

Weinberg:1942:GFB

[78] Alvin M. Weinberg. Green’s functions in biological potential problems. Bulletin of Mathematical Biophysics, 4(3):107–115, September 1942. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477940.

Rashevsky:1942:FCM

[79] N. Rashevsky. Further contributions to the mathematical biophysics of visual aesthetics. Bulletin of Mathematical Biophysics, 4(3):117– 120, September 1942. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477941.

Pitts:1942:SOS

[80] Walter Pitts. Some observations on the simple neuron circuit. Bulletin of Mathematical Biophysics, 4(3):121–129, September 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477942.

Branson:1942:DAD

[81] Herman Branson. Delayed adsorption and diffusion in colloidal media. Bulletin of Mathematical Biophysics, 4(3):131–137, September 1942. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477943.

Landahl:1942:ERR

[82] H. D. Landahl. An expression for the rate of return of an egg after artificial deformation. Bulletin of Mathematical Biophysics, 4(3): 139–147, September 1942. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477944.

Hirsh:1942:ESW

[83] Frederick R. Hirsh, Jr. An explanation of spastic walk due to Encephalitis lethargica. Bulletin of Mathematical Biophysics, 4(3):149– 150, September 1942. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477945.

Landahl:1942:ASC

[84] H. D. Landahl. An analysis of the shapes of a cell during division with particular reference to the role of surface tension. Bulletin of Mathematical Biophysics, 4(4):151–154, December 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478109.

Landahl:1942:ESN

[85] H. D. Landahl. Equilibrium shapes in non-uniform fields of concentration. Bulletin of Mathematical Biophysics, 4(4):155–158, December 1942. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478110.

Householder:1942:CFT

[86] Alston S. Householder. Cellular forms: The tri-axial cell. I. Bulletin of Mathematical Biophysics, 4(4):159–168, December 1942. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478111.

Pitts:1942:LTN

[87] Walter Pitts. The linear theory of neuron networks: The static problem. Bulletin of Mathematical Biophysics, 4(4):169–175, December 1942. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478112.

Rashevsky:1942:SPM

[88] N. Rashevsky. Some problems in mathematical biophysics of visual perception and aesthetics. Bulletin of Mathematical Biophysics, 4(4):

177–191, December 1942. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478113.

Bloch:1943:STC

[89] Ingram Bloch. Some theoretical considerations concerning the interchange of metabolites between capillaries and tissue. Bulletin of Mathematical Biophysics, 5(1):1–14, March 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478114.

Stanton:1943:TCP

[90] Henry E. Stanton. Theory of cellular pulsations. Bulletin of Mathematical Biophysics, 5(1):15–21, March 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478115.

Pitts:1943:LTN

[91] Walter Pitts. The linear theory of neuron networks: The dynamic problem. Bulletin of Mathematical Biophysics, 5(1):23–31, March 1943. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478116.

Rashevsky:1943:ONMa

[92] N. Rashevsky. Outline of a new mathematical approach to general biology: I. Bulletin of Mathematical Biophysics, 5(1):33–47, March 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478117.

Rashevsky:1943:ONMb

[93] N. Rashevsky. Outline of a new mathematical approach to general biology: II. Bulletin of Mathematical Biophysics, 5(2):49–64, June 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478014.

Rashevsky:1943:NHP

[94] N. Rashevsky. Note on the Hamiltonian Principle in biology and in physics. Bulletin of Mathematical Biophysics, 5(2):65–68, June 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478015.

Rashevsky:1943:FPA

[95] N. Rashevsky. On the form of plants and animals. Bulletin of Mathematical Biophysics, 5(2):69–73, June 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478016.

Landahl:1943:MDC

[96] H. D. Landahl. A mechanism of division of a cell with an impermeable membrane. Bulletin of Mathematical Biophysics, 5(2):75–81, June 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478017.

Householder:1943:KEI

[97] Alston S. Householder and George Gomori. The kinetics of enzyme inactivation. Bulletin of Mathematical Biophysics, 5(3):83–90, September 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478330.

Householder:1943:NAE

[98] Alston S. Householder. Note on anodal excitation. Bulletin of Mathematical Biophysics, 5(3):91–94, September 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478331.

Rashevsky:1943:TAM

[99] N. Rashevsky. On the theory of amoeboid movements. Bulletin of Mathematical Biophysics, 5(3):95–98, September 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478332.

Rashevsky:1943:MBC

[100] N. Rashevsky. Mathematical biophysics of cell division. Bulletin of Mathematical Biophysics, 5(3):99–102, September 1943. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478333.

Landahl:1943:SMB

[101] H. D. Landahl. Studies in the mathematical biophysics of discrimination and conditioning: III. Bulletin of Mathematical Biophysics, 5(3): 103–110, September 1943. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478334.

vonBonin:1943:GF

[102] Gerhardt von Bonin. On growth and form. Bulletin of Mathematical Biophysics, 5(3):111–114, September 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478335.

McCulloch:1943:LCI

[103] Warren S. McCulloch and Walter Pitts. A logical calculus of the ideas immanent in nervous activity. Bulletin of Mathematical Biophysics, 5(4): 115–133, December 1943. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478259.

Landahl:1943:SCL

[104] H. D. Landahl, W. S. McCulloch, and Walter Pitts. A statistical consequence of the logical calculus of nervous nets. Bulletin of Mathematical Biophysics, 5(4):135–137, December 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478260.

Lettvin:1943:MTA

[105] Jerome Y. Lettvin and Walter Pitts. A mathematical theory of the affective psychoses. Bulletin of Mathematical Biophysics, 5(4):139– 148, December 1943. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478261.

Householder:1943:SSP

[106] Alston S. Householder. On synchronous sporulation with possible reference to malarial parasites. Bulletin of Mathematical Biophysics, 5(4): 149–154, December 1943. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478262.

Householder:1943:TIS

[107] Alston S. Householder. A theory of the induced size effect. Bulletin of Mathematical Biophysics, 5(4):155–160, December 1943. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478263.

Householder:1943:NRE

[108] Alston S. Householder. Note on Rashevsky’s equation for cellular growth. Bulletin of Mathematical Biophysics, 5(4):161–163, December 1943. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478264.

Rashevsky:1943:OL

[109] N. Rashevsky. On the origin of life. Bulletin of Mathematical Biophysics, 5(4):165–169, December 1943. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478265.

Rashevsky:1944:SPT

[110] N. Rashevsky. Studies in the physicomathematical theory of organic form. Bulletin of Mathematical Biophysics, 6(1):1–59, March 1944. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478284.

Buchsbaum:1944:NMB

[111] Ralph Buchsbaum, N. Rashevsky, and Henry E. Stanton. A note on the mathematical biophysics of ameboid movements. Bulletin of Mathematical Biophysics, 6(2):61–63, June 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478482.

Morales:1944:PMB

[112] Manuel F. Morales. On a possible mechanism for biological periodicity. Bulletin of Mathematical Biophysics, 6(2):65–70, June 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478483.

Stanton:1944:TCD

[113] Henry E. Stanton. On the theory of cell division from the point of view of the principle of maximum energy exchange. Bulletin of Mathematical Biophysics, 6(2):71–76, June 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478484.

Landahl:1944:NNN

[114] H. D. Landahl. A note on neural nets. Bulletin of Mathematical Biophysics, 6(2):77–78, June 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478485.

Householder:1944:ETN

[115] Alston S. Householder. Equivalence of the theories of nervous excitation of Hill, monnier, and rashevsky. Bulletin of Mathematical Biophysics, 6(2):79–81, June 1944. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478486.

Householder:1944:BM

[116] Alston S. Householder. Biochemistry and morphogenesis. Bulletin of Mathematical Biophysics, 6(2):83–84, June 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478487.

Bloch:1944:TMP

[117] Ingram Bloch. A theory of membrane permeability: I. Bulletin of Mathematical Biophysics, 6(3):85–92, September 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478238.

Landahl:1944:OMT

[118] H. D. Landahl. Outline of a mathematical theory of the removal of malarial parasites from the blood stream. Bulletin of Mathematical Biophysics, 6(3):93–104, September 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478239.

Wilde:1944:SED

[119] Walter S. Wilde. Simplified equations for the distribution of chloride in body water. Bulletin of Mathematical Biophysics, 6(3):105– 111, September 1944. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478240.

Opatowski:1944:FSTa

[120] I. Opatowski. On the form and strength of trees: Part I. The trunk. Bulletin of Mathematical Biophysics, 6(3):113–118, September 1944. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478241.

Rashevsky:1944:CMBa

[121] N. Rashevsky and Virginia Brown. A contribution to the mathematical biophysics of visual perception and aesthetics. Bulletin of Mathematical Biophysics, 6(3):119–124, September 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478242.

Smith:1944:TBTa

[122] R. E. Smith, Lieut. (j.g.) and Ens M. F. Morales, H-V (S) USNR. On the theory of blood-tissue exchanges: I. Fundamental equations. Bulletin of Mathematical Biophysics, 6(3):125–131, September 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478243.

Smith:1944:TBTb

[123] R. E. Smith, Lieut. and M. F. Morales, Ens., H-V (S), USNR. On the theory of blood-tissue exchanges: II. Applications. Bulletin of Mathematical Biophysics, 6(4):133–139, December 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478430.

Morales:1944:TBT

[124] M. F. Morales, ENS. and R. E. Smith Lieut.H-V(S), USNR. On the theory of blood-tissue exchanges: III. Circulation and inert-gas exchanges at the lung with special reference to saturation. Bulletin of Mathematical Biophysics, 6(4):141–152, December 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478431.

Opatowski:1944:FSTb

[125] I. Opatowski. On the form and strength of trees: Part II. The primary branches. Bulletin of Mathematical Biophysics, 6(4):153–156, December 1944. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478432.

Householder:1944:TFT

[126] Alston S. Householder. The two-factor theory of nervous excitation with non-normal accommodation. Bulletin of Mathematical Biophysics, 6(4): 157–161, December 1944. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478433.

Rashevsky:1944:CMBb

[127] N. Rashevsky and Virginia Brown. Contributions to the mathematical biophysics of visual aesthetics. Bulletin of Mathematical Biophysics, 6(4): 163–168, December 1944. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478434.

Opatowski:1945:FST

[128] I. Opatowski. On the form and strength of trees: Part III. The secondary branches and their relation to the primary branches. Bulletin of Mathematical Biophysics, 7(1):1–4, March 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478253.

Householder:1945:MDM

[129] Alston S. Householder. Muscular dynamics and muscular efficiency: I. The isometric length-tension diagram of striated skeletal muscle. Bulletin of Mathematical Biophysics, 7(1):5–13, March 1945. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478254.

Morales:1945:SNE

[130] M. F. Morales, Ens.H-V(S) and A. S. F. L. Kreutzer, V-12(S), USNR. Some nutritional and excretional interactions and the growth of an organ or colony. Bulletin of Mathematical Biophysics, 7(1):15–24, March 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478255.

Rashevsky:1945:PMBa

[131] N. Rashevsky. A problem in the mathematical biophysics of blood circulation: I. Bulletin of Mathematical Biophysics, 7(1):25–33, March 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478256.

Rashevsky:1945:PMBb

[132] N. Rashevsky. A problem in the mathematical biophysics of blood circulation: II. Relation between pressure and flow of a viscous fluid in an elastic distensible tube. Bulletin of Mathematical Biophysics, 7 (1):35–39, March 1945. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478257.

Rashevsky:1945:CMB

[133] N. Rashevsky. A contribution to the mathematical biophysics of visual aesthetics. Bulletin of Mathematical Biophysics, 7(1):41–45, March 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478258.

Morales:1945:NPA

[134] M. F. Morales, Ens. and R. E. Smith, Lieut.H-V (S), USNR. A note on the physiological arrangement of tissues. Bulletin of Mathematical Biophysics, 7(2):47–51, June 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478452.

Householder:1945:DQL

[135] Alston S. Householder. Dynamics of quadrupedal locomotion. Bulletin of Mathematical Biophysics, 7(2):53–57, June 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478453.

Runge:1945:TPS

[136] Richard Runge. A theory of photosensitivity of some lower animals. Bulletin of Mathematical Biophysics, 7(2):59–67, June 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478454.

Rashevsky:1945:OMA

[137] N. Rashevsky. Outline of a mathematical approach to the cancer problem. Bulletin of Mathematical Biophysics, 7(2):69–81, June 1945. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478455.

Landahl:1945:NMC

[138] H. D. Landahl. Neural mechanisms for the concepts of difference and similarity. Bulletin of Mathematical Biophysics, 7(2):83–88, June 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478456.

McCulloch:1945:HVD

[139] Warren S. McCulloch. A heterarchy of values determined by the topology of nervous nets. Bulletin of Mathematical Biophysics, 7(2):89–93, June 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478457.

Smith:1945:TBT

[140] R. E. Smith and M. F. Morales. On the theory of blood-tissue exchanges: I. Fundamental equations. Bulletin of Mathematical Biophysics, 7(2):95, June 1945. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478458;   http://link.springer.com/content/pdf/10.
1007/BF02478458.pdf.

Branson:1945:MBC

[141] Herman Branson. Mathematical biophysics of the central nervous system. Bulletin of Mathematical Biophysics, 7(2):97–98, June 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478459.

Morales:1945:PFW

[142] Manuel F. Morales, Lt. (j.g.) and Robert E. Smith, Lt., H(S), USNR. The physiological factors which govern inert gas exchange. Bulletin of Mathematical Biophysics, 7(3):99–106, September 1945. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478311.

Hill:1945:CTD

[143] Terrell L. Hill and Laura E. Hill. Contribution to the theory of discrimination learning. Bulletin of Mathematical Biophysics, 7(3):107– 114, September 1945. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478312.

Rashevsky:1945:MBS

[144] N. Rashevsky. The mathematical biophysics of some mental phenomena. Bulletin of Mathematical Biophysics, 7(3):115–131, September 1945. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478313.

Rashevsky:1945:MBA

[145] N. Rashevsky. Mathematical biophysics of abstraction and logical thinking. Bulletin of Mathematical Biophysics, 7(3):133–148, September 1945. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478314.

Wilkins:1945:DDE

[146] J. Ernest Wilkins, Jr. The differential difference equation for epidemics. Bulletin of Mathematical Biophysics, 7(3):149–150, September 1945. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478315.

Rashevsky:1945:RMB

[147] N. Rashevsky. A reinterpretation of the mathematical biophysics of the central nervous system in the light of neurophysiological findings. Bulletin of Mathematical Biophysics, 7(3):151–160, September 1945. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478316.

Rashevsky:1946:MBS

[148] N. Rashevsky. The mathematical biophysics of some mental phenomena: II. Anxieties and elations. Bulletin of Mathematical Biophysics, 8(1):1–6, March 1946. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478465.

Opatowski:1946:CPTa

[149] I. Opatowski. Chain processes and their biophysical applications: Part II. The effect of recovery. Bulletin of Mathematical Biophysics, 8 (1):7–15, March 1946. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478466.

Runge:1946:CMB

[150] Richard Runge. A contribution to the mathematical biophysics of cell growth and shapes: II. Bulletin of Mathematical Biophysics, 8 (1):17–20, March 1946. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478467.

Bloch:1946:TMP

[151] Ingram Bloch. A theory of membrane permeability: II. Diffusion in the presence of water-flow. Bulletin of Mathematical Biophysics, 8 (1):21–28, March 1946. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478468.

Rashevsky:1946:NML

[152] N. Rashevsky. The neural mechanism of logical thinking. Bulletin of Mathematical Biophysics, 8(1):29–40, March 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478469.

Opatowski:1946:OGT

[153] I. Opatowski. On oblique growth of trees under the action of winds. Bulletin of Mathematical Biophysics, 8(2):41–49, June 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478460.

Rashevsky:1946:FCP

[154] N. Rashevsky. Further contributions to a probabilistic interpretation of the mathematical biophysics of the central nervous system. Bulletin of Mathematical Biophysics, 8(2):51–57, June 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478461.

Runge:1946:NEN

[155] Richard Runge. A note on the excitation of nerve pathways. Bulletin of Mathematical Biophysics, 8(2):59–63, June 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478462.

Esser:1946:TTBa

[156] Martinus H. M. Esser. Tree trunks and branches as optimum mechanical supports of the crown: I. The trunk. Bulletin of Mathematical Biophysics, 8(2):65–74, June 1946. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478463.

Landahl:1946:OMC

[157] H. D. Landahl and Richard Runge. Outline of a matrix calculus for neural nets. Bulletin of Mathematical Biophysics, 8(2):75–81, June 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478464.

Rashevsky:1946:SCS

[158] N. Rashevsky. Some considerations on the structure of quadruped extremities. Bulletin of Mathematical Biophysics, 8(3):83–93, September 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Esser:1946:TTBb

[159] Martinus H. M. Esser. Tree trunks and branches as optimum mechanical supports of the crown: II. The branches. Bulletin of Mathematical Biophysics, 8(3):95–100, September 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Opatowski:1946:PAE

[160] I. Opatowski. The probabilistic approach to the effects of radiations and variability of sensitivity. Bulletin of Mathematical Biophysics, 8(3): 101–119, September 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Landahl:1946:RBR

[161] H. D. Landahl. On the relationship between response time and dosage of a drug as a function of its mode of entry. Bulletin of Mathematical Biophysics, 8(3):121–127, September 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Landahl:1946:HLD

[162] H. D. Landahl. The half-life of a drug in relation to its therapeutic index. Bulletin of Mathematical Biophysics, 8(4):129–133, December 1946. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478447.

Shimbel:1946:CMT

[163] Alfonso Shimbel. Contributions to the mathematical theory of the stability of metabolizing systems with variable surface tension. Bulletin of Mathematical Biophysics, 8(4):135–145, December 1946. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478448.

Morales:1946:ALR

[164] Manuel F. Morales. Asynchrony of labyrinthine receptors as a physical factor in motion sickness. Bulletin of Mathematical Biophysics, 8(4): 147–157, December 1946. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478449.

Branson:1946:MDM

[165] Herman Branson. A mathematical description of metabolizing systems I. Bulletin of Mathematical Biophysics, 8(4):159–165, December 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478450.

Opatowski:1946:CPTb

[166] I. Opatowski. Chain processe and their biophysical applications: Part I. General theory. Bulletin of Mathematical Biophysics, 8(4):167, December 1946. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478451;
http://link.springer.com/content/pdf/10.1007/BF02478451.pdf.

Rashevsky:1947:SMB

[167] N. Rashevsky. Suggestions for a mathematical biophysics of some psychoses. Bulletin of Mathematical Biophysics, 9(1):1–8, March 1947. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478285.

Rashevsky:1947:PMB

[168] N. Rashevsky. A problem in the mathematical biophysics of interaction of two or more individuals which may be of interest in mathematical sociology. Bulletin of Mathematical Biophysics, 9(1):9–15, March 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478286.

Rapoport:1947:MTMa

[169] Anatol Rapoport. Mathematical theory of motivation interactions of two individuals: I. Bulletin of Mathematical Biophysics, 9(1):17–28, March 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478287.

Esser:1947:MMEa

[170] Martinus H. M. Esser. The mechanism of the middle ear: I. The two piston problem. Bulletin of Mathematical Biophysics, 9(1):29–40, March 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478288.

Rapoport:1947:MTMb

[171] Anatol Rapoport. Mathematical theory of motivation interactions of two individuals: II. Bulletin of Mathematical Biophysics, 9(2):41–61, June 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478292.

Bloch:1947:TMP

[172] Ingram Bloch. A theory of membrane permeability: III. The effect of hydrostatic pressure. Bulletin of Mathematical Biophysics, 9(2):63–74, June 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478293.

Esser:1947:MMEb

[173] Martinus H. M. Esser. The mechanism of the middle ear: Part II. The drum. Bulletin of Mathematical Biophysics, 9(2):75–91, June 1947. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478294.

Branson:1947:MDM

[174] Herman Branson. A mathematical description of metabolizing systems: II. Bulletin of Mathematical Biophysics, 9(2):93–98, June 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478295.

Landahl:1947:MCN

[175] H. D. Landahl. A matrix calculus for neural nets: II. Bulletin of Mathematical Biophysics, 9(2):99–108, June 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478296.

Rapoport:1947:FOD

[176] Anatol Rapoport. Forms of output distribution between two individuals motivated by a satisfaction function. Bulletin of Mathematical Biophysics, 9(3):109–122, September 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478289.

Rashevsky:1947:NMT

[177] N. Rashevsky. A note on the mathematical theory of vibrations of walls of blood vessels. Bulletin of Mathematical Biophysics, 9(3):123– 126, September 1947. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478290.

Pitts:1947:HWK

[178] Walter Pitts and Warren S. McCulloch. How we know universals the perception of auditory and visual forms. Bulletin of Mathematical Biophysics, 9(3):127–147, September 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478291.

Landahl:1947:ANI

[179] H. D. Landahl. An approach to a neurobiophysical interpretation of motivational interactions. Bulletin of Mathematical Biophysics, 9(4): 149–157, December 1947. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478306.

Shimbel:1947:CTO

[180] Alfonso Shimbel. Contributions to the theory of organic form: The intestinal tract. Bulletin of Mathematical Biophysics, 9(4):159– 168, December 1947. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478307.

Rapoport:1947:SEP

[181] Anatol Rapoport and Alfonso Shimbel. Suggested experimental procedure for determining the satisfaction function of animals. Bulletin of Mathematical Biophysics, 9(4):169–177, December 1947. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478308.

Rashevsky:1947:TAM

[182] N. Rashevsky. Theory of ameboid movements. Bulletin of Mathematical Biophysics, 9(4):179–184, December 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478309.

Householder:1947:MAB

[183] Alston S. Householder. Mathematical analysis of binocular vision. Bulletin of Mathematical Biophysics, 9(4):185–187, December 1947. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478310.

Roberts:1948:MBS

[184] J. B. Roberts. Mathematical biophysics of some neural nets. Bulletin of Mathematical Biophysics, 10(1):1–10, March 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478325.

Rashevsky:1948:LM

[185] N. Rashevsky. On the locomotion of mammals. Bulletin of Mathematical Biophysics, 10(1):11–23, March 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478326.

Brody:1948:SMF

[186] Daniel A. Brody and J. P. Quigley. Some mechanical factors in the dynamics of the thin-walled, spherical viscus. Bulletin of Mathematical Biophysics, 10(1):25–30, March 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478327.

Culbertson:1948:MONa

[187] James T. Culbertson. A mechanism for optic nerve conduction and form perception: I. Bulletin of Mathematical Biophysics, 10(1):31–40, March 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478328.

Shimbel:1948:SAT

[188] Alfonso Shimbel and Anatol Rapoport. A statistical approach to the theory of the central nervous system. Bulletin of Mathematical Biophysics, 10(1):41–55, March 1948. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478329.

Opatowski:1948:HA

[189] I. Opatowski and Alice M. Christiansen. On herbicidal actions. Bulletin of Mathematical Biophysics, 10(2):57–61, June 1948. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477432.

Wei:1948:MNN

[190] Telson Wei. On matrices of neural nets. Bulletin of Mathematical Biophysics, 10(2):63–67, June 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477433.

Kesselman:1948:MSO

[191] Russell H. Kesselman. A mathematical study of organism growth as an approach to the cancer problem. Bulletin of Mathematical Biophysics, 10(2):69–83, June 1948. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477434.

Anonymous:1948:A

[192] Anonymous. Addendum. Bulletin of Mathematical Biophysics, 10 (2):83–84, June 1948. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477435.

Rashevsky:1948:CES

[193] N. Rashevsky. Can elastic stresses in gels cause the elongation and division of a cell? Bulletin of Mathematical Biophysics, 10(2):85–89, June 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477436.

Landahl:1948:ECS

[194] H. D. Landahl and J. B. Kahn, Jr. On the effects of a constant subthreshold conditioning stimulus upon the response to a constant current test stimulus. Bulletin of Mathematical Biophysics, 10(2):91–95, June 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477437.

Culbertson:1948:MONb

[195] James T. Culbertson. A mechanism for optic nerve conduction and form perception: II. Bulletin of Mathematical Biophysics, 10(2):97–102, June 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477438.

Lotka:1948:PAO

[196] Alfred J. Lotka. Physical aspects of organic evolution. Bulletin of Mathematical Biophysics, 10(3):103–115, September 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477485.

Stephenson:1948:TMB

[197] John L. Stephenson. Theory of the measurement of blood flow by the dilution of an indicator. Bulletin of Mathematical Biophysics, 10(3): 117–121, September 1948. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477486.

Roberts:1948:IC

[198] J. B. Roberts. Interaction of cycles. Bulletin of Mathematical Biophysics, 10(3):123–129, September 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477487.

Shimbel:1948:ATS

[199] Alfonso Shimbel. An analysis of theoretical systems of differentiating nervous tissue. Bulletin of Mathematical Biophysics, 10(3):131– 143, September 1948. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477488.

Rapoport:1948:CDR

[200] Anatol Rapoport. Cycle distributions in random nets. Bulletin of Mathematical Biophysics, 10(3):145–157, September 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477489.

Rashevsky:1948:PMS

[201] N. Rashevsky. On periodicities in metabolizing systems. Bulletin of Mathematical Biophysics, 10(3):159–174, September 1948. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477490.

Hearon:1948:KBC

[202] John Z. Hearon. The kinetics of blood coagulation. Bulletin of Mathematical Biophysics, 10(3):175–186, September 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477491.

Landahl:1948:NUM

[203] H. D. Landahl. A note on the units of membrane permeability to water. Bulletin of Mathematical Biophysics, 10(3):187–190, September 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477492.

Morales:1948:TBT

[204] Manuel F. Morales and Robert E. Smith. On the theory of blood-tissue exchange of inert gases: VI. Validity of approximate uptake expressions. Bulletin of Mathematical Biophysics, 10(3):191–200, September 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477493.

Rashevsky:1948:NBP

[205] N. Rashevsky. A note on biological periodicities. Bulletin of Mathematical Biophysics, 10(4):201–204, December 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477501.

Rashevsky:1948:NPA

[206] N. Rashevsky. A note on the probabilistic approach to the cancer problem. Bulletin of Mathematical Biophysics, 10(4):205–210, December 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477502.

Rapoport:1948:SSRa

[207] Anatol Rapoport and Alfonso Shimbel. Steady states in random nets: I. Bulletin of Mathematical Biophysics, 10(4):211–220, December 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477503.

Rapoport:1948:SSRb

[208] Anatol Rapoport. Steady states in random nets: II. Bulletin of Mathematical Biophysics, 10(4):221–226, December 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477504.

Hearon:1948:CE

[209] John Z. Hearon. Crystalline enzymes. Bulletin of Mathematical Biophysics, 10(4):227–229, December 1948. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477505.

Rashevsky:1949:TVO

[210] N. Rashevsky. Time variable osmotic pressures produced by coupled reactions as a possible cause of cell division. Bulletin of Mathematical Biophysics, 11(1):1–7, March 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477908.

Rashevsky:1949:NDD

[211] N. Rashevsky. A note on the diffusion drag forces. Bulletin of Mathematical Biophysics, 11(1):9–13, March 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477909.

Rashevsky:1949:NCN

[212] N. Rashevsky. Note on a case of nonlinear diffusion. Bulletin of Mathematical Biophysics, 11(1):15–17, March 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477910.

Landahl:1949:VCN

[213] H. D. Landahl and R. J. Podolsky. On the velocity of conduction in nerve fibers with saltatory transmission. Bulletin of Mathematical Biophysics, 11(1):19–27, March 1949. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477911.

Hearon:1949:SSKa

[214] John Z. Hearon. The steady state kinetics of some biological systems: I. Bulletin of Mathematical Biophysics, 11(1):29–50, March 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477912.

Roberts:1949:GSS

[215] J. B. Roberts. The group structure of some neural nets. Bulletin of Mathematical Biophysics, 11(1):51–58, March 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477913.

Geiringer:1949:CLTa

[216] Hilda Geiringer. Contribution to the linkage theory of autopolyploids: I. Bulletin of Mathematical Biophysics, 11(2):59–82, June 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477494.

Hearon:1949:SSKb

[217] John Z. Hearon. The steady state kinetics of some biological systems: II. Bulletin of Mathematical Biophysics, 11(2):83–95, June 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477495.

Rashevsky:1949:DP

[218] N. Rashevsky. A diffusion problem. Bulletin of Mathematical Biophysics, 11(2):97–103, June 1949. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477496.

Rashevsky:1949:MBSa

[219] N. Rashevsky. Mathematical biology of social behavior. Bulletin of Mathematical Biophysics, 11(2):105–113, June 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477497.

Kesselman:1949:TGC

[220] Russell H. Kesselman. Tissue growth and cancer. Bulletin of Mathematical Biophysics, 11(2):115–138, June 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477498.

Hoffman:1949:TMI

[221] Joseph G. Hoffman. Theory of the mitotic index and its application to tissue growth measurement. Bulletin of Mathematical Biophysics, 11 (2):139–144, June 1949. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477499.

vonBonin:1949:CCC

[222] Gerhardt von Bonin. Cybernetics or control and communication in the animal and the machine. Bulletin of Mathematical Biophysics, 11 (2):145–147, June 1949. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477500.

Patlak:1949:NTF

[223] C. S. Patlak and H. D. Landahl. A note on the two-factor theory with rectification as applied to alternating current stimulation. Bulletin of Mathematical Biophysics, 11(3):149–155, September 1949. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478360.

Rashevsky:1949:MBSb

[224] N. Rashevsky. Mathematical biological of social behavior: II. Bulletin of Mathematical Biophysics, 11(3):157–163, September 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478361.

Shimbel:1949:IOP

[225] Alfonso Shimbel. Input-output problems in simple nerve-ganglion systems. Bulletin of Mathematical Biophysics, 11(3):165–171, September 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478362.

Rashevsky:1949:SSP

[226] N. Rashevsky. Some suggestions for a possible approach to a mathematical biophysics of mitosis. Bulletin of Mathematical Biophysics, 11(3): 173–182, September 1949. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478363.

Rapoport:1949:OPA

[227] Anatol Rapoport. Outline of a probabilistic approach to animal sociology: I. Bulletin of Mathematical Biophysics, 11(3):183–196, September 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478364.

Geiringer:1949:CLTb

[228] Hilda Geiringer. Contribution to the linkage theory of autopolyploids: II. Bulletin of Mathematical Biophysics, 11(3):197–219, September 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478365.

Best:1949:MAT

[229] J. Boyd Best, Jr. A mechanism for active transport of coenzyme-like substances with possible reference to auxin transport in plants. Bulletin of Mathematical Biophysics, 11(3):221–234, September 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478366.

Rashevsky:1949:MBA

[230] N. Rashevsky. The mathematical basis of the arts. Bulletin of Mathematical Biophysics, 11(3):235–237, September 1949. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478367.

Landahl:1950:NMD

[231] H. D. Landahl. Note on a mechanism of the distribution of wealth. Bulletin of Mathematical Biophysics, 12(1):1–6, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477339.

Rapoport:1950:OPA

[232] Anatol Rapoport. Outline of a probabilistic approach to animal sociology: III. Bulletin of Mathematical Biophysics, 12(1):7–17, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477340.

Opatowski:1950:SEH

[233] I. Opatowski and Alice M. Christiansen. On the single event hypothesis in radiogenetics. Bulletin of Mathematical Biophysics, 12(1):19–26, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477341.

Landau:1950:PRD

[234] H. G. Landau. A problem in radiobiology: Diffusion and recombination of ions. Bulletin of Mathematical Biophysics, 12(1):27–34, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477342.

Karreman:1950:NGC

[235] George Karreman. Note on the glycolytic coefficient. Bulletin of Mathematical Biophysics, 12(1):35–37, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477343.

Rashevsky:1950:NLT

[236] N. Rashevsky. A note on Landahl’s theory of psychophysical discrimination. Bulletin of Mathematical Biophysics, 12(1):39–42, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477344.

Landahl:1950:RABa

[237] H. D. Landahl. On the removal of air-borne droplets by the human respiratory tract: I. The lung. Bulletin of Mathematical Biophysics, 12(1):43–56, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477345.

Hearon:1950:SSKa

[238] John Z. Hearon. The steady state kinetics of some biological systems: III. Thermodynamic aspects. Bulletin of Mathematical Biophysics, 12 (1):57–83, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477346.

Hearon:1950:SSKb

[239] John Z. Hearon. The steady state kinetics of some biological systems: IV. Thermodynamic aspects. Bulletin of Mathematical Biophysics, 12 (1):85–106, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477347.

Rapoport:1950:MAL

[240] Anatol Rapoport. The mathematical analysis of logic. Bulletin of Mathematical Biophysics, 12(1):107, March 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477348.

Rapoport:1950:CPT

[241] Anatol Rapoport. Contribution to the probabilistic theory of neural nets: I. Randomization of refractory periods and of stimulus intervals. Bulletin of Mathematical Biophysics, 12(2):109–121, June 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478248.

Opatowski:1950:BTE

[242] I. Opatowski. On Blair’s theory of excitation and the role of internal energy sources. Bulletin of Mathematical Biophysics, 12(2):123–133, June 1950. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478249.

Hearon:1950:SCD

[243] John Z. Hearon. Some cellular diffusion problems based on Onsager’s generalization of Fick’s law. Bulletin of Mathematical Biophysics, 12 (2):135–159, June 1950. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478250.

Landahl:1950:RABb

[244] H. D. Landahl. On the removal of air-borne droplets by the human respiratory tract: II. The nasal passages. Bulletin of Mathematical Biophysics, 12(2):161–169, June 1950. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478251.

Landahl:1950:NTM

[245] H. D. Landahl. A note on a theoretical mechanism to represent the kinetics of some alkaline phosphatases. Bulletin of Mathematical Biophysics, 12(2):171–175, June 1950. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478252.

Landau:1951:DRSa

[246] H. G. Landau. On dominance relations and the structure of animal societies: I. Effect of inherent characteristics. Bulletin of Mathematical Biophysics, 13(1):1–19, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478336.

Epstein:1951:NMP

[247] Herman T. Epstein and Anatol Rapoport. A note on the McCulloch– Pitts neural net for heat–cold discrimination. Bulletin of Mathematical Biophysics, 13(1):21–22, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478337.

Hearon:1951:NEC

[248] John Z. Hearon. A note on the equations of conditioned reflex. Bulletin of Mathematical Biophysics, 13(1):23–26, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478338.

Opatowski:1951:PDE

[249] I. Opatowski. On permeability determination of the erythrocyte. Bulletin of Mathematical Biophysics, 13(1):27–29, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478339.

Reid:1951:DMI

[250] A. T. Reid. On the diffusion of metabolic intermediates. Bulletin of Mathematical Biophysics, 13(1):31–37, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478340.

Levinsohn:1951:NRA

[251] S. Levinsohn. A note on Rapoport’s approximate formula for the inputoutput curve in the case of bilateral pre-inhibition. Bulletin of Mathematical Biophysics, 13(1):39–40, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478341.

Opatowski:1951:MTE

[252] I. Opatowski. On the mathematical theories of excitation. Bulletin of Mathematical Biophysics, 13(1):41–45, March 1951. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478342.

Anderson:1951:NFT

[253] Norman Anderson and Anatol Rapoport. A note on facilitation and threshold phenomena. Bulletin of Mathematical Biophysics, 13(1): 47–49, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478343.

Rashevsky:1951:SMB

[254] N. Rashevsky. Suggestions for a mathematical biology of some cultural developments. Bulletin of Mathematical Biophysics, 13(1):51–59, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478344.

Reashevsky:1951:OMB

[255] N. Reashevsky. Outline of a mathematical biosociology of beliefs and prejudices. Bulletin of Mathematical Biophysics, 13(1):61–68, March 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478345.

McGill:1951:ERT

[256] William J. McGill and Walter A. Rosenblith. Electrical responses to two clicks: A simple statistical interpretation. Bulletin of Mathematical Biophysics, 13(2):69–77, June 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478353.

Rashevsky:1951:SMA

[257] N. Rashevsky. A suggestion for a mathematical approach to the problem of the effect of climatic factors on health. Bulletin of Mathematical Biophysics, 13(2):79–84, June 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478354.

Rapoport:1951:NDB

[258] Anatol Rapoport. Nets with distance bias. Bulletin of Mathematical Biophysics, 13(2):85–91, June 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478355.

Ackerman:1951:RBC

[259] Eugene Ackerman. Resonances of biological cells at audible frequencies. Bulletin of Mathematical Biophysics, 13(2):93–106, June 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478356.

Solomonoff:1951:CRN

[260] Ray Solomonoff and Anatol Rapoport. Connectivity of random nets. Bulletin of Mathematical Biophysics, 13(2):107–117, June 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478357.

Opatowski:1951:SDP

[261] I. Opatowski. On simultaneous determinations of the permeability of a membrane and of the diffusion coefficient in an adjacent medium. Bulletin of Mathematical Biophysics, 13(2):119–132, June 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478358.

Rapoport:1951:PDD

[262] Anatol Rapoport. The probability distribution of distinct hits on closely packed targets. Bulletin of Mathematical Biophysics, 13(2):133–138, June 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478359.

Rashevsky:1951:NTC

[263] N. Rashevsky. A note on the theory of communication through social channels. Bulletin of Mathematical Biophysics, 13(3):139– 146, September 1951. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478222.

Rashevsky:1951:NIB

[264] N. Rashevsky. A note on imitative behavior and information. Bulletin of Mathematical Biophysics, 13(3):147–151, September 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478223.

Reid:1951:SCP

[265] A. T. Reid and H. G. Landau. A suggested chain process for radiation damage. Bulletin of Mathematical Biophysics, 13(3):153–163, September 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478224.

Shimbel:1951:AMA

[266] Alfonso Shimbel. Applications of matrix algebra to communication nets. Bulletin of Mathematical Biophysics, 13(3):165–178, September 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478225.

Rapoport:1951:AMT

[267] Anatol Rapoport. “Addition” and “multiplication” theorems for the inputs of two neurons converging on a third. Bulletin of Mathematical Biophysics, 13(3):179–188, September 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478226.

Karreman:1951:CMB

[268] George Karreman. Contributions to the mathematical biology of excitation with particular emphasis on changes in membrane permeability and on threshold phenomena. Bulletin of Mathematical Biophysics, 13(3): 189–243, September 1951. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478227.

Landau:1951:DRSb

[269] H. G. Landau. On dominance relations and the structure of animal societies: II. Some effects of possible social factors. Bulletin of Mathematical Biophysics, 13(4):245–262, December 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477920.

Opatowski:1951:DSV

[270] I. Opatowski. Determination of the sensitive volume by irradiation and the ionization density. Bulletin of Mathematical Biophysics, 13(4): 263–271, December 1951. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477921.

Opatowski:1951:DRP

[271] I. Opatowski and F. R. Tangherlini. Diffusion and recombination of products of ionizing radiations. Bulletin of Mathematical Biophysics, 13(4): 273–288, December 1951. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477922.

Rashevsky:1951:MBB

[272] N. Rashevsky. Mathematical biosociology of beliefs and prejudices. Bulletin of Mathematical Biophysics, 13(4):289–301, December 1951. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477923.

Plunkett:1951:KIA

[273] R. L. Plunkett and C. L. Gemmill. The kinetics of invertase action. Bulletin of Mathematical Biophysics, 13(4):303–312, December 1951. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477924.

Reich:1951:GGA

[274] Edgar Reich. The game of “gossip” analyzed by the theory of information. Bulletin of Mathematical Biophysics, 13(4):313–318, December 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477925.

Shimbel:1951:NPC

[275] A. Shimbel. A note on the problem of cycles in random nets. Bulletin of Mathematical Biophysics, 13(4):319–321, December 1951. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477926.

Landahl:1951:NIC

[276] H. D. Landahl. A neurobiophysical interpretation of certain aspects of the problem of risks. Bulletin of Mathematical Biophysics, 13(4): 323–335, December 1951. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477927.

Rapoport:1951:CBN

[277] Anatol Rapoport. Consciousness and behavior: A neural analysis of behavior and consciousness. Bulletin of Mathematical Biophysics, 13(4): 337–338, December 1951. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477928.

Chilton:1952:MAC

[278] Arthur B. Chilton and Ralph W. Stacy. A mathematical analysis of carbon dioxide respiration in man. Bulletin of Mathematical Biophysics, 14(1):1–18, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477818.

Morgan:1952:SLF

[279] G. W. Morgan. On the steady laminar flow of a viscous incompressible fluid in an elastic tube. Bulletin of Mathematical Biophysics, 14 (1):19–26, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477819.

Mayne:1952:AIM

[280] Robert Mayne. The audiogyral illusion and the mechanism of spatial representation. Bulletin of Mathematical Biophysics, 14(1):27–34, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477820.

Rapoport:1952:IPR

[281] Anatol Rapoport. “Ignition” phenomena in random nets. Bulletin of Mathematical Biophysics, 14(1):35–44, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477821.

Opatowski:1952:DDP

[282] I. Opatowski and George W. Schmidt. Determination of diffusion and permeability coefficients in muscle. Bulletin of Mathematical Biophysics, 14(1):45–65, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477822.

Shimbel:1952:SEC

[283] A. Shimbel. Some elementary considerations of neural models. Bulletin of Mathematical Biophysics, 14(1):67–71, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477823.

Rapoport:1952:IOC

[284] Anatol Rapoport. Input output curves of aggregates of “simple counter” neurons. Bulletin of Mathematical Biophysics, 14(1):73–83, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477824.

Opatowski:1952:TDE

[285] I. Opatowski. On the theory of diffusion of electrolytes. Bulletin of Mathematical Biophysics, 14(1):85–94, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477825.

Rashevsky:1952:PDI

[286] N. Rashevsky. Prolegomena to a dynamics of ideologies. Bulletin of Mathematical Biophysics, 14(1):95–118, March 1952. CODEN BMBIAO.

ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477826.

Anonymous:1952:APDa

[287] Anonymous. Awards for post-doctoral study in statistics at the University of Chicago. Bulletin of Mathematical Biophysics, 14(1):119, March 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477827.

Wijsman:1952:PTD

[288] Robert A. Wijsman. Probability that two diffusing molecules will collide. Bulletin of Mathematical Biophysics, 14(2):121–126, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477710.

MacKay:1952:LIC

[289] Donald M. MacKay and Warren S. McCulloch. The limiting information capacity of a neuronal link. Bulletin of Mathematical Biophysics, 14 (2):127–135, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477711.

Rashevsky:1952:PEB

[290] N. Rashevsky. The problem of exchange between two or more individuals, motivated by hedonistic considerations. Bulletin of Mathematical Biophysics, 14(2):137–140, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477712.

Shimbel:1952:CHN

[291] Alfonso Shimbel. Communication in a hierarchical network. Bulletin of Mathematical Biophysics, 14(2):141–151, June 1952. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477713.

Solomonoff:1952:EMC

[292] R. Solomonoff. An exact method for the computation of the connectivity of random nets. Bulletin of Mathematical Biophysics, 14(2): 153–157, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477714.

Rapoport:1952:CMT

[293] Anatol Rapoport. Contribution to the mathematical theory of mass behavior: I. The propagation of single acts. Bulletin of Mathematical Biophysics, 14(2):159–169, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477715.

Rapoport:1952:POL

[294] Anatol Rapoport. Periodicities of open linear systems with positive steady states. Bulletin of Mathematical Biophysics, 14(2):171–183, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477716.

Macey:1952:EHM

[295] Robert Macey. The excitation of the heart and its modification under the influence of the chemical mediators and the cardiac nerves: I. Some general consequences of the one-factor theory. Bulletin of Mathematical Biophysics, 14(2):185–192, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477717.

Rashevsky:1952:EEF

[296] N. Rashevsky. The effect of environmental factors on the rates of cultural developments. Bulletin of Mathematical Biophysics, 14(2):193–201, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477718.

Landau:1952:SPR

[297] H. G. Landau. On some problems of random nets. Bulletin of Mathematical Biophysics, 14(2):203–212, June 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477719.

Rashevsky:1952:MBD

[298] N. Rashevsky. Mathematical biology of division of labor between two individuals or two social groups. Bulletin of Mathematical Biophysics, 14(3):213–227, September 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477815.

Schmidt:1952:MTC

[299] George W. Schmidt. A mathematical theory of capillary exchange as a function of tissue structure. Bulletin of Mathematical Biophysics, 14(3):

229–263, September 1952. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477816.

Taylor:1952:DMC

[300] Robert E. Taylor. The distribution of membrane current in nerve with longitudinal linearly increasing applied current. Bulletin of Mathematical Biophysics, 14(3):265–292, September 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477817.

Rashevsky:1952:SSN

[301] N. Rashevsky. Some suggestions for a new theory of cell division. Bulletin of Mathematical Biophysics, 14(4):293–305, December 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477846.

Goodman:1952:OAS

[302] Leo A. Goodman. On optimal arrangements in some social learning situations. Bulletin of Mathematical Biophysics, 14(4):307–312, December 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477847.

Reid:1952:NGB

[303] A. T. Reid. Note on the growth of bacterial populations. Bulletin of Mathematical Biophysics, 14(4):313–316, December 1952. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477848.

Landahl:1952:MBC

[304] H. D. Landahl. Mathematical biophysics of color vision. Bulletin of Mathematical Biophysics, 14(4):317–325, December 1952. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477849.

Karreman:1952:SCM

[305] George Karreman. Some contributions to the mathematical biology of blood circulation. Reflections of pressure waves in the arterial system. Bulletin of Mathematical Biophysics, 14(4):327–350, December 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477850.

Rapoport:1952:RTT

[306] Anatol Rapoport. Response time and threshold of a random net. Bulletin of Mathematical Biophysics, 14(4):351–363, December 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477851.

Trucco:1952:SVA

[307] Ernesto Trucco. The smallest value of the axon density for which ‘ignition’ can occur in a random net. Bulletin of Mathematical Biophysics, 14(4):365–374, December 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477852.

Rapoport:1952:MTR

[308] Anatol Rapoport and Lionel I. Rebhun. On the mathematical theory of rumor spread. Bulletin of Mathematical Biophysics, 14(4):375– 383, December 1952. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477853.

Bartlett:1952:DE

[309] J. H. Bartlett and R. A. Kromhout. The Donnan equilibrium. Bulletin of Mathematical Biophysics, 14(4):385–391, December 1952. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477854. See note [329].

Anonymous:1952:APDb

[310] Anonymous. Awards for post-doctoral study in statistics at the University of Chicago. Bulletin of Mathematical Biophysics, 14(4): 393, December 1952. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477855.

Bay:1953:TEM

[311] Z. Bay, M. C. Goodall, and A. Szent-Györgyi. The transmission of excitation from the membrane to actomyosin. Bulletin of Mathematical Biophysics, 15(1):1–13, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476362.

Hearon:1953:NDM

[312] John Z. Hearon. Nonlinear diffusion in metabolic systems. Bulletin of Mathematical Biophysics, 15(1):15–21, March 1953. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476363.

Hearon:1953:CASa

[313] John Z. Hearon. Comments on the approximate solution of the diffusion equation: I. Bulletin of Mathematical Biophysics, 15(1):23–31, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476364.

Eadie:1953:KS

[314] G. S. Eadie and F. Bernheim. The kinetics of the succinoxidase. Bulletin of Mathematical Biophysics, 15(1):33–42, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476365.

Hall:1953:IED

[315] T. A. Hall. The interpretation of exponential dose-effect curves. Bulletin of Mathematical Biophysics, 15(1):43–47, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476366.

Landahl:1953:AMS

[316] H. D. Landahl. An approximation method for the solution of diffusion and related problems. Bulletin of Mathematical Biophysics, 15 (1):49–61, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476367.

Rashevsky:1953:IBN

[317] N. Rashevsky. Imitative behavior in nonuniformly spatially distributed populations. Bulletin of Mathematical Biophysics, 15(1):63–71, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476368.

Isenberg:1953:CDS

[318] Irvin Isenberg. Cell division by swelling stresses. Bulletin of Mathematical Biophysics, 15(1):73–81, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476369.

Karreman:1953:SDO

[319] G. Karreman and H. D. Landahl. On spontaneous discharge obtained from a physicochemical model of excitation. Bulletin of Mathematical Biophysics, 15(1):83–91, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476370.

Cohn:1953:SCC

[320] David L. Cohn. Some considerations on capillary adjustment to changing metabolism of tissue. Bulletin of Mathematical Biophysics, 15(1): 93–101, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476371.

Wijsman:1953:NTM

[321] Robert A. Wijsman. Note on a theory of membrane permeability. Bulletin of Mathematical Biophysics, 15(1):103–104, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476372.

Kimball:1953:HUC

[322] A. W. Kimball. Heredity in uterine cancer. Bulletin of Mathematical Biophysics, 15(1):105–106, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476373.

MacKay:1953:LIC

[323] D. M. MacKay and W. S. McCulloch. The limiting information capacity of a neuronal link. Bulletin of Mathematical Biophysics, 15(1):107, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476374;
http://link.springer.com/content/pdf/10.1007/BF02476374.pdf.

Karreman:1953:SCM

[324] G. Karreman. Some contributions to the mathematical biology of blood circulation. Reflections of pressure waves in the arterial system. Bulletin of Mathematical Biophysics, 15(1):109, March 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476375;
http://link.springer.com/content/pdf/10.1007/BF02476375.pdf.

Hearon:1953:CASb

[325] John Z. Hearon. Comments on the approximate solution of the diffusion equation: II. Bulletin of Mathematical Biophysics, 15(2):111–119, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476376.

Hearon:1953:KLS

[326] John Z. Hearon. The kinetics of linear systems with special reference to periodic reactions. Bulletin of Mathematical Biophysics, 15 (2):121–141, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476377.

Landau:1953:DRS

[327] H. G. Landau. On dominance relations and the structure of animal societies: III the condition for a score structure. Bulletin of Mathematical Biophysics, 15(2):143–148, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476378.

Isenberg:1953:NFB

[328] Irvin Isenberg. A note on the flow of blood in capillary tubes. Bulletin of Mathematical Biophysics, 15(2):149–152, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476379.

Landahl:1953:NDE

[329] H. D. Landahl. Note on the Donnan equilibrium. Bulletin of Mathematical Biophysics, 15(2):153–159, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476380. See [309].

Karreman:1953:SCC

[330] George Karreman. Some considerations on a carrier mechanism as a model for ion accumulation. Bulletin of Mathematical Biophysics, 15 (2):161–165, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476381.

Landahl:1953:ATF

[331] H. D. Landahl and G. Karreman. Approximate treatment of the formation of a difference in the concentrations of an ion at the two sides of a membrane. Bulletin of Mathematical Biophysics, 15(2):167–171, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476382.

Landau:1953:CMT

[332] H. G. Landau and A. Rapoport. Contribution to the mathematical theory of contagion and spread of information: I. Spread through a thoroughly mixed population. Bulletin of Mathematical Biophysics, 15 (2):173–183, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476383.

Karreman:1953:CMB

[333] George Karreman. Contributions to the mathematical biophysics of the cardiovascular system. Bulletin of Mathematical Biophysics, 15 (2):185–195, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476384.

Rashevsky:1953:OMA

[334] N. Rashevsky. Outline of a mathematical approach to history. Bulletin of Mathematical Biophysics, 15(2):197–234, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476385.

Hearon:1953:CASc

[335] John Z. Hearon. Comments on the approximate solution of the diffusion equation: I. Bulletin of Mathematical Biophysics, 15(2):235, June 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476386;
http://link.springer.com/content/pdf/10.1007/BF02476386.pdf.

deVries:1953:SVC

[336] Hissel de Vries. On the subnormal variance in the counts of randomly distributed particles: I. Approximate treatment of the three-dimensional case and discussion of experiments. Bulletin of Mathematical Biophysics, 15(3):237–244, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476399.

Nijboer:1953:SVC

[337] B. R. A. Nijboer. On the subnormal variance in the counts of randomly distributed particles: II. Exact treatment of the one-dimensional case. Bulletin of Mathematical Biophysics, 15(3):245–250, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476400.

Wette:1953:ONT

[338] Reimut Wette. Outlines of a new theory of photoreception. Bulletin of Mathematical Biophysics, 15(3):251–260, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476401.

Wijsman:1953:CII

[339] Robert A. Wijsman. A critical investigation of the integral description of metabolizing systems. Bulletin of Mathematical Biophysics, 15(3): 261–268, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476402.

Hearon:1953:NIE

[340] John Z. Hearon. A note on the integral-equation description of metabolizing systems. Bulletin of Mathematical Biophysics, 15(3):269– 276, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476403.

Tricomi:1953:SDM

[341] Francesco G. Tricomi. On the statistical distribution of mutant bacteria. Bulletin of Mathematical Biophysics, 15(3):277–292, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476404.

Landahl:1953:NNI

[342] H. D. Landahl. Note on the neurobiophysical interpretation of imitative behavior. Bulletin of Mathematical Biophysics, 15(3):293– 300, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476405.

Karreman:1953:BFB

[343] George Karreman. Blood flow in branching circulatory systems during rest and activity. Bulletin of Mathematical Biophysics, 15(3):301– 309, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476406.

Patlak:1953:RWP

[344] Clifford S. Patlak. Random walk with persistence and external bias. Bulletin of Mathematical Biophysics, 15(3):311–338, September 1953. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476407.

Rashevsky:1953:SQA

[345] N. Rashevsky. Some quantitative aspects of history. Bulletin of Mathematical Biophysics, 15(3):339–359, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476408.

Reid:1953:ADS

[346] A. T. Reid. An age-dependent stochastic model of population growth. Bulletin of Mathematical Biophysics, 15(3):361–365, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476409.

Landahl:1953:SIT

[347] H. D. Landahl. On the spread of information with time and distance. Bulletin of Mathematical Biophysics, 15(3):367–381, September 1953. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476410.

Landau:1953:MSA

[348] H. G. Landau. Methods of statistical analysis. Bulletin of Mathematical Biophysics, 15(3):383, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476411.

Bey:1953:TEM

[349] Z. Bey, M. C. Goodall, and A. Szent-Györgyi. The transmission of excitation from the membrane to actomyosin. Bulletin of Mathematical Biophysics, 15(3):385, September 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476412;   http://link.springer.com/
content/pdf/10.1007/BF02476412.pdf.

Sangren:1953:MDE

[350] W. C. Sangren and C. W. Sheppard. A mathematical derivation of the exchange of a labeled substance between a liquid flowing in a vessel and an external compartment. Bulletin of Mathematical Biophysics, 15(4): 387–394, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476432.

Jacobs:1953:PDT

[351] Robert B. Jacobs. On the propagation of a disturbance through a viscous liquid flowing in a distensible tube of appreciable mass. Bulletin of Mathematical Biophysics, 15(4):395–409, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476433.

Stephenson:1953:TVD

[352] John L. Stephenson. Theory of the vacuum drying of frozen tissues. Bulletin of Mathematical Biophysics, 15(4):411–429, December 1953. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476434.

Patlak:1953:MCS

[353] Clifford S. Patlak. A mathematical contribution to the study of orientation of organisms. Bulletin of Mathematical Biophysics, 15(4):431– 476, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476435.

Schmidt:1953:TCC

[354] George W. Schmidt. The time course of capillary exchange. Bulletin of Mathematical Biophysics, 15(4):477–488, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476436.

Schmidt:1953:DDP

[355] George W. Schmidt. Determination of diffusion and permeability coefficients in nerve trunks. Bulletin of Mathematical Biophysics, 15(4): 489–500, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476437.

Shimbel:1953:SPC

[356] Alfonso Shimbel. Structural parameters of communication networks. Bulletin of Mathematical Biophysics, 15(4):501–507, December 1953. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476438.

Bierman:1953:ELM

[357] Arthur Bierman. Enzyme localization as a mechanism of apparent active transport. Bulletin of Mathematical Biophysics, 15(4):509– 522, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476439.

Rapoport:1953:SITa

[358] Anatol Rapoport. Spread of information through a population with sociostructural bias: I. Assumption of transitivity. Bulletin of Mathematical Biophysics, 15(4):523–533, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476440.

Rapoport:1953:SITb

[359] Anatol Rapoport. Spread of information through a population with sociostructural bias: II. Various models with partial transitivity. Bulletin of Mathematical Biophysics, 15(4):535–546, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476441.

Macey:1953:EHM

[360] Robert I. Macey. The excitation of the heart and its modification under the influence of the chemical mediators and the cardiac nerves: II. A particular case of the one-factor theory. Bulletin of Mathematical Biophysics, 15(4):547–559, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476442.

Landahl:1953:SRB

[361] H. D. Landahl. Symposium on radiobiology. The basic spects of radiation effects on living systems. Bulletin of Mathematical Biophysics, 15(4): 561–563, December 1953. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476443.

Chilton:1954:MAO

[362] Arthur B. Chilton, Delbert S. Barth, and Ralph W. Stacy. A mathematical analysis of oxygen respiration in man. Bulletin of Mathematical Biophysics, 16(1):1–14, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02481808.

Danziger:1954:MTP

[363] Lewis Danziger and George L. Elmergreen. Mathematical theory of periodic relapsing catatonia. Bulletin of Mathematical Biophysics, 16 (1):15–21, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481809.

Stephenson:1954:TDA

[364] John L. Stephenson. Theory for the design of apparatus for drying frozen tissues. Bulletin of Mathematical Biophysics, 16(1):23–43, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02481810.

Horwitz:1954:STC

[365] Leonard Horwitz. Some theoretical considerations relating to the meaning and measurement of chloroplast reducing potential. Bulletin of Mathematical Biophysics, 16(1):45–53, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481811.

Landahl:1954:TPC

[366] H. D. Landahl. Transient phenomena in capillary exchange. Bulletin of Mathematical Biophysics, 16(1):55–58, March 1954. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02481812.

Cohn:1954:OSV

[367] David L. Cohn. Optimal systems: I. The vascular system. Bulletin of Mathematical Biophysics, 16(1):59–74, March 1954. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02481813.

Rapoport:1954:SIT

[368] Anatol Rapoport. Spread of information through a population with socio-structural bias: III. Suggested experimental procedures. Bulletin of Mathematical Biophysics, 16(1):75–81, March 1954. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02481814.

Isenberg:1954:TNP

[369] Irvin Isenberg. On the theory of the nematic phase and its possible relation to the mitotic spindle structure. Bulletin of Mathematical Biophysics, 16(1):83–96, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02481815.

Bierman:1954:NTK

[370] Arthur Bierman. A note on the thermodynamics and kinetics of open and steady state systems. Bulletin of Mathematical Biophysics, 16(1):

97–101, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481816.

Karreman:1954:VPP

[371] George Karreman. On the velocity of propagation of pressure waves in an incompressible viscous fluid enclosed in a tube with an elastomeric wall. Bulletin of Mathematical Biophysics, 16(1):103–109, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02481817.

Karreman:1954:CMB

[372] George Karreman. Contributions to the mathematical biophysics of branched circulatory systems. Bulletin of Mathematical Biophysics, 16 (1):111–116, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481818.

Rashevsky:1954:CMB

[373] N. Rashevsky. A contribution to the mathematical biology of the rates of historical development. Bulletin of Mathematical Biophysics, 16(1): 117–128, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481819.

Landau:1954:CMT

[374] H. G. Landau and A. Rapoport. Contribution to the mathematical theory of contagion and spread of information: I. Spread through a thoroughly mixed population. Bulletin of Mathematical Biophysics, 16(1):129, March 1954. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481820;   http://link.springer.com/content/pdf/10.
1007/BF02481820.pdf.

Davis:1954:SDR

[375] Robert L. Davis. Structures of dominance relations. Bulletin of Mathematical Biophysics, 16(2):131–140, June 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478368.

Ackerman:1954:ETR

[376] Eugene Ackerman. An extension of the theory of resonances of biological cells: I. Effects of viscosity and compressibility. Bulletin of Mathematical Biophysics, 16(2):141–150, June 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478369. See erratum [447].

Landahl:1954:NIT

[377] H. D. Landahl. Note on the interpretation of tracer experiments in biological systems. Bulletin of Mathematical Biophysics, 16(2):151–154, June 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478370.

Landahl:1954:DEM

[378] H. D. Landahl. A difference equation method for the calculation of the probability that two molecules will collide. Bulletin of Mathematical Biophysics, 16(2):155–158, June 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478371.

Karreman:1954:RAS

[379] George Karreman. The resonance of the arterial system. Bulletin of Mathematical Biophysics, 16(2):159–170, June 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478372.

Landahl:1954:AMD

[380] H. D. Landahl. An approximation method for the determination of diffusion and permeability coefficients in nerve trunks. Bulletin of Mathematical Biophysics, 16(2):171–179, June 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478373.

Finn:1954:APB

[381] R. K. Finn. Accounting for periodicities in biology. Bulletin of Mathematical Biophysics, 16(2):181–182, June 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478374.

Rashevsky:1954:ITB

[382] N. Rashevsky. Information theory in biology. Bulletin of Mathematical Biophysics, 16(2):183–185, June 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478375.

Landau:1954:DCT

[383] H. G. Landau. The distribution of completion times for random communication in a task-oriented group. Bulletin of Mathematical Biophysics, 16(3):187–201, September 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478413.

Bierman:1954:SES

[384] Arthur Bierman. Studies on the effects of structure on the behavior of enzyme systems. Bulletin of Mathematical Biophysics, 16(3):203– 257, September 1954. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478414.

Hearon:1954:EKI

[385] John Z. Hearon and Robert Katzman. Enzyme kinetics: Inhibition equations derived from the inelastic collision hypothesis. Bulletin of Mathematical Biophysics, 16(4):259–277, December 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02484493.

Trucco:1954:SIB

[386] Ernesto Trucco. Studies in imitative behavior: A generalization of the Rashevsky model; its mathematical properties. Bulletin of Mathematical Biophysics, 16(4):279–316, December 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02484494.

Rashevsky:1954:TLS

[387] N. Rashevsky. Topology and life: In search of general mathematical principles in biology and sociology. Bulletin of Mathematical Biophysics, 16(4):317–348, December 1954. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02484495.

Anonymous:1954:APD

[388] Anonymous. Awards for post-doctoral study in statistics at the University of Chicago. Bulletin of Mathematical Biophysics, 16(4): 349, December 1954. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02484496.

Karlsson:1955:NSS

[389] Georg Karlsson. Note on the spread of a state in small social groups. Bulletin of Mathematical Biophysics, 17(1):1–5, March 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02481832.

Hearon:1955:NTM

[390] John Z. Hearon. Note on the theory of mass behavior. Bulletin of Mathematical Biophysics, 17(1):7–13, March 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481833.

Rapoport:1955:AINa

[391] Anatol Rapoport. Application of information networks to a theory of vision. Bulletin of Mathematical Biophysics, 17(1):15–33, March 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02481834.

Ackerman:1955:ETR

[392] Eugene Ackerman. An extension of the theory of resonances of biological cells: II. Cross-section in a plane wave. Bulletin of Mathematical Biophysics, 17(1):35–40, March 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02481835. See erratum [448].

Landau:1955:NIC

[393] H. G. Landau. Note on an inequality for the coefficients in Legendre polynomial expansions and its application to the theory of liquid phase transitions. Bulletin of Mathematical Biophysics, 17(1):41–44, March 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02481836.

Rashevsky:1955:NCP

[394] N. Rashevsky. Note on a combinatorial problem in topological biology. Bulletin of Mathematical Biophysics, 17(1):45–50, March 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02481837.

Zimmer:1955:SPP

[395] L. T. Zimmer. Some possible phase transitions in dilute colloidal solutions. Bulletin of Mathematical Biophysics, 17(1):51–61, March 1955.

CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02481838.

Landahl:1955:MMTa

[396] H. D. Landahl. A mathematical model for the temporal pattern of a population structure, with particular reference to the flour beetle. Bulletin of Mathematical Biophysics, 17(1):63–77, March 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02481839.

Rashevsky:1955:TLS

[397] N. Rashevsky. Topology and life: In search of general mathematical principles in biology and sociology. Bulletin of Mathematical Biophysics, 17(1):79, March 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481840;   http://link.springer.com/content/pdf/10.
1007/BF02481840.pdf.

Morowitz:1955:SOD

[398] Harold J. Morowitz. Some order-disorder considerations in living systems. Bulletin of Mathematical Biophysics, 17(2):81–86, June 1955. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477985.

Hart:1955:ATE

[399] H. E. Hart. Analysis of tracer experiments in non-conservative steadystate systems. Bulletin of Mathematical Biophysics, 17(2):87–94, June 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477986.

Karreman:1955:CMT

[400] George Karreman. A countercurrent mechanism with transverse diffusion. Bulletin of Mathematical Biophysics, 17(2):95–100, June 1955. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477987.

Cunningham:1955:SNE

[401] W. J. Cunningham. Simultaneous nonlinear equations of growth. Bulletin of Mathematical Biophysics, 17(2):101–110, June 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477988.

Rashevsky:1955:STT

[402] N. Rashevsky. Some theorems in topology and a possible biological implication. Bulletin of Mathematical Biophysics, 17(2):111–126, June 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477989.

Landahl:1955:NPP

[403] H. D. Landahl. Note on the periodic properties of biochemical systems represented by linear equations. Bulletin of Mathematical Biophysics, 17(2):127–129, June 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477990.

Landahl:1955:MMTb

[404] H. D. Landahl. A mathematical model for the temporal pattern of a population structure, with particular reference to the flour beetle: II. Competition between species. Bulletin of Mathematical Biophysics, 17 (2):131–140, June 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477991.

OSullivan:1955:DSCa

[405] D. G. O’Sullivan. Diffusion and simultaneous chemical reactions: I. A method for solving the equations of some systems in which a fixed concentration exists at a boundary. Bulletin of Mathematical Biophysics, 17(2):141–153, June 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477992.

Rapoport:1955:STC

[406] Anatol Rapoport. Some theoretical consequences of the allometric growth equations. Bulletin of Mathematical Biophysics, 17(3):155– 167, September 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477856.

Macey:1955:SRB

[407] Robert Macey. Studies on the relations between pressure and flow in the circulation. Bulletin of Mathematical Biophysics, 17(3):169– 205, September 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477857.

Rashevsky:1955:SRT

[408] N. Rashevsky. Some remarks on topological biology. Bulletin of Mathematical Biophysics, 17(3):207–218, September 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477858.

Cohn:1955:OSIa

[409] David L. Cohn. Optimal systems: II. The vascular system. Bulletin of Mathematical Biophysics, 17(3):219–227, September 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477859.

Rashevsky:1955:LIT

[410] N. Rashevsky. Life, information theory, and topology. Bulletin of Mathematical Biophysics, 17(3):229–235, September 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477860.

Trucco:1955:MTS

[411] Ernesto Trucco. Mathematical thinking in the social sciences. Bulletin of Mathematical Biophysics, 17(3):237–241, September 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477861.

OSullivan:1955:DSCb

[412] D. G. O’Sullivan. Diffusion and simultaneous chemical reactions: II. The equations of those systems in which transport occurs from one region to an adjoining region. Bulletin of Mathematical Biophysics, 17(4): 243–255, December 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477752.

FitzHugh:1955:MMT

[413] Richard FitzHugh. Mathematical models of threshold phenomena in the nerve membrane. Bulletin of Mathematical Biophysics, 17(4):257– 278, December 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477753.

Karreman:1955:TIC

[414] George Karreman. Topological information content and chemical reactions. Bulletin of Mathematical Biophysics, 17(4):279–285, December 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477754.

Patlak:1955:PCD

[415] Clifford S. Patlak. Potential and current distribution in nerve: The effect of the nerve sheath, the number of fibers, and the frequency of alternating current stimulation. Bulletin of Mathematical Biophysics, 17(4): 287–307, December 1955. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477755.

Cohn:1955:OSIb

[416] David L. Cohn. Optimal systems: III. Muscle. Bulletin of Mathematical Biophysics, 17(4):309–329, December 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477756.

Rapoport:1955:FS

[417] Anatol Rapoport. The foundations of statistics. Bulletin of Mathematical Biophysics, 17(4):331–335, December 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477757.

Rapoport:1955:AINb

[418] Anatol Rapoport. Application of information networks to a theory of vision. Bulletin of Mathematical Biophysics, 17(4):337, December 1955. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477758;
http://link.springer.com/content/pdf/10.1007/BF02477758.pdf.

Danziger:1956:TPH

[419] Lewis Danziger and George L. Elmergreen. The thyroid-pituitary homeostatic mechanism. Bulletin of Mathematical Biophysics, 18(1): 1–13, March 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477840.

Rapoport:1956:SGT

[420] Anatol Rapoport. Some game-theoretical aspects of parasitism and symbiosis. Bulletin of Mathematical Biophysics, 18(1):15–30, March 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477841.

Rashevsky:1956:GB

[421] N. Rashevsky. The geometrization of biology. Bulletin of Mathematical Biophysics, 18(1):31–56, March 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/ article/10.1007/BF02477842. See correction [436] and note [467].

Monk:1956:BCD

[422] G. W. Monk and R. D. Mattuck. Biological cloud dynamics. Bulletin of Mathematical Biophysics, 18(1):57–64, March 1956. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477843.

Trucco:1956:NRT

[423] Ernesto Trucco. A note on Rashevsky’s theorem about point-bases in topological biology. Bulletin of Mathematical Biophysics, 18(1): 65–85, March 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477844.

Trucco:1956:FBD

[424] Ernesto Trucco. French bibliographical digest. Bulletin of Mathematical Biophysics, 18(1):87, March 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477845.

Christie:1956:DST

[425] Lee S. Christie and R. Duncan Luce. Decision structure and time relations in simple choice behavior. Bulletin of Mathematical Biophysics, 18(2):89–112, June 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477834.

Rashevsky:1956:CTB

[426] N. Rashevsky. Contributions to topological biology: Some considerations on the primordial graph and on some possible transformations. Bulletin of Mathematical Biophysics, 18(2):113–128, June 1956. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477835.

Trucco:1956:NIC

[427] Ernesto Trucco. A note on the information content of graphs. Bulletin of Mathematical Biophysics, 18(2):129–135, June 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477836.

Landahl:1956:MBC

[428] H. D. Landahl. Mathematical biophysics of color vision II: Theory of color changes induced by the alternation of colors at various frequencies. Bulletin of Mathematical Biophysics, 18(2):137–149, June 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477837.

Gutstein:1956:GHL

[429] William H. Gutstein. A generalization of Hooke’s Law in muscle elasticity. Bulletin of Mathematical Biophysics, 18(2):151–170, June 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477838.

Anonymous:1956:ANE

[430] Anonymous. Announcement of new examinations. Bulletin of Mathematical Biophysics, 18(2):171, June 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477839.

Rashevsky:1956:WTE

[431] N. Rashevsky. What type of empirically verifiable predictions can topological biology make? Bulletin of Mathematical Biophysics, 18(3):173– 188, September 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481853. See note [452].

Rashevsky:1956:NMA

[432] N. Rashevsky. A neural mechanism for adjustment to optimal conditions, with possible reference to visual accommodation. Bulletin of Mathematical Biophysics, 18(3):189–198, September 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481854.

OSullivan:1956:DSCa

[433] D. G. O’Sullivan. Diffusion and simultaneous chemical reactions: III. The degree of localization achieved in cytochemical staining procedures. Bulletin of Mathematical Biophysics, 18(3):199–203, September 1956. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02481855.

Macey:1956:PAS

[434] Robert Macey. A probabilistic approach to some problems in bloodtissue exchange. Bulletin of Mathematical Biophysics, 18(3):205– 217, September 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481856.

Foster:1956:PSP

[435] Caxton Foster and Anatol Rapoport. Parasitism and symbiosis in an N-person non-constant-sum continuous game. Bulletin of Mathematical Biophysics, 18(3):219–231, September 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481857.

Rashevsky:1956:GBC

[436] N. Rashevsky. The geometrization of biology: A correction. Bulletin of Mathematical Biophysics, 18(3):233–235, September 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02481858;
http://link.springer.com/content/pdf/10.1007/BF02481858.pdf.

See [421].

Trucco:1956:ICG

[437] Ernesto Trucco. On the information content of graphs: Compound symbols; different states for each point. Bulletin of Mathematical Biophysics, 18(3):237–253, September 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02481859.

Harris:1956:DIM

[438] Joseph D. Harris. On the diffusion of ions in membranes. Bulletin of Mathematical Biophysics, 18(3):255–261, September 1956. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02481860.

OSullivan:1956:DSCb

[439] D. G. O’Sullivan. Diffusion and simultaneous chemical reactions: I. A method for solving the equations of some systems in which a fixed concentration exists at a boundary. Bulletin of Mathematical Biophysics, 18 (3):263, September 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481861;   http://link.springer.com/content/pdf/10.
1007/BF02481861.pdf.

OSullivan:1956:DSCc

[440] D. G. O’Sullivan. Diffusion and simultaneous chemical reactions: II. The equations of those systems in which transport occurs from one region to an adjoining region. Bulletin of Mathematical Biophysics, 18 (3):263, September 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481862;   http://link.springer.com/content/pdf/10.
1007/BF02481862.pdf.

Epstein:1956:MEI

[441] H. T. Epstein. A model of the effects of irradiation on bacterial ability to support phage growth. Bulletin of Mathematical Biophysics, 18(4): 265–270, December 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481875.

Patlak:1956:CTA

[442] Clifford S. Patlak. Contributions to the theory of active transport. Bulletin of Mathematical Biophysics, 18(4):271–315, December 1956. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02481876.

Rapoport:1956:AIC

[443] Anatol Rapoport. On the application of the information concept to learning theory. Bulletin of Mathematical Biophysics, 18(4):317– 321, December 1956. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02481877.

Rashevsky:1956:SMB

[444] N. Rashevsky. Studies in mathematical biosociology of imitative behavior: I. Effects of income distribution. Bulletin of Mathematical Biophysics, 18(4):323–336, December 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02481878.

George:1956:LNB

[445] F. H. George. Logical networks and behavior. Bulletin of Mathematical Biophysics, 18(4):337–348, December 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481879.

Ackerman:1956:RBC

[446] Eugene Ackerman. Resonances of biological cells at audible frequencies. Bulletin of Mathematical Biophysics, 18(4):349, December 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02481880;
http://link.springer.com/content/pdf/10.1007/BF02481880.pdf.

Ackerman:1956:EETa

[447] Eugene Ackerman. Erratum: An extension of the theory of resonances of biological cells: I. Effects of viscosity and compressibility. Bulletin of Mathematical Biophysics, 18(4):349–350, December 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02481881;
http://link.springer.com/content/pdf/10.1007/BF02481881.pdf.

See [376].

Ackerman:1956:EETb

[448] Eugene Ackerman. Erratum: An extension of the theory of resonances of biological cells: II. Cross-section in a plane wave. Bulletin of Mathematical Biophysics, 18(4):350, December 1956. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02481882;   http://link.springer.com/
content/pdf/10.1007/BF02481882.pdf. See [392].

Anonymous:1956:PFS

[449] Anonymous. Postdoctoral fellowship of sigma delta epsilon. Bulletin of Mathematical Biophysics, 18(4):351, December 1956. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02481883.

Ackerman:1957:ETR

[450] Eugene Ackerman. An extension of the theory of resonances of biological cells: III. Relationship of breakdown curves and mechanical Q. Bulletin of Mathematical Biophysics, 19(1):1–7, March 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02668287.

Danziger:1957:MME

[451] Lewis Danziger and George L. Elmergreen. Mathematical models of endocrine systems. Bulletin of Mathematical Biophysics, 19(1):9–18, March 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02668288.

Trucco:1957:TBN

[452] Ernesto Trucco. Topological biology: A note on Rashevsky’s transformation T. Bulletin of Mathematical Biophysics, 19(1):19–21, March 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02668289. See [431].

Cole:1957:TOI

[453] D. F. Cole, B.Sc., Ph.D., (I.C.I.Fellow) and J. F. Meredith, B.Sc. Targetorgan interaction: Considerations of the effects of sodium and water retaining hormones on renal function. Bulletin of Mathematical Biophysics, 19(1):23–39, March 1957. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02668290.

Rashevsky:1957:OPM

[454] N. Rashevsky. Outline of a possible mathematical approach to the problem of the effects of environment upon the incidence of some psychoneuroses. Bulletin of Mathematical Biophysics, 19(1):41–59, March 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02668291.

Anonymous:1957:ATE

[455] Anonymous. Analysis of tracer experiments: II. Non-conservative non-steady-state systems. Bulletin of Mathematical Biophysics, 19(1): 61–72, March 1957. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02668292.

Rapoport:1957:MRM

[456] Anatol Rapoport. Medical research: A midcentury survey. Bulletin of Mathematical Biophysics, 19(1):73–76, March 1957. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02668293.

Anonymous:1957:SCS

[457] Anonymous. The scientists’ committee on security. Bulletin of Mathematical Biophysics, 19(1):77–78, March 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02668294.

Vass:1957:IAB

[458] C. C. N. Vass, M.Sc., Ph.D., M.B., Ch.B. International abstracts of biological sciences. Bulletin of Mathematical Biophysics, 19(1):79–80, March 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02668295.

Beckmann:1957:EDP

[459] Martin J. Beckmann. On the equilibrium distribution of population in space. Bulletin of Mathematical Biophysics, 19(2):81–90, June 1957. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477881.

Rashevsky:1957:CTI

[460] N. Rashevsky. Contributions to the theory of imitative behavior. Bulletin of Mathematical Biophysics, 19(2):91–119, June 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477882.

Kerner:1957:SMI

[461] Edward H. Kerner. A statistical mechanics of interacting biological species. Bulletin of Mathematical Biophysics, 19(2):121–146, June 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477883.

Greene:1957:FVA

[462] Peter H. Greene. Factors in visual acuity: I. Neural inhibition and the visual perception of contours. Bulletin of Mathematical Biophysics, 19 (2):147–156, June 1957. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477884.

Landahl:1957:IEA

[463] H. D. Landahl. On the interpretation of the effect of area on the critical flicker frequency. Bulletin of Mathematical Biophysics, 19(2): 157–162, June 1957. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477885.

Sprott:1957:PDA

[464] D. A. Sprott. Probability distributions associated with distinct hits on targets. Bulletin of Mathematical Biophysics, 19(3):163–170, September 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477759.

Landahl:1957:PGU

[465] H. D. Landahl. Population growth under the influence of random dispersal. Bulletin of Mathematical Biophysics, 19(3):171–186, September 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477760.

George:1957:LNP

[466] F. H. George. Logical networks and probability. Bulletin of Mathematical Biophysics, 19(3):187–199, September 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477761.

Rashevsky:1957:NGB

[467] N. Rashevsky. A note on the geometrization of biology. Bulletin of Mathematical Biophysics, 19(3):201–204, September 1957. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477762. See [421].

Rashevsky:1957:RIP

[468] N. Rashevsky. Remark on an interesting problem in topological biology. Bulletin of Mathematical Biophysics, 19(3):205–208, September 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477763.

Patlak:1957:CTA

[469] Clifford S. Patlak. Contributions to the theory of active transport: II. The gate type non-carrier mechanism and generalizations concerning tracer flow, efficiency, and measurement of energy expenditure. Bulletin of Mathematical Biophysics, 19(3):209–235, September 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477764.

Rashevsky:1957:EEN

[470] N. Rashevsky. Experimental ecology of nutrition of fishes. Bulletin of Mathematical Biophysics, 19(3):237–240, September 1957. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477765.

Davison:1957:CSP

[471] B. Davison. The compatibility of the survival plateaux hypothesis with Gaussian distribution of population. Bulletin of Mathematical Biophysics, 19(4):241–246, December 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478415.

Kesselman:1957:PRB

[472] Russell H. Kesselman and Arthur Grishman. The physical rationale of ballistocardiography. Bulletin of Mathematical Biophysics, 19(4):247– 255, December 1957. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478416.

Rapoport:1957:CTR

[473] Anatol Rapoport. Contribution to the theory of random and biased nets. Bulletin of Mathematical Biophysics, 19(4):257–277, December 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478417.

Isenberg:1957:UOD

[474] Irvin Isenberg. The use of organic dyes in electron microscopy. Bulletin of Mathematical Biophysics, 19(4):279–292, December 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478418.

Yeh:1957:CSP

[475] Gordon C. K. Yeh and Johann Martinek. Comparison of surface potentials due to several singularity representations of the human heart. Bulletin of Mathematical Biophysics, 19(4):293–308, December 1957. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478419.

Trucco:1957:NCP

[476] Ernesto Trucco. Note on a combinatorial problem. Bulletin of Mathematical Biophysics, 19(4):309–336, December 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478420.

Anonymous:1957:PSS

[477] Anonymous. Postdoctoral study in statistics. Bulletin of Mathematical Biophysics, 19(4):337, December 1957. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478421.

Landahl:1958:TCP

[478] H. D. Landahl. Theoretical considerations on potentiation in drug interaction. Bulletin of Mathematical Biophysics, 20(1):1–23, March 1958.

CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476558.

Kaltofen:1958:BTW

[479] Herbert Kaltofen. Beitrag zur Theorie der Wirkung Mineralischer Nährstoffe. (German) [Contribution to the theory of the effect of mineral nutrients]. Bulletin of Mathematical Biophysics, 20(1):25–32, March 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476559.

Rigas:1958:KII

[480] Demetrios A. Rigas. Kinetics of isotope incorporation into the desoxyribonucleic acid (DNA) of tissues: Life span and generation time of cells. Bulletin of Mathematical Biophysics, 20(1):33–70, March 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476560.

Rashevsky:1958:CSG

[481] N. Rashevsky. A contribution to the search of general mathematical principles in biology. Bulletin of Mathematical Biophysics, 20(1): 71–93, March 1958. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476561.

Anonymous:1958:APB

[482] Anonymous. Announcement by perspectives in biology and medicine. Bulletin of Mathematical Biophysics, 20(1):95, March 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476562.

Bartholomay:1958:LBD

[483] Anthony F. Bartholomay. On the linear birth and death processes of biology as Markoff chains. Bulletin of Mathematical Biophysics, 20 (2):97–118, June 1958. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477571.

Lovgren:1958:MTD

[484] Birger Lövgren. A mathematical treatment of the development of colonies of different kinds of social wasps. Bulletin of Mathematical Biophysics, 20(2):119–148, June 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477572.

Westheimer:1958:NRC

[485] Gerald Westheimer. A note on the response characteristics of the extraocular muscle system. Bulletin of Mathematical Biophysics, 20(2): 149–153, June 1958. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477573.

Williams:1958:NCS

[486] C. M. Williams and H. D. Landahl. A note on the control of spontaneous activity in nervous systems. Bulletin of Mathematical Biophysics, 20 (2):155–160, June 1958. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477574.

Landahl:1958:ECB

[487] H. D. Landahl. On the estimation of cerebral blood flow obtained from electroplethysmographic observations. Bulletin of Mathematical Biophysics, 20(2):161–166, June 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477575.

Rashevsky:1958:NPM

[488] N. Rashevsky. A note on a possible mathematical approach to the theory of individual freedom. Bulletin of Mathematical Biophysics, 20 (2):167–174, June 1958. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477576.

Bartholomay:1958:SMC

[489] Anthony F. Bartholomay. Stochastic models for chemical reactions: I. Theory of the unimolecular reaction process. Bulletin of Mathematical Biophysics, 20(3):175–190, September 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478297.

Rapoport:1958:NRB

[490] Anatol Rapoport. Nets with reciprocity bias. Bulletin of Mathematical Biophysics, 20(3):191–201, September 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478298.

Yeh:1958:MRC

[491] G. C. K. Yeh, J. Martinek, and H. de Beaumont. Multipole representations of current generators in a volume conductor. Bulletin of Mathematical Biophysics, 20(3):203–216, September 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478299.

Williams:1958:RLB

[492] Clyde M. Williams. Representation of locality in a biological information system. Bulletin of Mathematical Biophysics, 20(3):217–230, September 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478300.

Scheer:1958:ATD

[493] Bradley T. Scheer. Active transport: Definitions and criteria. Bulletin of Mathematical Biophysics, 20(3):231–244, September 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478301.

Rosen:1958:RTB

[494] Robert Rosen. A relational theory of biological systems. Bulletin of Mathematical Biophysics, 20(3):245–260, September 1958. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478302.

Landahl:1958:NIP

[495] H. D. Landahl. A note on the interpretation of perseveration in discrimination. Bulletin of Mathematical Biophysics, 20(3):261–265, September 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478303.

Rashevsky:1958:CST

[496] N. Rashevsky. A comparison of set-theoretical and graph-theoretical approaches in topological biology. Bulletin of Mathematical Biophysics, 20(3):267–273, September 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478304.

Rashevsky:1958:NBR

[497] N. Rashevsky. A note on biotopology of reproduction. Bulletin of Mathematical Biophysics, 20(3):275–280, September 1958. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478305.

Roston:1959:MFC

[498] Sidney Roston. Mathematical formulation of cardiovascular dynamics by use of the Laplace transform. Bulletin of Mathematical Biophysics, 21(1):1–11, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476454.

Williamson:1959:SEU

[499] M. H. Williamson. Some extensions of the use of matrices in population theory. Bulletin of Mathematical Biophysics, 21(1):13–17, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476455.

Macey:1959:QSS

[500] Robert I. Macey. A quasi-steady-state approximation method for diffusion problems: I. Concentration dependent diffusion coefficients. Bulletin of Mathematical Biophysics, 21(1):19–32, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476456.

Yeh:1959:MRE

[501] G. C. K. Yeh and J. Martinek. Multipole representation of an eccentric dipole and an eccentric double-layer. Bulletin of Mathematical Biophysics, 21(1):33–60, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476457.

Martinek:1959:IPD

[502] J. Martinek, G. C. K. Yeh, and H. de Beaumont. The inconsistencies of present-day mathematical-physical methods pertaining to current generators in volume conductors used in electrocardiography. Bulletin of Mathematical Biophysics, 21(1):61–70, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476458.

Rosen:1959:DPC

[503] Robert Rosen. The DNA-protein coding problem. Bulletin of Mathematical Biophysics, 21(1):71–95, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476459.

Rashevsky:1959:NTB

[504] N. Rashevsky. A note on topological biology. Bulletin of Mathematical Biophysics, 21(1):97–100, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476460.

Rashevsky:1959:STA

[505] N. Rashevsky. A set-theoretical approach to biology. Bulletin of Mathematical Biophysics, 21(1):101–106, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476461.

Anonymous:1959:SRM

[506] Anonymous. Statistical research monographs. Bulletin of Mathematical Biophysics, 21(1):107, March 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476462.

Rosen:1959:RTB

[507] Robert Rosen. A relational theory of biological systems II. Bulletin of Mathematical Biophysics, 21(2):109–128, June 1959. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476354.

Patlak:1959:CSD

[508] Clifford S. Patlak. A contribution to the study of diffusion of neutral particles through pores. Bulletin of Mathematical Biophysics, 21 (2):129–140, June 1959. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476355.

Barakat:1959:NTS

[509] Richard Barakat. A note on the transient stage of the random dispersal of logistic populations. Bulletin of Mathematical Biophysics, 21 (2):141–151, June 1959. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476356.

Landahl:1959:NPG

[510] H. D. Landahl. A note on population growth under random dispersal. Bulletin of Mathematical Biophysics, 21(2):153–159, June 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476357.

Rashevsky:1959:SRMa

[511] N. Rashevsky. Some remarks on the mathematical theory of nutrition of fishes. Bulletin of Mathematical Biophysics, 21(2):161–183, June 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476358.

Rashevsky:1959:NNO

[512] N. Rashevsky. A note on the nature and origin of life. Bulletin of Mathematical Biophysics, 21(2):185–193, June 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476359.

Adolph:1959:FPS

[513] Alan R. Adolph. Feedback in physiological systems: An application of feedback analysis and stochastic models to neurophysiology. Bulletin of Mathematical Biophysics, 21(2):195–216, June 1959. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476360.

Kerner:1959:FCS

[514] Edward H. Kerner. Further considerations on the statistical mechanics of biological associations. Bulletin of Mathematical Biophysics, 21 (2):217–255, June 1959. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476361.

Altshuler:1959:CRD

[515] Bernard Altshuler. Calculation of regional deposition of aerosol in the respiratory tract. Bulletin of Mathematical Biophysics, 21(3):257– 270, September 1959. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477914.

Roston:1959:MRS

[516] Sidney Roston. Mathematical representation of some endocrinological systems. Bulletin of Mathematical Biophysics, 21(3):271–282, September 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477915.

Landahl:1959:NFF

[517] H. D. Landahl. Note on the fusion frequency for stimuli of alternating duration. Bulletin of Mathematical Biophysics, 21(3):283–287, September 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477916.

Rosen:1959:SFC

[518] Robert Rosen. Some further comments on the DNA–protein coding problem. Bulletin of Mathematical Biophysics, 21(3):289–297, September 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477917. See

comments [543].

Rashevsky:1959:SRMb

[519] N. Rashevsky. Some remarks on the mathematical aspects of automobile driving. Bulletin of Mathematical Biophysics, 21(3):299–308, September 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477918. See

remarks [684, 524, 547, 548, 561, 626, 519].

Rashevsky:1959:SPA

[520] N. Rashevsky. Suggestion for a possible approach to molecular biology. Bulletin of Mathematical Biophysics, 21(3):309–326, September 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477919.

Valentinuzzi:1959:MIM

[521] Maximo Valentinuzzi. A mathematical interpretation of motor reactions to fixed and moving stimulation of the eye. Bulletin of Mathematical Biophysics, 21(4):327–341, December 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477893.

Wise:1959:MUV

[522] M. E. Wise and J. G. Defares. A model for unequal ventilation of the lungs assuming a common dead space and two separate dead spaces. Bulletin of Mathematical Biophysics, 21(4):343–362, December 1959. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477894.

Bartholomay:1959:SMC

[523] Anthony F. Bartholomay. Stochastic models for chemical reactions: II. The unimolecular rate constant. Bulletin of Mathematical Biophysics, 21(4):363–373, December 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477895.

Rashevsky:1959:MBA

[524] N. Rashevsky. Mathematical biophysics of automobile driving. Bulletin of Mathematical Biophysics, 21(4):375–385, December 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477896. See [519].

Rosen:1959:LPI

[525] Robert Rosen. On a logical paradox implicit in the notion of a selfreproducing automation. Bulletin of Mathematical Biophysics, 21(4): 387–394, December 1959. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477897.

Landahl:1959:MBC

[526] H. D. Landahl. Mathematical biophysics of color vision: III. Color constancy. Bulletin of Mathematical Biophysics, 21(4):395–402, December 1959. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477898.

Anonymous:1959:ICL

[527] Anonymous. 1960 international Congress for logic, methodology and philosophy of science. Bulletin of Mathematical Biophysics, 21(4): 403, December 1959. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477899.

Stephenson:1960:TTLa

[528] John L. Stephenson. Theory of transport in linear biological systems: I. Fundamental integral equation. Bulletin of Mathematical Biophysics, 22(1):1–17, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Stephenson:1960:TTL

[529] John L. Stephenson. Theory of transport in linear biological systems: I. Fundamental integral equation. Bulletin of Mathematical Biophysics, 22(1):1–17, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477968.

Cope:1960:ERT

[530] Freeman W. Cope. An elastic reservoir theory of the human systemic arterial system using current data on aortic elasticity. Bulletin of Mathematical Biophysics, 22(1):19–40, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477969.

Hart:1960:ATE

[531] H. E. Hart. Analysis of tracer experiments: IV. The kinetics of general N compartment systems. Bulletin of Mathematical Biophysics, 22 (1):41–52, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477970.

Gutstein:1960:DPS

[532] William H. Gutstein and Heinz Scharfstein. Dilation-producing stresses and strains obtained with the aid of density measurements. Bulletin of Mathematical Biophysics, 22(1):53–61, March 1960. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477971.

Rashevsky:1960:IB

[533] N. Rashevsky. On imitative behavior. Bulletin of Mathematical Biophysics, 22(1):63–71, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477972.

Rashevsky:1960:CRB

[534] N. Rashevsky. Contributions to relational biology. Bulletin of Mathematical Biophysics, 22(1):73–84, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477973.

Rapoport:1960:DRL

[535] Anatol Rapoport. A derivation of a rote learning curve from the total uncertainty of a task. Bulletin of Mathematical Biophysics, 22 (1):85–97, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477974.

Roston:1960:NCQ

[536] Sidney Roston. A note concerning quantum mechanics and medicine. Bulletin of Mathematical Biophysics, 22(1):99–102, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477975.

Jardetzky:1960:DAT

[537] Oleg Jardetzky. On the definition of active transport. Bulletin of Mathematical Biophysics, 22(1):103–110, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477976.

Anonymous:1960:EI

[538] Anonymous. Events of interest. Bulletin of Mathematical Biophysics, 22(1):111, March 1960. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477977.

Stephenson:1960:TTLb

[539] John L. Stephenson. Theory of transport in linear biological systems: II. Multiflux problems. Bulletin of Mathematical Biophysics, 22(2): 113–138, June 1960. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478002.

Roy:1960:MSI

[540] Archie E. Roy. On a method of storing information. Bulletin of Mathematical Biophysics, 22(2):139–168, June 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478003.

Trucco:1960:NLS

[541] Ernesto Trucco. Note on a linear system of differential equations. Bulletin of Mathematical Biophysics, 22(2):169–180, June 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478004.

Bellman:1960:SMA

[542] Richard Bellman, John A. Jacquez, and Robert Kalaba. Some mathematical aspects of chemotherapy: I. One-organ models. Bulletin of Mathematical Biophysics, 22(2):181–198, June 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478005.

Rosen:1960:SFC

[543] Robert Rosen. Some further comments on the DNA–protein coding problem: A correction and a note. Bulletin of Mathematical Biophysics, 22 (2):199–205, June 1960. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478006. See [518].

Rashevsky:1960:RIS

[544] N. Rashevsky. The role of the individual in social dynamics. Bulletin of Mathematical Biophysics, 22(2):207–215, June 1960. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478007.

Macey:1960:KUH

[545] Robert I. Macey and A. V. Wolf. Kinetics of ultrafiltration hemodialysis. Bulletin of Mathematical Biophysics, 22(3):217–226, September 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478346.

Rosen:1960:QTA

[546] Robert Rosen. A quantum-theoretic approach to genetic problems. Bulletin of Mathematical Biophysics, 22(3):227–255, September 1960. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478347.

Rashevsky:1960:FCM

[547] N. Rashevsky. Further contributions to the mathematical biophysics of automobile driving. Bulletin of Mathematical Biophysics, 22(3):257– 262, September 1960. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478348. See [519].

Rashevsky:1960:CMB

[548] N. Rashevsky. A contribution to the mathematical biology of automobile driving: II. Passing as a case of psychophysical discrimination. Bulletin of Mathematical Biophysics, 22(3):263–267, September 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478349. See [519].

Scheer:1960:FFR

[549] Bradley T. Scheer. The flux-force relations across complex membranes with active transport. Bulletin of Mathematical Biophysics, 22(3):269– 284, September 1960. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478350.

Bartholomay:1960:MST

[550] Anthony F. Bartholomay. Molecular set theory: A mathematical representation for chemical reaction mechanisms. Bulletin of Mathematical Biophysics, 22(3):285–307, September 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478351.

Jacquez:1960:SMA

[551] John A. Jacquez, Richard Bellman, and Robert Kalaba. Some mathematical aspects of chemotherapy — II: The distribution of a drug in the body. Bulletin of Mathematical Biophysics, 22(3):309–322, September 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478352.

Gilbert:1960:RBI

[552] Daniel L. Gilbert. Relationship between ion distribution and membrane potential during a steady state. Bulletin of Mathematical Biophysics, 22(4):323–349, December 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476718.

Rashevsky:1960:LIT

[553] N. Rashevsky. Life, information theory, probability, and physics. Bulletin of Mathematical Biophysics, 22(4):351–364, December 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476719.

Rashevsky:1960:SRM

[554] N. Rashevsky. Some remarks on molecular biology. Bulletin of Mathematical Biophysics, 22(4):365–370, December 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476720.

Ivlev:1960:UFP

[555] V. S. Ivlev. On the utilization of food by planktophage fishes. Bulletin of Mathematical Biophysics, 22(4):371–389, December 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476721.

Bernard:1960:MSI

[556] S. R. Bernard. Mathematical studies of the interaction of respiratory gases with whole blood I. O_{2} absorption. Bulletin of Mathematical Biophysics, 22(4):391–415, December 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476722.

Sherman:1960:TID

[557] Herbert Sherman. On the theory of indicator-dilution methods under varying blood-flow conditions. Bulletin of Mathematical Biophysics, 22(4):417–424, December 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476723.

Anonymous:1960:SMT

[558] Anonymous. Symposium on mathematical theories of biological phenomena. Bulletin of Mathematical Biophysics, 22(4):425, December 1960. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476724.

Bernard:1961:MSI

[559] S. R. Bernard. Mathematical studies of the interaction of respiratory gases with whole blood II. CO_{2} absorption. Bulletin of Mathematical Biophysics, 23(1):1–14, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476569.

Rashevsky:1961:MII

[560] N. Rashevsky. On the mutual influence of individuals in a society. Bulletin of Mathematical Biophysics, 23(1):15–18, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476570.

Rashevsky:1961:CMB

[561] N. Rashevsky. Contribution to the mathematical biophysics of automobile driving. Bulletin of Mathematical Biophysics, 23(1):19–29, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476571. See

[519].

Brody:1961:EEL

[562] Daniel A. Brody, J. Carl Brashaw, and John W. Evans. The elements of an electrocardiographic lead tensor theory. Bulletin of Mathematical Biophysics, 23(1):31–42, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476572.

White:1961:FCA

[563] Harold White. The formation of cell assemblies. Bulletin of Mathematical Biophysics, 23(1):43–53, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476573.

Karreman:1961:CQB

[564] George Karreman. Contributions to quantum biology. I. Mobile electronic characteristics of riboflavin radicals. Bulletin of Mathematical Biophysics, 23(1):55–68, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476574.

Heinmets:1961:QAMa

[565] F. Heinmets and A. Herschman. Quantitative analysis of metabolic processes III. A model-system for enzyme synthesis by sequential induction and mathematical formulation of the process. Bulletin of Mathematical Biophysics, 23(1):69–89, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476575.

Landahl:1961:NMM

[566] H. D. Landahl. A note on mathematical models for the interaction of neural elements. Bulletin of Mathematical Biophysics, 23(1):91–97, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476576.

Roston:1961:GMM

[567] Sidney Roston. Glucose metabolism — a “mixed-cycle” system. Bulletin of Mathematical Biophysics, 23(1):99–103, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476577.

Harshbarger:1961:SRG

[568] Dr. Boyd Harshbarger. Southern Regional Graduate Summer Session in Statistics at Virginia Polytechnic Institute. Bulletin of Mathematical Biophysics, 23(1):105–106, March 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476578.

Anonymous:1961:SMT

[569] Anonymous. Symposium on mathematical theories of biological phenomena. Bulletin of Mathematical Biophysics, 23(1):107–108, March 1961.

CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476579.

Rashevsky:1961:BEA

[570] N. Rashevsky. Biological epimorphism, adequate design, and the problem of regeneration. Bulletin of Mathematical Biophysics, 23(2): 109–113, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477465.

Rashevsky:1961:CTE

[571] N. Rashevsky. A contribution to the theory of egoistic and altruistic interactions. Bulletin of Mathematical Biophysics, 23(2):115–134, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477466.

Karreman:1961:SQB

[572] George Karreman. Studies in quantum biology. II. The mobile electron characteristics of tryptophan^{+} in relation to those of FMN^{−}, FMNH and FMNH_{2}^{+}. Bulletin of Mathematical Biophysics, 23(2):135–140, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477467.

Kerner:1961:VLP

[573] Edward H. Kerner. On the Volterra–Lotka principle. Bulletin of Mathematical Biophysics, 23(2):141–157, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477468.

Landahl:1961:MCE

[574] H. D. Landahl. A model for conditioning and the effect of spaced trials in learning. Bulletin of Mathematical Biophysics, 23(2):159–163, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477469.

Rosen:1961:RTS

[575] Robert Rosen. A relational theory of the structural changes induced in biological systems by alterations in environment. Bulletin of Mathematical Biophysics, 23(2):165–171, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477470.

Patlak:1961:CTA

[576] Clifford S. Patlak. Contributions to the theory of active transport: III. The effect of drugs on active transport systems. Bulletin of Mathematical Biophysics, 23(2):173–195, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477471.

Kaltofen:1961:ESN

[577] Herbert Kaltofen. Über den Ertragsverlauf bei Steigen der Nährstoffkonzentration. (German) [On the yield curve when the nutrient concentration increases]. Bulletin of Mathematical Biophysics, 23(2):197–210, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477472.

Rashevsky:1961:LIT

[578] N. Rashevsky. Life, information theory, probability, and physics. Bulletin of Mathematical Biophysics, 23(2):211, June 1961. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477473;
http://link.springer.com/content/pdf/10.1007/BF02477473.pdf.

Anonymous:1961:PAU

[579] Anonymous. Pan American Union’s request for old scientific journal. Bulletin of Mathematical Biophysics, 23(2):213, June 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477474.

Anonymous:1961:GTM

[580] Anonymous. Granduate traineeship in mathematical biology and biometry. Bulletin of Mathematical Biophysics, 23(2):215–216, June 1961. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477475.

Hearon:1961:RCT

[581] John Z. Hearon. On the roots of a certain type of matrix. Bulletin of Mathematical Biophysics, 23(3):217–221, September 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476735.

Collins:1961:TGT

[582] R. E. Collins. Transport of gases through hemoglobin solution. Bulletin of Mathematical Biophysics, 23(3):223–232, September 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476736.

Rashevsky:1961:RBS

[583] N. Rashevsky. On relations between sets. Bulletin of Mathematical Biophysics, 23(3):233–235, September 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476737.

Rashevsky:1961:AMMa

[584] N. Rashevsky. Abstract mathematical molecular biology. Bulletin of Mathematical Biophysics, 23(3):237–260, September 1961. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476738.

Utz:1961:EPG

[585] W. R. Utz. The equations of population growth. Bulletin of Mathematical Biophysics, 23(3):261–262, September 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476739.

Yeh:1961:EMR

[586] Gordon C. K. Yeh. Eccentric multipole representations of current generators in a spherical volume conductor. Bulletin of Mathematical Biophysics, 23(3):263–276, September 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476740.

Heinmets:1961:QAMb

[587] F. Heinmets and A. Herschman. Quantitative analysis of metabolic processes. II. A model-system for the synthesis of the dissociable enzyme and mathematical formulation of the process. Bulletin of Mathematical Biophysics, 23(3):277–295, September 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476741.

Liu:1961:DLE

[588] In-Mao Liu. A derivation of latency equation. Bulletin of Mathematical Biophysics, 23(3):297–303, September 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476742.

Rosen:1961:HFD

[589] Robert Rosen. An hypothesis of Freese and the DNA–protein coding problem. Bulletin of Mathematical Biophysics, 23(3):305–318, September 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476743.

Anonymous:1961:PPFa

[590] Anonymous. Post-and predoctoral fellowships in mathematical biology. Bulletin of Mathematical Biophysics, 23(3):319, September 1961. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476744.

Rapoport:1961:PL

[591] Anatol Rapoport. The perfect learner. Bulletin of Mathematical Biophysics, 23(4):321–335, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476490.

Cope:1961:MCA

[592] Freeman W. Cope. A method for the computation of aortic distensibility in the living human patient and its use for the determination of the aortic effects of aging, drugs, and exercise. Bulletin of Mathematical Biophysics, 23(4):337–353, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476491.

Stahl:1961:DAM

[593] Walter R. Stahl. Dimensional analysis in mathematical biology I. General discussion. Bulletin of Mathematical Biophysics, 23(4):355– 376, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476492.

Nubar:1961:MPB

[594] Yves Nubar and Renato Contini. A minimal principle in biomechanics. Bulletin of Mathematical Biophysics, 23(4):377–391, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476493.

Rosen:1961:RCS

[595] Robert Rosen. On the role of chemical systems in the microphysical aspects of primary genetic mechanisms. Bulletin of Mathematical Biophysics, 23(4):393–403, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476494.

Rashevsky:1961:IBI

[596] N. Rashevsky. On imitative behavior. II. Time course of change from one behavior to another. Bulletin of Mathematical Biophysics, 23(4): 405–411, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476495.

Rashevsky:1961:UJC

[597] N. Rashevsky. On the unreliability of judgments of certain social situations by minority groups. Bulletin of Mathematical Biophysics, 23(4): 413–416, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476496.

Rashevsky:1961:AMMb

[598] N. Rashevsky. Abstract mathematical molecular biology. Bulletin of Mathematical Biophysics, 23(4):417, December 1961. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476497;
http://link.springer.com/content/pdf/10.1007/BF02476497.pdf.

Anonymous:1961:PPFb

[599] Anonymous. Post- and predoctoral fellowships in mathematical biology. Bulletin of Mathematical Biophysics, 23(4):419, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476498.

Anonymous:1961:A

[600] Anonymous. Announcements. Bulletin of Mathematical Biophysics, 23(4):421–422, December 1961. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476499.

Rashevsky:1962:CTI

[601] N. Rashevsky. Contributions to the theory of imitative behavior: The number of political parties as determined by biological and social factors. Bulletin of Mathematical Biophysics, 24(1):1–3, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477862.

Yilmaz:1962:CP

[602] Huseyin Yilmaz. On color perception. Bulletin of Mathematical Biophysics, 24(1):5–29, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477863.

Rosen:1962:NAR

[603] Robert Rosen. A note on abstract relational biologies. Bulletin of Mathematical Biophysics, 24(1):31–38, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477864.

Roy:1962:MSI

[604] Archie E. Roy. On a method of storing information: II. A further study of model properties. Bulletin of Mathematical Biophysics, 24 (1):39–69, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477865.

Hommes:1962:ICD

[605] F. A. Hommes and F. W. Zilliken. Induction of cell differentiation: III. A quantitative approach in the analysis of induction. Bulletin of Mathematical Biophysics, 24(1):71–80, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477866.

Stahl:1962:DAM

[606] Walter R. Stahl. Dimensional analysis in mathematical biology. II. Bulletin of Mathematical Biophysics, 24(1):81–108, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477867.

Rigas:1962:MPB

[607] Demetrios A. Rigas. Mathematical principles in biology and their applications. Bulletin of Mathematical Biophysics, 24(1):109–110, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477868.

Anonymous:1962:CSO

[608] Anonymous. Conference on self-organizing systems. Bulletin of Mathematical Biophysics, 24(1):111, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477869.

Anonymous:1962:GTM

[609] Anonymous. Graduate traineeship in mathematical biology and biometry. Bulletin of Mathematical Biophysics, 24(1):113–114, March 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477870.

Vos:1962:TRB

[610] J. J. Vos. A theory of retinal burns. Bulletin of Mathematical Biophysics, 24(2):115–128, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477421.

Roston:1962:VPC

[611] Sidney Roston. Variation of pressure with cycle length and duration of systole in the two-chambered cardiovascular model. Bulletin of Mathematical Biophysics, 24(2):129–135, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477422.

Cope:1962:MCH

[612] Freeman W. Cope. A method for calculation of human systolic and diastolic blood pressures using an elastic reservoir theory of the systemic arterial system, and some clinical and physiological applications. Bulletin of Mathematical Biophysics, 24(2):137–157, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477423.

Nelson:1962:MDN

[613] Pierre P. Nelson. Un modèle du neurone. (French) [a model of the neuron]. Bulletin of Mathematical Biophysics, 24(2):159–181, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477424.

Brody:1962:EGC

[614] Daniel A. Brody and J. Carl Bradshaw. The equivalent generator components of uniform double layers. Bulletin of Mathematical Biophysics, 24(2):183–195, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477425.

Yeh:1962:MRC

[615] Gordon C. K. Yeh. On multipole representation of current generators. Bulletin of Mathematical Biophysics, 24(2):197–207, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477426.

Mullin:1962:MS

[616] A. A. Mullin. On mutant sets. Bulletin of Mathematical Biophysics, 24(2):209–215, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477427.

Rashevsky:1962:NEE

[617] N. Rashevsky. A note on energy expenditure in walking on level ground and uphill. Bulletin of Mathematical Biophysics, 24(2):217–227, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477428.

Rashevsky:1962:FDL

[618] N. Rashevsky. On the function and design of the lung. Bulletin of Mathematical Biophysics, 24(2):229–242, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477429.

Rosen:1962:SRA

[619] Robert Rosen. Self-reproducing automation. Bulletin of Mathematical Biophysics, 24(2):243–245, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477430.

Anonymous:1962:PPF

[620] Anonymous. Post- and predoctoral fellowships in mathematical biology. Bulletin of Mathematical Biophysics, 24(2):246, June 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477431.

Greene:1962:LNNa

[621] Peter H. Greene. On looking for neural networks and “cell assemblies” that underlie behavior: I. A mathematical model. Bulletin of Mathematical Biophysics, 24(3):247–275, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477957.

Krohn:1962:EFE

[622] Lawrence H. Krohn. The electrocardiographic field equation. Bulletin of Mathematical Biophysics, 24(3):277–278, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477958.

Rosen:1962:DAT

[623] Robert Rosen. The derivation of D’Arcy Thompson’s theory of transformations from the theory of optimal design. Bulletin of Mathematical Biophysics, 24(3):279–290, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477959.

Macey:1962:AME

[624] Robert I. Macey. Approximation method for electrolyte diffusion through a fixed charge medium. Bulletin of Mathematical Biophysics, 24(3): 291–302, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477960.

Kelman:1962:TNE

[625] R. B. Kelman. A theoretical note on exponential flow in the proximal part of the mammalian nephron. Bulletin of Mathematical Biophysics, 24(3): 303–317, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477961.

Rashevsky:1962:ADP

[626] N. Rashevsky. Automobile driving as psychophysical discrimination. Bulletin of Mathematical Biophysics, 24(3):319–325, September 1962. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477962. See [519].

Rashevsky:1962:AMM

[627] N. Rashevsky. Abstract mathematical molecular biology: II some relational consequences of the “one gene-one enzyme” hypothesis. Bulletin of Mathematical Biophysics, 24(3):327–334, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477963.

Towber:1962:CFN

[628] Jacob Towber. A canonical form for neural nets without circles. Bulletin of Mathematical Biophysics, 24(3):335–343, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477964.

Trucco:1962:CFT

[629] Ernesto Trucco. La cinetica dei farmaci e dei traccianti radioattivi. (Italian) [The kinetics of drugs and radioactive tracers]. Bulletin of Mathematical Biophysics, 24(3):345–346, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477965.

Anonymous:1962:IIR

[630] Anonymous. Institute on Information Retrieval. Bulletin of Mathematical Biophysics, 24(3):347, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477966.

Anonymous:1962:SCI

[631] Anonymous. Short course for industrial information specialists. Bulletin of Mathematical Biophysics, 24(3):349, September 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477967.

White:1962:VAP

[632] Harold White. On visual adaptation: I. Photochemistry. Bulletin of Mathematical Biophysics, 24(4):351–359, December 1962. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477993.

Marmasse:1962:NGM

[633] Claude Marmasse. On a new graphical method in the theory of multiple fixations and its application to the Wurmserg theory of isohemagglutination. Bulletin of Mathematical Biophysics, 24(4):361–367, December 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477994.

Roston:1962:BG

[634] Sidney Roston. On biological growth. Bulletin of Mathematical Biophysics, 24(4):369–373, December 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477995.

Rosen:1962:CTR

[635] Robert Rosen. Church’s thesis and its relation to the concept of realizability in biology and physics. Bulletin of Mathematical Biophysics, 24(4): 375–393, December 1962. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477996.

Greene:1962:LNNb

[636] Peter H. Greene. On looking for neural networks and “cell assemblies” that underlie behavior: II. Neural realization of the mathematical model. Bulletin of Mathematical Biophysics, 24(4):395–411, December 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477997.

Karreman:1962:EPI

[637] George Karreman and George Eisenman. Electrical potentials and ionic fluxes in ion exchangers: I. “n type” non-ideal systems with zero current. Bulletin of Mathematical Biophysics, 24(4):413–427, December 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477998.

Klemm:1962:WQW

[638] Uwe Klemm. Über die Wahrscheinlichkeit eines quasiergodischen Wortes im Grenzfall unendlich grosser Wortlänge. (German) [On the probability of a quasiergodic word in the limit of infinitely large word length]. Bulletin of Mathematical Biophysics, 24(4):429–439, December 1962. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477999.

Rosen:1962:TGA

[639] Robert Rosen. The theory of graphs and its applications. Bulletin of Mathematical Biophysics, 24(4):441–443, December 1962. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478000.

Anonymous:1962:BTN

[640] Anonymous. Biomathematics training at North Carolina state College. Bulletin of Mathematical Biophysics, 24(4):445–446, December 1962. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478001.

Macey:1963:PFP

[641] Robert I. Macey. Pressure flow patterns in a cylinder with reabsorbing walls. Bulletin of Mathematical Biophysics, 25(1):1–9, March 1963. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477766.

Valentinuzzi:1963:ICC

[642] Maximo Valentinuzzi and M. Eugenio Valentinuzzi. Information content of chemical structures and some possible biological applications. Bulletin of Mathematical Biophysics, 25(1):11–27, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477767.

Landahl:1963:PRR

[643] H. D. Landahl. Particle removal by the respiratory system note on the removal of airborne particulates by the human respiratory tract with particular reference to the role of diffusion. Bulletin of Mathematical Biophysics, 25(1):29–39, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477768.

Rosen:1963:REI

[644] Robert Rosen. On the reversibility of environmentally induced alterations in abstract biological systems. Bulletin of Mathematical Biophysics, 25 (1):41–50, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477769.

Rashevsky:1963:MBL

[645] N. Rashevsky. Mathematical biology of learning to drive an automobile. Bulletin of Mathematical Biophysics, 25(1):51–58, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477770.

Rashevsky:1963:PADa

[646] N. Rashevsky. The principle of adequate design and the cardiovascular system. Bulletin of Mathematical Biophysics, 25(1):59–74, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477771.

Utz:1963:PBS

[647] W. R. Utz and P. E. Waltman. Periodicity and boundedness of solutions of generalized differential equations of growth. Bulletin of Mathematical Biophysics, 25(1):75–93, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477772.

Meretz:1963:WHG

[648] Wolfgang Meretz. Die Wabenzelle der Honigbiene. (German) [The honeycomb of honey bees]. Bulletin of Mathematical Biophysics, 25(1): 95–110, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477773.

Griffith:1963:FTN

[649] J. S. Griffith. A field theory of neural nets: I: Derivation of field equations. Bulletin of Mathematical Biophysics, 25(1):111–120, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477774.

Anonymous:1963:GTB

[650] Anonymous. Graduate traineeships in biometry. Bulletin of Mathematical Biophysics, 25(1):121, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477775.

Anonymous:1963:CUI

[651] Anonymous. Course on the uses of information theory in science and engineering. Bulletin of Mathematical Biophysics, 25(1):123, March 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477776.

White:1963:VAI

[652] Harold White. On visual adaptation: II. The electroretinogram and the bipolar cells. Bulletin of Mathematical Biophysics, 25(2):125–139, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478271.

Licko:1963:SNT

[653] Vojtech Ličko. Some notes on the theory of reaction rates: Enzymatic reactions. Bulletin of Mathematical Biophysics, 25(2):141–154, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478272.

Rashevsky:1963:NPD

[654] N. Rashevsky. Note on psychophysical discrimination with more than two stimuli. Bulletin of Mathematical Biophysics, 25(2):155–163, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478273.

Cope:1963:KTE

[655] Freeman W. Cope. A kinetic theory of enzymatic oxidation-reduction reactions based on a postulate of electron conduction in a macromolecular enzyme with an application to active transport of small ions across biological membranes. Bulletin of Mathematical Biophysics, 25 (2):165–176, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478274.

Sommerfield:1963:EPR

[656] M. Sommerfield. Elementary propositions on relations between sets. Bulletin of Mathematical Biophysics, 25(2):177–181, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478275.

Rosen:1963:NQT

[657] Robert Rosen. A note on the quantum-theoretic basis of primary genetic activity. Bulletin of Mathematical Biophysics, 25(2):183–187, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478276. See com-

ments [696].

Leibovic:1963:IPO

[658] K. N. Leibovic. Information processing organs, mathematical mappings and self-organizing systems. Bulletin of Mathematical Biophysics, 25 (2):189–201, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478277.

Thompson:1963:TBP

[659] W. A. Thompson, Jr. and G. H. Weiss. Transient behavior of population density with competition for resources. Bulletin of Mathematical Biophysics, 25(2):203–211, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478278.

Urbakh:1963:AII

[660] V. Yu. Urbakh. An account of intersymbol influences in the information theory: Application to linguistics and biology. Bulletin of Mathematical Biophysics, 25(2):213–229, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478279.

Rosen:1963:SRG

[661] Robert Rosen. Some results in graph theory and their application to abstract relational biology. Bulletin of Mathematical Biophysics, 25 (2):231–241, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478280.

Rosen:1963:PSA

[662] Robert Rosen. Proceedings of symposia in applied mathematics XIV: “Mathematical problems in the biological sciences”. Bulletin of Mathematical Biophysics, 25(2):243–246, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478281.

Anonymous:1963:ICL

[663] Anonymous. International Congress for logic, methodology and philosophy of science. Bulletin of Mathematical Biophysics, 25(2):247, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478282.

Anonymous:1963:PPF

[664] Anonymous. Post- and predoctoral fellowships in mathematical biology. Bulletin of Mathematical Biophysics, 25(2):249, June 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478283.

Bradner:1963:BLD

[665] Hugh Bradner and R. Stuart Mackay. Biophysical limitations on deep diving: Some limiting performance expectations. Bulletin of Mathematical Biophysics, 25(3):251–272, September 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02484022.

Bierman:1963:GTTa

[666] Arthur Bierman. A general target theory of radiobiological action. Bulletin of Mathematical Biophysics, 25(3):273–296, September 1963. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02484023.

Rashevsky:1963:NVP

[667] N. Rashevsky. A note on a variational problem in biology and sociology. Bulletin of Mathematical Biophysics, 25(3):297–301, September 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02484024.

Trucco:1963:FPEa

[668] Ernesto Trucco. On the Fokker–Planck equation in the stochastic theory of mortality: I. Bulletin of Mathematical Biophysics, 25(3):303– 323, September 1963. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02484025.

Apter:1963:AAP

[669] Julia T. Apter. An analysis of the aortic pressure curve during diastole. Bulletin of Mathematical Biophysics, 25(3):325–339, September 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02484026.

Rashevsky:1963:PADb

[670] N. Rashevsky. The principle of adequate design and the cardiovascular system. Bulletin of Mathematical Biophysics, 25(3):341, September 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02484027;
http://link.springer.com/content/pdf/10.1007/BF02484027.pdf.

Trucco:1963:FPEb

[671] Ernesto Trucco. On the Fokker–Planck equation in the stochastic theory of mortality: II. Bulletin of Mathematical Biophysics, 25(4):343– 366, December 1963. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476563.

Bierman:1963:GTTb

[672] Arthur Bierman. A general target theory of radiobiological action: II. Bulletin of Mathematical Biophysics, 25(4):367–385, December 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476564.

Filitti-Wurmser:1963:TFI

[673] S. Filitti-Wurmser, Y. Jacquot-Armand, and R. Wurmser. On the theory of fixation of isohemagglutinins on the erythrocytes. Bulletin of Mathematical Biophysics, 25(4):387–392, December 1963. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476565.

Martinez:1963:SSL

[674] H. M. Martinez. Studies in stochastic learning theory: I. Path-dependent linear models for experimenter–subject controlled experiments. Bulletin of Mathematical Biophysics, 25(4):393–419, December 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476566.

Nelson:1963:RCA

[675] P. P. Nelson. Reflexes et centres analyseurs dans les microstructures isotropes. (French) [Reflexes and analyzer centers in isotropic microstructures]. Bulletin of Mathematical Biophysics, 25(4):421–469, December 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476567.

Anonymous:1963:TPC

[676] Anonymous. Training program in computer use for biomedical scientists. Bulletin of Mathematical Biophysics, 25(4):471, December 1963. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476568.

Roston:1964:TBF

[677] Sidney Roston. Tissue blood flow and the sympathomimetic amines. Bulletin of Mathematical Biophysics, 26(1):1–7, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476615.

Richards:1964:AFC

[678] Whitman Richards. An analysis of the 2^{o} and 10^{o} field color-matching functions. Bulletin of Mathematical Biophysics, 26(1):9–24, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476616.

Hershberg:1964:PED

[679] Philip I. Hershberg. On the possibilities of error-detecting codes in nucleic acid molecules. Bulletin of Mathematical Biophysics, 26(1):25–29, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476617.

Rescigno:1964:STP

[680] Aldo Rescigno and Giorgio Segre. On some topological properties of the systems of compartments. Bulletin of Mathematical Biophysics, 26 (1):31–38, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476618.

Waltman:1964:EG

[681] P. E. Waltman. The equations of growth. Bulletin of Mathematical Biophysics, 26(1):39–43, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476619.

Lachlan:1964:NEP

[682] A. H. Lachlan. A note on the elementary α-perceptron. Bulletin of Mathematical Biophysics, 26(1):45–47, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476620.

Karreman:1964:AICa

[683] George Karreman. On the adsorption of ions at charged sites in an electric field: I. Bulletin of Mathematical Biophysics, 26(1):49–55, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476621.

Rashevsky:1964:TRM

[684] N. Rashevsky. Two remarks on the mathematical biology of automobile driving. Bulletin of Mathematical Biophysics, 26(1):57–61, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476622. See [519].

Martinez:1964:SSL

[685] H. M. Martinez. Studies in stochastic learning theory: II. Interpretation of response probabilities. Bulletin of Mathematical Biophysics, 26 (1):63–75, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476623.

Marmasse:1964:ASA

[686] Claude Marmasse and John Paul Brady. Analysis of a simple avoidance situation. II. A model. Bulletin of Mathematical Biophysics, 26 (1):77–81, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476624.

Landahl:1964:ALS

[687] H. D. Landahl. An avoidance learning situation. A neural net model. Bulletin of Mathematical Biophysics, 26(1):83–89, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476625.

Shikata:1964:GRG

[688] Morikazu Shikata. Group representation of genetic recombinations. Bulletin of Mathematical Biophysics, 26(1):91–100, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476626.

Marmasse:1964:CTM

[689] Claude Marmasse. Comparison of two models for a simple avoidance learning situation. Bulletin of Mathematical Biophysics, 26(1): 101, March 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476627.

Rosen:1964:ABSa

[690] Robert Rosen. Abstract biological systems as sequential machines. Bulletin of Mathematical Biophysics, 26(2):103–111, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476831.

Roston:1964:ACH

[691] Sidney Roston. The arterial circulation of the heart. Bulletin of Mathematical Biophysics, 26(2):113–120, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476832.

Goguen:1964:CPU

[692] J. A. Goguen, Jr. Color perception using a single cone type with distributed maximum sensitivity. Bulletin of Mathematical Biophysics, 26 (2):121–138, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476833.

Karreman:1964:AICb

[693] George Karreman. On the adsorption of ions at charged sites in an electric field: II. Bulletin of Mathematical Biophysics, 26(2):139–146, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476834.

Wissler:1964:MMH

[694] Eugene H. Wissler. A mathematical model of the human thermal system. Bulletin of Mathematical Biophysics, 26(2):147–166, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476835.

Rashevsky:1964:NMS

[695] N. Rashevsky. A neurobiophysical model of schizophrenias and of their possible treatment. Bulletin of Mathematical Biophysics, 26(2): 167–185, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476836.

Rashevsky:1964:SRR

[696] N. Rashevsky. Some remarks on Rosen’s quantum-mechanical approach to genetics. Bulletin of Mathematical Biophysics, 26(2):187–191, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476837. See

[657].

Landahl:1964:PMS

[697] H. D. Landahl, David Baldwin, and Theodore B. Schwartz. A physicomathematical model of sodium balance. Bulletin of Mathematical Biophysics, 26(2):193–198, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476838.

Anonymous:1964:JAP

[698] Anonymous. Journal of applied probability. Bulletin of Mathematical Biophysics, 26(2):199, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476839.

Anonymous:1964:TPC

[699] Anonymous. Training program in computer use for biomedical scientists. Bulletin of Mathematical Biophysics, 26(2):201, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476840.

Anonymous:1964:ICL

[700] Anonymous. 1964 international Congress for logic, methodology, and philosophy of science. Bulletin of Mathematical Biophysics, 26(2):203, June 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476841.

Mode:1964:SMD

[701] Charles J. Mode. A stochastic model of the dynamics of host-pathogen systems with mutation. Bulletin of Mathematical Biophysics, 26(3): 205–233, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02479043.

Yilmaz:1964:LP

[702] Huseyin Yilmaz. On the laws of psychophysics. Bulletin of Mathematical Biophysics, 26(3):235–237, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02479044.

Rosen:1964:ABSb

[703] Robert Rosen. Abstract biological systems as sequential machines II: Strong connectedness and reversibility. Bulletin of Mathematical Biophysics, 26(3):239–246, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02479045.

Scott:1964:AMN

[704] A. C. Scott. Analysis of a myelinated nerve model. Bulletin of Mathematical Biophysics, 26(3):247–254, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02479046.

Nelson:1964:IDS

[705] P. P. Nelson. Influence de la digestion sur le sommeil. (French) [Influence of digestion on sleep]. Bulletin of Mathematical Biophysics, 26(3): 255–263, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02479047.

Shea:1964:SIP

[706] Stephen M. Shea. On the stochastic interpretation of a pathological cell renewal system, I: Theoretical basis. Bulletin of Mathematical Biophysics, 26(3):265–274, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02479048.

Karreman:1964:AICc

[707] George Karreman. Adsorption of ions at charged sites and phase boundary potentials. Bulletin of Mathematical Biophysics, 26(3):275– 290, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02479049.

Ling:1964:SES

[708] Rufus C. Ling. On the stimulation and excitation of a system. Bulletin of Mathematical Biophysics, 26(3):291–294, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02479050.

Rosen:1964:BR

[709] Robert Rosen. Book review. Bulletin of Mathematical Biophysics, 26(3): 295–297, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02479051.

Lewis:1964:BR

[710] Robert V. Lewis, M.D. Book review. Bulletin of Mathematical Biophysics, 26(3):299–302, September 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02479052.

Trucco:1964:SAS

[711] Ernesto Trucco. Self-absorption in spheres and cylinders of radioactive material. Bulletin of Mathematical Biophysics, 26(4):303–325, December 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02484232.

Rashevsky:1964:MBA

[712] N. Rashevsky. Mathematical biology of automobile driving: I. The shape of the tracking curve on an empty straight road. Bulletin of Mathematical Biophysics, 26(4):327–332, December 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02484233.

Kerner:1964:DAK

[713] Edward H. Kerner. Dynamical aspects of kinetics. Bulletin of Mathematical Biophysics, 26(4):333–349, December 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02484234.

Martinez:1964:TOD

[714] H. M. Martinez. Toward an optimal design principle in relational biology. Bulletin of Mathematical Biophysics, 26(4):351–365, December 1964.

CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02484235.

Apter:1964:MDP

[715] Julia T. Apter. Mathematical development of a physical model of some visco-elastic properties of the aorta. Bulletin of Mathematical Biophysics, 26(4):367–388, December 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02484236.

Rosen:1964:CTB

[716] Robert Rosen. Control theory and biological systems. Bulletin of Mathematical Biophysics, 26(4):389–390, December 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02484237.

Karreman:1964:AICd

[717] George Karreman. Adsorption of ions at charged sites and phase boundary potentials. Bulletin of Mathematical Biophysics, 26(4):391, December 1964. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02484238;
https://link.springer.com/content/pdf/10.1007/BF02484238.pdf.

Anonymous:1965:BSN

[718] Anonymous. A biographical sketch of Nicolas Rashevsky. Bulletin of Mathematical Biophysics, 27(1S):3–4, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477255.

Landahl:1965:LRP

[719] H. D. Landahl. A letter read to Professor Rashevsky at a dinner in his honor December 11, 1964. Bulletin of Mathematical Biophysics, 27 (1S):5–10, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477256.

Rosen:1965:SCR

[720] Robert Rosen. Some comments on re-establishability. Bulletin of Mathematical Biophysics, 27(1S):11–14, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477257.

Licko:1965:C

[721] Vojtech Ličko. On compartmentalization. Bulletin of Mathematical Biophysics, 27(1S):15–19, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477258.

Ackerman:1965:MSB

[722] Eugene Ackerman, Laël C. Gatewood, John W. Rosevear, and George D. Molnar. Model studies of blood-glucose regulation. Bulletin of Mathematical Biophysics, 27(1S):21–37, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02477259.

Trucco:1965:UFE

[723] Ernesto Trucco. On the use of the Von Foerster equation for the solution and generalization of a problem studied by S. A. Tyler and R. Baserga. Bulletin of Mathematical Biophysics, 27(1S):39–48, January 1965. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477260.

Segre:1965:CSG

[724] Giorgio Segre. Compartmental systems and generating functions. Bulletin of Mathematical Biophysics, 27(1S):49–63, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477261.

Sheppard:1965:PAP

[725] C. W. Sheppard and G. B. Spurr. Periodic and aperiodic phases of circulatory mixing. Bulletin of Mathematical Biophysics, 27(1S): 65–70, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477262.

Defares:1965:TCS

[726] J. G. Defares, J. J. Osborn, and H. H. Hara. On the theory of the cardiovascular system. Bulletin of Mathematical Biophysics, 27(1S): 71–83, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477263.

Rescigno:1965:CEP

[727] Aldo Rescigno and Irvin W. Richardson. On the competitive exclusion principle. Bulletin of Mathematical Biophysics, 27(1S):85–89, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477264.

Karreman:1965:CSA

[728] George Karreman. Cooperative specific adsorption of ions at charged sites in an electric field. Bulletin of Mathematical Biophysics, 27(1S): 91–104, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477265.

Reiner:1965:PMS

[729] John M. Reiner. A proposed mechanism for the sol-gel transformation. Bulletin of Mathematical Biophysics, 27(1S):105–112, January 1965. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477266.

Cope:1965:KCS

[730] Freeman W. Cope. Kinetic criteria of solid state mechanisms in biology. Bulletin of Mathematical Biophysics, 27(1S):113–118, January 1965. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477267.

Apter:1965:ASE

[731] Julia T. Apter. Analysis of some effects of aortic smooth muscle on in vivo aortic pressure curves. Bulletin of Mathematical Biophysics, 27 (1S):119–130, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477268.

Martinez:1965:BSR

[732] Hugo M. Martinez. Biological similarity relative to allometric quantities. Bulletin of Mathematical Biophysics, 27(1S):131–133, January 1965. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477269.

White:1965:ECD

[733] Harold White. The entropy of a continuous distribution. Bulletin of Mathematical Biophysics, 27(1S):135–143, January 1965. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477270.

Woodger:1965:TRE

[734] J. H. Woodger. Theorems on random evolution. Bulletin of Mathematical Biophysics, 27(1S):145–150, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477271.

Landau:1965:DSS

[735] H. G. Landau. Development of structure in a society with a dominance relation when new members are added successively. Bulletin of Mathematical Biophysics, 27(1S):151–160, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477272.

Rapoport:1965:NEM

[736] Anatol Rapoport. A note on equivalent mathematical models. Bulletin of Mathematical Biophysics, 27(1S):161–175, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477273.

Shimbel:1965:ITG

[737] A. Shimbel. Information theory and genetics. Bulletin of Mathematical Biophysics, 27(1S):177–181, January 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477274.

Roston:1965:ECC

[738] Sidney Roston. Effects of the collateral circulation of the heart. Bulletin of Mathematical Biophysics, 27(1):1–8, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Spencer:1965:STT

[739] Richard P. Spencer. Simple theory of transfusions. Bulletin of Mathematical Biophysics, 27(1):9–20, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL https://link.springer.

com/article/10.1007/BF02476464.

Rashevsky:1965:SPQ

[740] N. Rashevsky. Some possible quantitative aspects of a neurophysiological model of schizophrenias. Bulletin of Mathematical Biophysics, 27(1):21– 26, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Apter:1965:AAP

[741] Julia T. Apter. An analysis of aortic pressure curves taking into account visco-elastic properties of the aorta and variations in peripheral resistance. Bulletin of Mathematical Biophysics, 27(1):27–52, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Kelman:1965:LDA

[742] R. B. Kelman. Longitudinal diffusion along the nephron during stop flow. Bulletin of Mathematical Biophysics, 27(1):53–56, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Kelman:1965:SSD

[743] R. B. Kelman. Steady-state diffusion through a finite pore into an infinite reservoir: An exact solution. Bulletin of Mathematical Biophysics, 27(1): 57–65, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). See erratum [776].

Brown:1965:EHD

[744] A. C. Brown. Equations of heat distribution within the body. Bulletin of Mathematical Biophysics, 27(1):67–78, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Hladky:1965:SFM

[745] Stephen B. Hladky. The single file model for the diffusion of ions through a membrane. Bulletin of Mathematical Biophysics, 27(1):79–86, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Hart:1965:DEC

[746] H. E. Hart. Determination of equilibrium constants and maximum binding capacities in complex In vitro systems: I. The mammillary system. Bulletin of Mathematical Biophysics, 27(1):87–98, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Cope:1965:TIT

[747] Freeman W. Cope. A theory of ion transport across cell surfaces by a process analogous to electron transport across liquid-solid interfaces. Bulletin of Mathematical Biophysics, 27(1):99–109, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Anonymous:1965:SRG

[748] Anonymous. Southern Regional Graduate Summer Session in Statistics at Oklahoma State University. Bulletin of Mathematical Biophysics, 27 (1):111, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Rashevsky:1965:IAO

[749] N. Rashevsky. Important announcement to our contributors and readers. Bulletin of Mathematical Biophysics, 27(1):113, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Ling:1965:LE

[750] Rufus C. Ling. Letter to the editor. Bulletin of Mathematical Biophysics, 27(1):114, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL https://link.springer.com/content/
pdf/10.1007/BF02476475.pdf.

Apter:1965:MDP

[751] Julia T. Apter. Mathematical development of a physical model of some visco-elastic properties of the aorta. Bulletin of Mathematical Biophysics, 27(1):115, March 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL https://link.springer.com/content/
pdf/10.1007/BF02476476.pdf.

Macey:1965:HRT

[752] Robert I. Macey. Hydrodynamics in the renal tubule. Bulletin of Mathematical Biophysics, 27(2):117–124, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Reiner:1965:EMD

[753] John M. Reiner. Evolution of macromolecules: I. Dual coding systems. Bulletin of Mathematical Biophysics, 27(2):125–132, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Norwich:1965:HCT

[754] Kenneth H. Norwich and Raymond Reiter. Homeostatic control of thyroxin concentration expressed by a set of linear differential equations. Bulletin of Mathematical Biophysics, 27(2):133–144, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

TenHoopen:1965:RIO

[755] M. Ten Hoopen and H. A. Reuver. Remark on the input–output relation of formalized neurons in case of randomized stimuli. Bulletin of Mathematical Biophysics, 27(2):145–152, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Martinez:1965:CTL

[756] H. M. Martinez. A convergence theorem for linear threshold elements. Bulletin of Mathematical Biophysics, 27(2):153–159, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Vayo:1965:SSV

[757] H. Westcott Vayo. The size of spherical virus capsomeres. Bulletin of Mathematical Biophysics, 27(2):161–166, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL https://link.

springer.com/article/10.1007/BF02498771.

Roston:1965:CEC

[758] Sidney Roston, M.D. The cardiovascular effects of the carotid sinus mechanism. Bulletin of Mathematical Biophysics, 27(2):167–173, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Rashevsky:1965:IB

[759] N. Rashevsky. On imitative behavior. Bulletin of Mathematical Biophysics, 27(2):175–185, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Griffith:1965:FTN

[760] J. S. Griffith. A field theory of neural nets: II. Properties of the field equations. Bulletin of Mathematical Biophysics, 27(2):187–195, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Hendrix:1965:TEF

[761] Charles E. Hendrix. Transmission of electric fields in cortical tissue: A model for the origin of the alpha rhythm. Bulletin of Mathematical Biophysics, 27(2):197–213, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Pavlidis:1965:NMS

[762] Theodosios Pavlidis. A new model for simple neural nets and its application in the design of a neural oscillator. Bulletin of Mathematical Biophysics, 27(2):215–229, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Rashevsky:1965:VSI

[763] N. Rashevsky. Victor Sergeevich Ivlev. Bulletin of Mathematical Biophysics, 27(2):231–232, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

Anonymous:1965:SUO

[764] Anonymous. Syracuse University offers new interdisciplinary academic programs. Bulletin of Mathematical Biophysics, 27(2):233, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

White:1965:VAP

[765] Harold White. On visual adaptation: I. Photochemistry. Bulletin of Mathematical Biophysics, 27(2):235, June 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL https://link.

springer.com/content/pdf/10.1007/BF02498779.pdf.

Cope:1965:GTP

[766] Freeman W. Cope. A generalized theory of particulate electron conduction enzymes applied to cytochrome oxidase. A theory of coupled electron and/or ion transport applied to pyruvate carboxylase. Bulletin of Mathematical Biophysics, 27(3):237–252, September 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478401.

Filmer:1965:APD

[767] David L. Filmer and John R. Cannon. Analysis of a procedure for the determination of kinetic rate constants. Bulletin of Mathematical Biophysics, 27(3):253–263, September 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478402.

Hearon:1965:CAA

[768] John Z. Hearon. A case of apparent active transport. I. Bulletin of Mathematical Biophysics, 27(3):265–269, September 1965. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478403.

Lewi:1965:FOW

[769] P. J. Lewi. Forced oscillations in a Windkessel model. Bulletin of Mathematical Biophysics, 27(3):271–280, September 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02478404.

Sherwood:1965:TCP

[770] S. L. Sherwood. A theoretical consideration of perception at lower levels. Bulletin of Mathematical Biophysics, 27(3):281–283, September 1965.

CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478405.

Rashevsky:1965:NCS

[771] N. Rashevsky. A note on the cybernetics of segregation. Bulletin of Mathematical Biophysics, 27(3):305–310, September 1965. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478407.

Rashevsky:1965:NIB

[772] N. Rashevsky. A note on imitative behavior. Bulletin of Mathematical Biophysics, 27(3):311–313, September 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02478408.

Rescigno:1965:SMP

[773] Aldo Rescigno and Giorgio Segre. On some metric properties of the systems of compartments. Bulletin of Mathematical Biophysics, 27(3): 315–323, September 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478409.

Weinbaum:1965:MME

[774] Sheldon Weinbaum. A mathematical model for the elastic and fluid mechanical behavior of the human eye. Bulletin of Mathematical Biophysics, 27(3):325–354, September 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478410.

Whirlow:1965:PFV

[775] D. K. Whirlow and W. T. Rouleau. Periodic flow of a viscous liquid in a thick-walled elastic tube. Bulletin of Mathematical Biophysics, 27(3): 355–370, September 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478411.

Kelman:1965:E

[776] R. B. Kelman. Erratum. Bulletin of Mathematical Biophysics, 27(3):371, September 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398

(electronic). URL http://link.springer.com/article/10.1007/
BF02478412;   https://link.springer.com/content/pdf/10.1007/
BF02478412.pdf. See [743].

Roston:1965:VC

[777] Sidney Roston. The venous circulation. Bulletin of Mathematical Biophysics, 27(4):373–377, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476842.

Herdan:1965:LMG

[778] R. Herdan. A logical model for growth and differentiation in multicelled organisms. Bulletin of Mathematical Biophysics, 27(4):379– 387, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476843.

Gurpide:1965:ESR

[779] Erlio Gurpide and Jonah Mann. Estimation of secretory rates of hormones from the specific activities of metabolites which have multiple secreted precursors. Bulletin of Mathematical Biophysics, 27(4):389– 406, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476844.

Stibitz:1965:MMU

[780] George R. Stibitz. A mathematical model of the urethra. Bulletin of Mathematical Biophysics, 27(4):407–415, December 1965. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476845.

Hart:1965:ATE

[781] H. E. Hart. Analysis of tracer experiments v: Integral equations of perturbation-tracer analysis. Bulletin of Mathematical Biophysics, 27(4): 417–429, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476846.

Leitner:1965:UTR

[782] Felix Leitner. The use of two radiotracers for transfer-rate determinations in closed two compartment systems. Bulletin of Mathematical Biophysics, 27(4):431–434, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476847.

Bernhard:1965:CWB

[783] R. Bernhard. Correlation “waves” in brain-like structures. Bulletin of Mathematical Biophysics, 27(4):435–447, December 1965. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476848.

Trucco:1965:MMCb

[784] E. Trucco. Mathematical models for cellular systems. The Von Foerster equation. Part II. Bulletin of Mathematical Biophysics, 27(4): 449–471, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476849.

Mulligan:1965:MMN

[785] James T. Mulligan and Arend Bouhuys. Mathematical models of nonuniform intrapulmonary gas distribution. Bulletin of Mathematical Biophysics, 27(4):473–476, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476850.

Rashevsky:1965:ROT

[786] N. Rashevsky. The representation of organisms in terms of predicates. Bulletin of Mathematical Biophysics, 27(4):477–491, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476851.

Pilkington:1965:EFE

[787] Theo C. Pilkington and John P. Boineau. On the electrocardiographic field equation. Bulletin of Mathematical Biophysics, 27(4):493– 495, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476852.

Rosen:1965:TOC

[788] Robert Rosen. Temporal organization in cells. Bulletin of Mathematical Biophysics, 27(4):497–500, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476853.

Rashevsky:1965:HWM

[789] N. Rashevsky. Harold A. White, March 4, 1933–July 27, 1965: Obituary notice. Bulletin of Mathematical Biophysics, 27(4):501–502, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476854.

Anonymous:1965:NMO

[790] Anonymous. 29th National Meeting of the Operations Research Society of America. Bulletin of Mathematical Biophysics, 27(4):503, December 1965. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476855.

Nelson:1966:EPE

[791] P. Nelson. Étude des premiers étages sensoriels. (French) [Study of the first sensory stages]. Bulletin of Mathematical Biophysics, 28 (1):1–10, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476387.

Cerny:1966:FVL

[792] Lawrence C. Cerny and Walter P. Walawender. The flow of a viscous liquid in a converging tube. Bulletin of Mathematical Biophysics, 28 (1):11–24, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476388.

Mode:1966:SMDa

[793] Charles J. Mode. Some multi-dimensional branching processes as motivated by a class of problems in mathematical genetics I. Bulletin of Mathematical Biophysics, 28(1):25–50, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476389.

Pavlidis:1966:DNN

[794] Theodosios Pavlidis. Design of neural nets with intermittent response and certain other relevant studies. Bulletin of Mathematical Biophysics, 28(1):51–74, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476390.

Houghton:1966:ALV

[795] Gerald Houghton. Academic learning, a variational model. Bulletin of Mathematical Biophysics, 28(1):75–90, March 1966. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476391.

Gunther:1966:TBS

[796] B. Günther and B. León de la Barra. Theory of biological similarities, nondimensional parameters and invariant numbers. Bulletin of Mathematical Biophysics, 28(1):91–102, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476392.

Leininger:1966:CRS

[797] C. W. Leininger. On chains of related sets. Bulletin of Mathematical Biophysics, 28(1):103–106, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476393.

Bramsen:1966:MAT

[798] John Bramsen. A matrix approach to the theory of biotopological mapping. Bulletin of Mathematical Biophysics, 28(1):107–116, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476394.

Rashevsky:1966:RBS

[799] N. Rashevsky. On relations between sets: II. Bulletin of Mathematical Biophysics, 28(1):117–124, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476395.

Rashevsky:1966:FSS

[800] N. Rashevsky. Further studies of some aspects of a neurobiophysical model of schizophrenia. Bulletin of Mathematical Biophysics, 28(1): 125–135, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476396.

Fox:1966:BEM

[801] Michael A. Fox. Biological effects of magnetic fields. Bulletin of Mathematical Biophysics, 28(1):137–138, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476397.

Anonymous:1966:BS

[802] Anonymous. Bionics symposium 1966. Bulletin of Mathematical Biophysics, 28(1):139, March 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476398.

Rosen:1966:ABS

[803] Robert Rosen. Abstract biological systems as sequential machines: III. Some algebraic aspects. Bulletin of Mathematical Biophysics, 28 (2):141–148, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476987.

Rosen:1966:NRS

[804] Robert Rosen. A note on replication in (M, R)-systems. Bulletin of Mathematical Biophysics, 28(2):149–151, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476988. See comments [842].

Demetrius:1966:ABS

[805] Lloyd A. Demetrius. Abstract biological systems as sequential machines: Behavioral reversibility. Bulletin of Mathematical Biophysics, 28(2):153–160, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476989.

Plonsey:1966:MTE

[806] Robert Plonsey. On multipole theory in electrocardiography. Bulletin of Mathematical Biophysics, 28(2):161–166, June 1966. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476990.

Thompson:1966:VSD

[807] Howard K. Thompson, Jr., Robert H. Peter, and Henry D. McIntosh. Ventricular septal defect with bidirectional shunting: Mathematical considerations. Bulletin of Mathematical Biophysics, 28(2):167–179, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476991.

Mode:1966:SMDb

[808] Charles J. Mode. Some multi-dimensional branching processes as motivated by a class of problems in mathematical genetics II. Bulletin of Mathematical Biophysics, 28(2):181–190, June 1966. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476992.

Guttman:1966:RRP

[809] Burton S. Guttman. A resolution of Rosen’s paradox for self-reproducing automata. Bulletin of Mathematical Biophysics, 28(2):191–194, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476993.

Roston:1966:CCM

[810] Sidney Roston. Control of the contractility of the myocardium in a feedback system. Bulletin of Mathematical Biophysics, 28(2):195–205, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476994.

Wu:1966:SDM

[811] Tai Te Wu. On the shape of a DNA molecule. Bulletin of Mathematical Biophysics, 28(2):207–216, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476995.

Shikata:1966:NGR

[812] Morikazu Shikata. A note on group ring expression of self-fertilized population. Bulletin of Mathematical Biophysics, 28(2):217–218, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476996.

Lefkovitch:1966:PGM

[813] L. P. Lefkovitch. A population growth model incorporating delayed responses. Bulletin of Mathematical Biophysics, 28(2):219–233, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476997.

Mosimann:1966:RPN

[814] J. E. Mosimann, M. B. Shapiro, C. R. Merril, D. F. Bradley, and J. E. Vinton. Reconstruction of protein and nucleic acid sequences: IV. The algebra of free monoids and the fragmentation stratagem. Bulletin of Mathematical Biophysics, 28(2):235–260, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476998.

Hart:1966:ATE

[815] H. E. Hart. Analysis of tracer experiments: VII. General multicompartment systems imbedded in non-homogeneous inaccessible media. Bulletin of Mathematical Biophysics, 28(2):261–282, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476999.

Rashevsky:1966:PBS

[816] N. Rashevsky. Physics, biology, and sociology: A reappraisal. Bulletin of Mathematical Biophysics, 28(2):283–308, June 1966. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477000.

Rashevsky:1966:NRB

[817] N. Rashevsky. A note on relations between sets. Bulletin of Mathematical Biophysics, 28(2):309–313, June 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477001.

Mode:1966:RTP

[818] Charles J. Mode. Restricted transition probabilities and their applications to some problems in the dynamics of biological populations. Bulletin of Mathematical Biophysics, 28(3):315–331, September 1966. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476813.

Mode:1966:MBP

[819] Charles J. Mode. A multidimensional birth process and its application to some problems in the dynamics of biological populations. Bulletin of Mathematical Biophysics, 28(3):333–345, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476814.

Nelson:1966:CMN

[820] P. Nelson. Comparaison des modeles de neurone de Hodgkin–Huxley et de Nelson. (French) [Comparison of neuron models of Hodgkin– Huxley and of Nelson]. Bulletin of Mathematical Biophysics, 28(3): 347–354, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476815.

Vayo:1966:MLV

[821] H. Westcott Vayo. The motion of the left ventricle. I. Bulletin of Mathematical Biophysics, 28(3):355–362, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476816.

Vayo:1966:TLV

[822] H. Westcott Vayo. The theory of the left ventricular aneurysm. Bulletin of Mathematical Biophysics, 28(3):363–370, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476817.

Foster:1966:REA

[823] B. L. Foster. Re-establishability in abstract biology. Bulletin of Mathematical Biophysics, 28(3):371–374, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476818.

Vargo:1966:NCC

[824] Louis G. Vargo. A note on crime control. Bulletin of Mathematical Biophysics, 28(3):375–378, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476819.

Licko:1966:IER

[825] Vojtech Ličko. Interactions in enzymatic reactions. Bulletin of Mathematical Biophysics, 28(3):379–390, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476820.

Hirsch:1966:DCS

[826] Henry R. Hirsch and Joseph Engelberg. Decay of cell synchronization: Solutions of the cell-growth equation. Bulletin of Mathematical Biophysics, 28(3):391–409, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476821.

Martinez:1966:DMG

[827] Hugo M. Martinez. On the derivation of a mean growth equation for cell cultures. Bulletin of Mathematical Biophysics, 28(3):411–416, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476822.

Zimmerman:1966:MTEa

[828] Stuart O. Zimmerman. A mathematical theory of enamel solubility and the onset of dental caries: I. The kinetics of dissolution of powdered enamel in acid buffer. Bulletin of Mathematical Biophysics, 28(3): 417–432, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476823.

Zimmerman:1966:MTEb

[829] Stuart O. Zimmerman. A mathematical theory of enamel solubility and the onset of dental caries: II. Some solubility equilibrium considerations of hydroxyapatite. Bulletin of Mathematical Biophysics, 28(3): 433–441, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476824.

Zimmerman:1966:MTEc

[830] Stuart O. Zimmerman. A mathematical theory of enamel solubility and the onset of dental caries: III. Development and computer simulation of a model of caries formation. Bulletin of Mathematical Biophysics, 28(3): 443–464, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476825.

Rashevsky:1966:MB

[831] N. Rashevsky. On mass behavior. Bulletin of Mathematical Biophysics, 28(3):465–475, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476826.

Rashevsky:1966:RPL

[832] N. Rashevsky. Reward and punishment in learning with possible reference to punishment for crime. Bulletin of Mathematical Biophysics, 28(3):477–481, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476827.

Karreman:1966:CSA

[833] George Karreman. Cooperative specific adsorption of iona at charged sites in an electric field. Bulletin of Mathematical Biophysics, 28(3): 483, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476829;   http://link.springer.com/content/pdf/10.
1007/BF02476829.pdf.

Leitner:1966:UTR

[834] Felix Leitner. The use of two radiotracers for transfer-rate determinations in closed two compartment systems. Bulletin of Mathematical Biophysics, 28(3):483, September 1966. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476828;   http://link.springer.com/content/pdf/10.
1007/BF02476828.pdf.

Anonymous:1966:ORA

[835] Anonymous. Operations research around the world meetings. Bulletin of Mathematical Biophysics, 28(3):485, September 1966. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476830.

Licko:1967:BSL

[836] Vojtech Ličko. Bimolecular systems: I. Linear systems of complexes. Bulletin of Mathematical Biophysics, 29(1):1–16, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476958.

Oldfield:1967:DHI

[837] D. G. Oldfield. DNA histograms of irradiated cells: The case of prophasetrapping. Bulletin of Mathematical Biophysics, 29(1):17–32, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476959.

Oldfield:1967:NAT

[838] D. G. Oldfield. Note on the analysis of thymidine labeling in cells with chromosome-number dispersion. Bulletin of Mathematical Biophysics, 29(1):33–40, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476960.

Oldfield:1967:DRD

[839] D. G. Oldfield. Dose-rate dependence of radiation-induced metaphase delay. Bulletin of Mathematical Biophysics, 29(1):41–56, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476961.

Stibitz:1967:BMM

[840] George R. Stibitz. Bistable mathematical and mechanical models related to the urethra. Bulletin of Mathematical Biophysics, 29(1):57–65, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476962.

Marsh:1967:TUF

[841] Donald J. Marsh, Robert B. Kelman, and Henry C. Howard. The theory of urine formation in water diuresis with implications for antidiuresis. Bulletin of Mathematical Biophysics, 29(1):67–89, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476963.

Rosen:1967:FCR

[842] Robert Rosen. Further comments on replication in (M, ℜ)-systems. Bulletin of Mathematical Biophysics, 29(1):91–94, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476964. See [804].

Bright:1967:BFEa

[843] Peter B. Bright. The basic flow equations of electrophysiology in the presence of chemical reactions: I. Development of the equations. Bulletin of Mathematical Biophysics, 29(1):95–121, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476965.

Bright:1967:BFEb

[844] Peter B. Bright. The basic flow equations of electrophysiology in the presence of chemical reactions: II. A practical application concerning the pH and voltage effects accompanying the diffusion of O_{2} through hemoglobin solution. Bulletin of Mathematical Biophysics, 29(1):123–138, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476966.

Rashevsky:1967:OSO

[845] N. Rashevsky. Organismic sets: Outline of a general theory of biological and social organisms. Bulletin of Mathematical Biophysics, 29 (1):139–152, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476967.

Cook:1967:DMH

[846] Gerald Cook and Lawrence Stark. Derivation of a model for the human eye-positioning mechanism. Bulletin of Mathematical Biophysics, 29(1):153–174, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476968.

Stibitz:1967:MUU

[847] George R. Stibitz. Model of urethra under uniform circumferential tension. Bulletin of Mathematical Biophysics, 29(1):175–179, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476969. See erra-

tum [873].

Rashevsky:1967:MBA

[848] N. Rashevsky. Mathematical biology of automobile driving. Bulletin of Mathematical Biophysics, 29(1):181–186, March 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476970.

Rashevsky:1967:NMB

[849] N. Rashevsky. A note on the mathematical biology of automobile driving. Bulletin of Mathematical Biophysics, 29(1):187–188, March 1967. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476971. See [519].

Rashevsky:1967:HGL

[850] N. Rashevsky. Hyman Garshin Landau. Bulletin of Mathematical Biophysics, 29(1):189–190, March 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476972.

Gregg:1967:FLS

[851] John R. Gregg. Finite Linnaean structures. Bulletin of Mathematical Biophysics, 29(2):191–206, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476893.

Robison:1967:APS

[852] Gerson B. Robison. Abstract power structures (APS) and related bumping orders. Bulletin of Mathematical Biophysics, 29(2):207–216, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476894.

Roberge:1967:SAM

[853] F. A. Roberge. Stability analysis of a mathematical neuron model. Bulletin of Mathematical Biophysics, 29(2):217–226, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:

//link.springer.com/article/10.1007/BF02476895. See comments

[954, 955].

Landahl:1967:NNM

[854] H. D. Landahl. A neural net model for masking phenomena. Bulletin of Mathematical Biophysics, 29(2):227–232, June 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476896.

Gluss:1967:MNF

[855] Brian Gluss. A model for neuron firing with exponential decay of potential resulting in diffusion equations for probability density. Bulletin of Mathematical Biophysics, 29(2):233–243, June 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476897.

Horsfield:1967:ABD

[856] K. Horsfield and G. Cumming. Angles of branching and diameters of branches in the human bronchial tree. Bulletin of Mathematical Biophysics, 29(2):245–259, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476898.

Carrasquer:1967:EAM

[857] Gaspar Carrasquer. An evaluation of anti-multiplier factors in kidneys with obstructed ureters. Bulletin of Mathematical Biophysics, 29 (2):261–266, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476899.

Valentinuzzi:1967:AMF

[858] Maximo Valentinuzzi. An analysis of the mechanical forces in each semicircular canal of the cat under single and combined rotations. Bulletin of Mathematical Biophysics, 29(2):267–289, June 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476900.

Pavlidis:1967:MML

[859] Theodosios Pavlidis. A mathematical model for the light affected system in the Drosophila eclosion rhythm. Bulletin of Mathematical Biophysics, 29(2):291–310, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476901.

Mirsky:1967:PVO

[860] I. Mirsky. Pulse velocities in an orthotropic elastic tube. Bulletin of Mathematical Biophysics, 29(2):311–318, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476902.

Hart:1967:ATE

[861] H. E. Hart. Analysis of tracer experiments: VIII. Integro-differential equation treatment of partly accessible, partly injectable multicompartment systems. Bulletin of Mathematical Biophysics, 29(2):319–333, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476903.

Darvey:1967:IBA

[862] I. G. Darvey and R. F. Matlak. An investigation of basic assumption in enzyme kinetics using results of the geometric theory of differential equations. Bulletin of Mathematical Biophysics, 29(2):335–341, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476904.

Mode:1967:PLB

[863] Charles J. Mode. On the probability a line becomes extinct before a favorable mutation appears. Bulletin of Mathematical Biophysics, 29 (2):343–348, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476905.

Reiner:1967:MBK

[864] John M. Reiner. Molecular biology and the kinetics of growth: I. Cell growth. Bulletin of Mathematical Biophysics, 29(2):349–361, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476906.

Scott:1967:MMN

[865] A. C. Scott. More on myelinated nerve model analysis. Bulletin of Mathematical Biophysics, 29(2):363–371, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476907.

Daley:1967:CSN

[866] D. J. Daley. Concerning the spread of news in a population of individuals who never forget. Bulletin of Mathematical Biophysics, 29 (2):373–376, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476908.

Rescigno:1967:SLT

[867] Aldo Rescigno and Irvin W. Richardson. The struggle for life: I. Two species. Bulletin of Mathematical Biophysics, 29(2):377–388, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476909.

Rashevsky:1967:OSB

[868] N. Rashevsky. Organismic sets and biological epimorphism. Bulletin of Mathematical Biophysics, 29(2):389–393, June 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476910.

Rashevsky:1967:MTP

[869] N. Rashevsky. Mathematical theory of the possible role of intercellular fluid and of vascularization on physiological periodicities. Bulletin of Mathematical Biophysics, 29(2):395–401, June 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476911.

Rescigno:1967:NRP

[870] Aldo Rescigno. A note on Rashevsky’s postulate of relational forces. Bulletin of Mathematical Biophysics, 29(2):403–404, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476912.

Cull:1967:GTF

[871] Paul Cull. General two factor models. Bulletin of Mathematical Biophysics, 29(2):405, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476913.

Anonymous:1967:PVB

[872] Anonymous. Out-of-print volumes and back issues of the Bulletin of Mathematical Biophysics. Bulletin of Mathematical Biophysics, 29 (2):407, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476914.

Stibitz:1967:EMU

[873] George R. Stibitz. Erratum: Model of urethra under uniform circumferential tension. Bulletin of Mathematical Biophysics, 29(2):409, June 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476915;
https://link.springer.com/content/pdf/10.1007/BF02476915.pdf.

See [847].

Wei:1967:PET

[874] Ling Y. Wei. Possibility of electron tunneling through a nerve membrane. Bulletin of Mathematical Biophysics, 29(3):411–418, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476580.

Barton:1967:AIC

[875] Charles Barton and Severin Raynor. Analytical investigation of cilia induced mucous flow. Bulletin of Mathematical Biophysics, 29(3):419– 428, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476581.

Bonnor:1967:GSC

[876] W. B. Bonnor. Growth of a single cell. Bulletin of Mathematical Biophysics, 29(3):429–436, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476582.

Elandt-Johnson:1967:ECP

[877] Regina C. Elandt-Johnson. Equilibrium conditions in polysomic inheritance for a panmictic population. Bulletin of Mathematical Biophysics, 29(3):437–449, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476583.

Singer:1967:MMA

[878] Alvin Singer. Mathematical model analysis of pressure pulse propagation of the aortic wall. Bulletin of Mathematical Biophysics, 29(3): 451–464, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476584.

Singer:1967:CVA

[879] Alvin Singer. Computed ventricular acoustic frequency. Bulletin of Mathematical Biophysics, 29(3):465–471, September 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476585.

Miekisz:1967:CMC

[880] Stanisław Miekisz. On a certain model of cardiovascular system. Bulletin of Mathematical Biophysics, 29(3):473–483, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476586.

Martisovits:1967:QDB

[881] V. Martišovitš and Š. Veis. Quantitative description of blood oxygenation process in oxygenator for extracorporal circulation. Bulletin of Mathematical Biophysics, 29(3):485–497, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476587.

Vayo:1967:MLV

[882] H. Westcott Vayo. The motion of the left ventricle: II. Bulletin of Mathematical Biophysics, 29(3):499–512, September 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476588.

Patlak:1967:ADM

[883] Clifford S. Patlak. Analysis of the distribution of materials within the blood-brain-cerebrospinal fluid system. Bulletin of Mathematical Biophysics, 29(3):513–531, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476589.

Holen:1967:EGE

[884] Jarle Holen and Seymour I. Schwartz. Effects of gradual expansion on arterial stenosis. Bulletin of Mathematical Biophysics, 29(3):533– 540, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476590.

Roston:1967:BFB

[885] Sidney Roston. The blood flow of the brain. Bulletin of Mathematical Biophysics, 29(3):541–548, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476591.

Deakin:1967:EDAa

[886] Michael A. B. Deakin. Erythrocyte distribution in arterial blood flow: I. Basic equations and anisotropic effects. Bulletin of Mathematical Biophysics, 29(3):549–563, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476592.

Deakin:1967:EDAb

[887] Michael A. B. Deakin. Erythrocyte distribution in arterial blood flow: II. The hypothesis of minimal energy dissipation. Bulletin of Mathematical Biophysics, 29(3):565–574, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476593.

Wu:1967:REX

[888] Tai Te Wu. A re-evaluation of X-ray diffraction patterns of DNA. Bulletin of Mathematical Biophysics, 29(3):575–581, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476594.

Cope:1967:TCH

[889] Freeman W. Cope. A theory of cell hydration governed by adsorption of water on cell proteins rather than by osmotic pressure. Bulletin of Mathematical Biophysics, 29(3):583–596, September 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476595.

Oldfield:1967:C

[890] D. G. Oldfield. Cytokinetics. Bulletin of Mathematical Biophysics, 29(3): 597–603, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476596.

Licko:1967:BSIa

[891] Vojtech Ličko. Bimolecular systems: II. Nonlinear systems of complexes. Bulletin of Mathematical Biophysics, 29(3):605–613, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476597.

Licko:1967:BSIb

[892] Vojtech Ličko. Bimolecular systems: III. Catalytic systems. Bulletin of Mathematical Biophysics, 29(3):615–623, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476598.

Dahnert:1967:ASV

[893] Klaus Dähnert. Die Abhängigkeit des Strategischen Verhaltens von Sportlern bei Wettläufen von den Möglichkeiten der Kommunikation und Informationsverarbeitung. (German) [The dependence of the strategic behavior of athletes in the course of events of possibilities of communication and information processing]. Bulletin of Mathematical Biophysics, 29(3):625–638, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476599.

Rashevsky:1967:RBS

[894] N. Rashevsky. On relations between sets: III. Bulletin of Mathematical Biophysics, 29(3):639–642, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476600.

Rashevsky:1967:PBS

[895] N. Rashevsky. Physics, biology and sociology: II. Suggestion for a synthesis. Bulletin of Mathematical Biophysics, 29(3):643–648, September 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476601.

Deakin:1967:EDAc

[896] Michael A. B. Deakin. Erythrocyte distribution in arterial blood flow: III. Physiological considerations. Bulletin of Mathematical Biophysics, 29(4):649–656, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476916.

Plonsey:1967:CQS

[897] Robert Plonsey and Dennis B. Heppner. Considerations of quasistationarity in electrophysiological systems. Bulletin of Mathematical Biophysics, 29(4):657–664, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476917.

Comorosan:1967:BPB

[898] S. Comorosan and O. Platica. On the biotopology of protein biosynthesis. Bulletin of Mathematical Biophysics, 29(4):665–675, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476918.

Galindo:1967:CTP

[899] Ignacio G. Galindo. A charge transfer process in the visual pigments. Bulletin of Mathematical Biophysics, 29(4):677–690, December 1967. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476919.

Cope:1967:NET

[900] Freeman W. Cope. A non-equilibrium thermodynamic theory of leakage of complexed Na^{+} from muscle, with NMR evidence that the noncomplexed fraction of muscle Na^{+} is intra-vacuolar rather than extracellular. Bulletin of Mathematical Biophysics, 29(4):691–704, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476920.

Pilkington:1967:ECI

[901] T. C. Pilkington, R. C. Barr, and C. L. Rogers. Effect of conductivity interfaces. Bulletin of Mathematical Biophysics, 29(4):705–710, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476921.

Evans:1967:DSC

[902] John W. Evans, David G. Cantor, and Joe R. Norman. The dead space in a compartmental lung model. Bulletin of Mathematical Biophysics, 29(4): 711–718, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476922.

Rubio:1967:MMR

[903] J. E. Rubio. A mathematical model of the respiratory center. Bulletin of Mathematical Biophysics, 29(4):719–736, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476923.

Collins:1967:MAM

[904] R. E. Collins, R. W. Kilpper, and Daniel J. Jenkins. A mathematical analysis of mechanical factors in the forced expiration. Bulletin of Mathematical Biophysics, 29(4):737–752, December 1967. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476924.

Gupta:1967:PSC

[905] P. C. Gupta. Probability of survival of a cell in a complex system when irradiated with ionizing radiations. Bulletin of Mathematical Biophysics, 29(4):753–757, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476925.

Harary:1967:NBS

[906] Frank Harary and Edgar M. Palmer. On the number of balanced signed graphs. Bulletin of Mathematical Biophysics, 29(4):759–765, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476926.

Reilly:1967:MSA

[907] Kevin D. Reilly. Models for scopolamine, atropine and amphetamine effects on an alternate bar pressing paradigm. Bulletin of Mathematical Biophysics, 29(4):767–779, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476927.

Pavlidis:1967:MCC

[908] Theodosios Pavlidis. A model for circadian clocks. Bulletin of Mathematical Biophysics, 29(4):781–791, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476928.

Yilmaz:1967:TSP

[909] Hüseyin Yilmaz. A theory of speech perception. Bulletin of Mathematical Biophysics, 29(4):793–825, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476929.

Martinez:1967:NDS

[910] H. M. Martinez. A note on the derivation of the Stevens power law. Bulletin of Mathematical Biophysics, 29(4):827–829, December 1967. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476930.

Holgate:1967:EOD

[911] P. Holgate and K. H. Lakhani. Effect of offspring distribution on population survival. Bulletin of Mathematical Biophysics, 29(4):831– 839, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476931. See note [977].

Houghton:1967:OMS

[912] Gerald Houghton. Optimal models for social change. Bulletin of Mathematical Biophysics, 29(4):841–862, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476932.

Rashevsky:1967:SNA

[913] N. Rashevsky. A suggestion for a new approach to the mathematical theory of imitative behavior. Bulletin of Mathematical Biophysics, 29(4): 863–877, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476933.

Uslenghi:1967:GFB

[914] Piergiorgio L. E. Uslenghi. On Green’s functions in biological potential problems. Bulletin of Mathematical Biophysics, 29(4):879–880, December 1967. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476934.

Anonymous:1968:NE

[915] Anonymous. Note from the editor. Bulletin of Mathematical Biophysics, 30(1):1, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476935.

Distefano:1968:NCT

[916] Joseph J. Distefano III and Edwin B. Stear. Neuroendocrine control of thyroid secretion in living systems: A feedback control system model. Bulletin of Mathematical Biophysics, 30(1):3–26, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:

//link.springer.com/article/10.1007/BF02476936. See comments

[1004].

Deakin:1968:EDAa

[917] Michael A. B. Deakin. Erythrocyte distribution in arterial blood flow: IV. Simplification of the proposed equations. Bulletin of Mathematical Biophysics, 30(1):27–32, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476937.

Legay:1968:ETG

[918] J. M. Legay. Éléments d’une théorie générale de la croissance d’une population. (French) [Elements of a general theory of population growth]. Bulletin of Mathematical Biophysics, 30(1):33–46, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476938.

Hills:1968:LBD

[919] B. A. Hills. Linear bulk diffusion into heterogeneous tissue. Bulletin of Mathematical Biophysics, 30(1):47–59, March 1968. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476939.

Houghton:1968:SMI

[920] Gerald Houghton. A systems-mathematical interpretation of psychoanalytic theory. Bulletin of Mathematical Biophysics, 30(1):61–86, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476940.

Harris:1968:IRS

[921] Thomas R. Harris. The identification of recirculating systems in the frequency domain. Bulletin of Mathematical Biophysics, 30(1): 87–104, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476941.

Demetrius:1968:CSG

[922] Lloyd Demetrius. Cellular systems as graphs. Bulletin of Mathematical Biophysics, 30(1):105–116, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476942.

Demetrius:1968:CSS

[923] Lloyd Demetrius. Cellular systems as sequential machines: Stability properties. Bulletin of Mathematical Biophysics, 30(1):117–122, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476943.

Comorosan:1968:TMC

[924] S. Comorosan and O. Platica. A theoretical model for the control mechanisms in biochemical reactions. Bulletin of Mathematical Biophysics, 30 (1):123–133, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476944.

Butz:1968:CRM

[925] Edward G. Butz. A contribution to Rashevsky’s mathematical theory of development. Bulletin of Mathematical Biophysics, 30(1):135–151, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476945.

Rashevsky:1968:MBA

[926] N. Rashevsky. Mathematical biology of automobile driving: III. Bulletin of Mathematical Biophysics, 30(1):153–162, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476946.

Rashevsky:1968:OSI

[927] N. Rashevsky. Organismic sets: II. Some general considerations. Bulletin of Mathematical Biophysics, 30(1):163–174, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476947.

Mowshowitz:1968:ECGa

[928] Abbe Mowshowitz. Entropy and the complexity of graphs: I. An index of the relative complexity of a graph. Bulletin of Mathematical Biophysics, 30(1):175–204, March 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476948.

Landau:1968:CSU

[929] H. G. Landau. Computer solution of the ultracentrifuge equation. Bulletin of Mathematical Biophysics, 30(2):205–213, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476690.

Landau:1968:MSS

[930] H. G. Landau. Models of social structure. Bulletin of Mathematical Biophysics, 30(2):215–224, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476691.

Mowshowitz:1968:ECGb

[931] Abbe Mowshowitz. Entropy and the complexity of graphs: II. The information content of digraphs and infinite graphs. Bulletin of Mathematical Biophysics, 30(2):225–240, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476692.

Deakin:1968:EDAb

[932] Michael A. B. Deakin and Alan S. Jones. Erythrocyte distribution in arterial blood flow: V. Flow with minimal energy dissipation. Bulletin of Mathematical Biophysics, 30(2):241–252, June 1968. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476693.

London:1968:SSK

[933] Wayne P. London. Steady state kinetics of an enzyme reaction with one substrate and one modifier. Bulletin of Mathematical Biophysics, 30(2):253–277, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476694.

Moore:1968:STA

[934] G. William Moore and Morris Goodman. A set theoretical approach to immunotaxonomy: Analysis of species comparisons in modified Ouchterlony plates. Bulletin of Mathematical Biophysics, 30(2):279–289, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476695.

Rescigno:1968:SLI

[935] Aldo Rescigno. The struggle for life: II. Three competitors. Bulletin of Mathematical Biophysics, 30(2):291–298, June 1968. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476696.

Mirsky:1968:PVI

[936] I. Mirsky. Pulse velocities in initially stressed cylindrical rubber tubes. Bulletin of Mathematical Biophysics, 30(2):299–308, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476697.

Mills:1968:MFE

[937] R. A. Mills. A macroscopic formulation of electromagnetophoresis. Bulletin of Mathematical Biophysics, 30(2):309–318, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476698.

Ramkrishna:1968:RBV

[938] D. Ramkrishna, A. G. Fredrickson, and H. M. Tsuchiya. On relationships between various distribution functions in balanced unicellular growth. Bulletin of Mathematical Biophysics, 30(2):319–323, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476699.

Hearon:1968:DES

[939] John Z. Hearon. The differential equations for specific activities in a distributed tracer system. Bulletin of Mathematical Biophysics, 30 (2):325–331, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476700.

Hearon:1968:PSD

[940] John Z. Hearon. Properties of solution of the diffusion-reaction equation. II. Bulletin of Mathematical Biophysics, 30(2):333–340, June 1968. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476701.

Rashevsky:1968:SPT

[941] N. Rashevsky. Some possible theoretical implications of experiments on the chemical transfer of memory. Bulletin of Mathematical Biophysics, 30(2):341–349, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476702.

Rashevsky:1968:RPU

[942] N. Rashevsky. A remark on the possible use of nonoriented graphs in biology. Bulletin of Mathematical Biophysics, 30(2):351–353, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476703.

Rashevsky:1968:NDO

[943] N. Rashevsky. A note on the development of organismic sets. Bulletin of Mathematical Biophysics, 30(2):355–357, June 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476704.

Young:1968:TBA

[944] E. D. Young. Transient behavior of arterial systems in response to flow pulses. Bulletin of Mathematical Biophysics, 30(3):359–385, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476602.

Mowshowitz:1968:ECGc

[945] Abbe Mowshowitz. Entropy and the complexity of graphs: III. Graphs with prescribed information content. Bulletin of Mathematical Biophysics, 30(3):387–414, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476603.

Lofgren:1968:AEC

[946] Lars Löfgren. An axiomatic explanation of complete self-reproduction. Bulletin of Mathematical Biophysics, 30(3):415–425, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476604.

Weiss:1968:EAS

[947] George H. Weiss. Equations for the age structure of growing populations. Bulletin of Mathematical Biophysics, 30(3):427–435, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476605.

Mow:1968:TMC

[948] C. C. Mow. A theoretical model of the cornea for use in studies of tonometry. Bulletin of Mathematical Biophysics, 30(3):437–453, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476606.

Yilmaz:1968:TSP

[949] Hüseyin Yilmaz. A theory of speech perception: II. Bulletin of Mathematical Biophysics, 30(3):455–479, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476607.

Rosen:1968:AS

[950] Robert Rosen. On analogous systems. Bulletin of Mathematical Biophysics, 30(3):481–492, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476608.

Rosen:1968:TMT

[951] Robert Rosen. Turing’s morphogens, two-factor systems and active transport. Bulletin of Mathematical Biophysics, 30(3):493–499, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476609.

Rashevsky:1968:CMB

[952] N. Rashevsky. A contribution to the mathematical biology of social behavior: Riots by oppressed groups. Bulletin of Mathematical Biophysics, 30(3):501–518, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476610.

Rashevsky:1968:PRC

[953] N. Rashevsky. On possible reversible changes between arational and rational behavior of society. Bulletin of Mathematical Biophysics, 30(3): 519–525, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476611.

Weissenberger:1968:CCF

[954] Stein Weissenberger. Comments concerning F. A. Roberge’s “Stability Analysis of a Mathematical Neuron Model”. Bulletin of Mathematical Biophysics, 30(3):527–528, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476612. See [853, 955].

Roberge:1968:RWC

[955] F. A. Roberge. Remarks on Weissenberger’s ‘Comments Concerning “Stability Analysis of a Mathematical Neuron Model” ’. Bulletin of Mathematical Biophysics, 30(3):529, September 1968. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476613. See [853, 954].

Anonymous:1968:A

[956] Anonymous. Announcement. Bulletin of Mathematical Biophysics, 30 (3):531, September 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476614.

Mowshowitz:1968:ECGd

[957] Abbe Mowshowitz. Entropy and the complexity of graphs: IV. Entropy measures and graphical structure. Bulletin of Mathematical Biophysics, 30(4):533–546, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476673.

D:1968:TES

[958] Sidney Roston M. D. Theory of emotional stress and mental illness. Bulletin of Mathematical Biophysics, 30(4):547–552, December 1968. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476674.

Coutinho:1968:CBD

[959] A. Bezerra Coutinho and F. A. Bezerra Coutinho. The conservation of the biomass density and the age structure of the populations. Bulletin of Mathematical Biophysics, 30(4):553–563, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476675.

Reilly:1968:NMC

[960] Kevin D. Reilly. A neural model for conditioned avoidance learning. Bulletin of Mathematical Biophysics, 30(4):565–579, December 1968. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476676.

Friedenberg:1968:ASD

[961] R. Friedenberg, A. J. Blatt, and R. Allen. Applications of a surface dipole model to biological cells rationale. Bulletin of Mathematical Biophysics, 30(4):581–614, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476677.

Hohl:1968:CHL

[962] Jakob Hohl and Stanley Rush. The complete heart–lead relationship in the Einthoven triangle. Bulletin of Mathematical Biophysics, 30(4): 615–623, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476678.

Baianu:1968:OSP

[963] I. Băianu and M. Marinescu. Organismic supercategories: I. Proposals for a general unitary theory of systems. Bulletin of Mathematical Biophysics, 30(4):625–635, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476679.

Pilkington:1968:ECI

[964] T. C. Pilkington, R. C. Barr, and C. L. Rogers. Effect of conductivity interfaces in electrocardiography. Bulletin of Mathematical Biophysics, 30(4):637–643, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476680.

Huckabe:1968:GAM

[965] Charles E. Huckabe, Ph.D. and Allen W. Hahn, D.V.M., Ph.D. A generalized approach to the modeling of arterial blood flow. Bulletin of Mathematical Biophysics, 30(4):645–662, December 1968. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476681.

Barton:1968:PFT

[966] Charles Barton and Severin Raynor. Peristaltic flow in tubes. Bulletin of Mathematical Biophysics, 30(4):663–680, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476682.

Wu:1968:SDR

[967] Tai Te Wu. Strandedness of DNA at 92% relative humidity. Bulletin of Mathematical Biophysics, 30(4):681–686, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476683.

Wu:1968:PCD

[968] Tai Te Wu. Periodic conformations of deoxyribonucleic acids. Bulletin of Mathematical Biophysics, 30(4):687–700, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476684.

Ridgeway:1968:TDS

[969] Don Ridgeway. Three-dimensional steady-state temperature distribution about cylinders and discs. Bulletin of Mathematical Biophysics, 30(4): 701–713, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476685.

Landahl:1968:NEM

[970] H. D. Landahl. Note on the effects of maternal selection in biological populations. Bulletin of Mathematical Biophysics, 30(4):715–726, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476686.

King:1968:DA

[971] Robert E. King. Diagnostic analysis. Bulletin of Mathematical Biophysics, 30(4):727–734, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476687.

Rashevsky:1968:MTB

[972] N. Rashevsky. Mathematical theory of biological periodicities: Formulation of the n-body case. Bulletin of Mathematical Biophysics, 30(4): 735–749, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476688.

Rashevsky:1968:SNA

[973] N. Rashevsky. A suggestion of a new approach to the theory of some biological periodicities. Bulletin of Mathematical Biophysics, 30(4):751– 760, December 1968. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476689.

Gupta:1969:TCM

[974] P. C. Gupta. A two compartment model for cell survival after irradiation with ionizing radiation. Bulletin of Mathematical Biophysics, 31(1):1–8, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478203.

Namerow:1969:CDA

[975] Norman S. Namerow, M.D. and Joseph Jude Kappl, M.S. Conduction in demyelinated axons — a simplified model. Bulletin of Mathematical Biophysics, 31(1):9–23, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02478204.

Hills:1969:RBD

[976] B. A. Hills, M.A., B.Sc., Ph.D. Radial bulk diffusion into heterogeneous tissue. Bulletin of Mathematical Biophysics, 31(1):25–34, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478205.

Daley:1969:EPB

[977] D. J. Daley. Extinction probabilities in branching processes: A note on Holgate and Lakhani’s paper. Bulletin of Mathematical Biophysics, 31(1):35–37, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478206. See [911].

Wei:1969:MMN

[978] Ling Y. Wei. Molecular mechanisms of nerve excitation and conduction. Bulletin of Mathematical Biophysics, 31(1):39–58, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478207.

Comorosan:1969:ARB

[979] S. Comorosan and I. Baianu. Abstract representations of biological systems in supercategories. Bulletin of Mathematical Biophysics, 31 (1):59–70, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478208.

Deakin:1969:NPT

[980] Michael A. B. Deakin. A note on the Poiseuille-type flow of a Casson fluid. Bulletin of Mathematical Biophysics, 31(1):71–74, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478209.

Ghista:1969:TVE

[981] Dhanjoo N. Ghista and H. Westcott Vayo. The time-varying elastic properties of the left ventricular muscle. Bulletin of Mathematical Biophysics, 31(1):75–92, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478210.

Conroy:1969:EAD

[982] M. F. Conroy. Estimation of aortic distensibility and instantaneous left ventricular volume in living man. Bulletin of Mathematical Biophysics, 31(1):93–104, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478211.

Norwich:1969:MMK

[983] K. H. Norwich. Mathematical models of the kinetics of glucose and insulin in plasma. Bulletin of Mathematical Biophysics, 31(1):

105–121, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478212.

Wickelgren:1969:LSC

[984] Wayne A. Wickelgren. Learned specification of concept neurons. Bulletin of Mathematical Biophysics, 31(1):123–142, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02478213.

Lih:1969:MMA

[985] Marshall M. Lih. A mathematical model for the axial migration of suspended particles in tube flow. Bulletin of Mathematical Biophysics, 31 (1):143–157, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02478214.

Rashevsky:1969:OUA

[986] N. Rashevsky. Outline of a unified approach to physics, biology and sociology. Bulletin of Mathematical Biophysics, 31(1):159–198, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02478215.

Rashevsky:1969:SCH

[987] N. Rashevsky. Some consequences of Henri Poincaré’s hypothesis about the biological nature of the three-dimensionality of our space. Bulletin of Mathematical Biophysics, 31(1):199–203, March 1969. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02478216.

Rashevsky:1969:CMB

[988] N. Rashevsky. A contribution to the mathematical biology of social behavior: Comparison of non-violent resistance and of riots by oppressed groups. Bulletin of Mathematical Biophysics, 31(1):205–213, March 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02478217.

Freudenstein:1969:KHK

[989] F. Freudenstein and L. S. Woo. Kinematics of the human knee joint. Bulletin of Mathematical Biophysics, 31(2):215–232, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477002.

Distefano:1969:HRF

[990] Joseph J. Distefano III, Ph.D. Hypothalamic and rate feedback in the thyroid hormone regulation system: An hypothesis. Bulletin of Mathematical Biophysics, 31(2):233–246, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477003.

Findler:1969:CSS

[991] Nicholas V. Findler and Wiley R. McKinzie. Computer simulation of a self-preserving and learning organism. Bulletin of Mathematical Biophysics, 31(2):247–253, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477004.

Marimont:1969:SCM

[992] Rosalind B. Marimont. System connectivity and matrix properties. Bulletin of Mathematical Biophysics, 31(2):255–274, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477005.

Houghton:1969:LGT

[993] Gerald Houghton. A Lie group topology for normal and abnormal human behavior. Bulletin of Mathematical Biophysics, 31(2):275–293, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477006.

Lock:1969:EEL

[994] G. S. H. Lock. Elastic expulsion from a long tube. Bulletin of Mathematical Biophysics, 31(2):295–306, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477007.

Cooksey:1969:PSS

[995] Jack C. Cooksey. The partial stream sampling assumption in intestinal perfusion studies. Bulletin of Mathematical Biophysics, 31(2): 307–317, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477008.

Capobianco:1969:TPD

[996] Michael F. Capobianco. Tensor products of digraphs and the structure of groups of pairs. Bulletin of Mathematical Biophysics, 31(2):

319–326, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477009.

Massaro:1969:ODL

[997] Thomas A. Massaro and Irving Fatt. Oxygen diffusion in large single-celled organisms. Bulletin of Mathematical Biophysics, 31(2): 327–340, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477010.

Roy:1969:AED

[998] Bhakta K. Roy and David R. Smith. Analysis of the exponential decay model of the neuron showing frequency threshold effects. Bulletin of Mathematical Biophysics, 31(2):341–357, June 1969. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477011.

Offner:1969:IFM

[999] Frankin F. Offner. Ionic forces and membrane phenomena. Bulletin of Mathematical Biophysics, 31(2):359–381, June 1969. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477012.

Voukydis:1969:PSG

[1000] Panagiotis C. Voukydis, M.D., D.M.Sc., Ph.D. Physiological significance of the geometrical shape of the left ventricle: Course and curvature of the individual myocardial fibers. Bulletin of Mathematical Biophysics, 31(2):383–393, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.

1007/BF02477013. See comments [1041].

Wu:1969:MTS

[1001] Tai Te Wu. A model for the tertiary structure of transfer ribonucleic acids. Bulletin of Mathematical Biophysics, 31(2):395–402, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477014.

Rashevsky:1969:NMT

[1002] N. Rashevsky. A note on the mathematical theory of the effect of nervous stress on coronary thrombosis. Bulletin of Mathematical Biophysics, 31 (2):403–416, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477015.

Rashevsky:1969:MRE

[1003] N. Rashevsky. Multiple relational equilibria: Polymorphism, metamorphosis and other possibly similar phenomena. Bulletin of Mathematical Biophysics, 31(2):417–427, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477016.

Anonymous:1969:CCD

[1004] Anonymous. Comments concerning Distefano and Stear’s “Neuroendocrine control of thyroid secretion in living systems: A feedback control mechanism”. Bulletin of Mathematical Biophysics, 31(2):429, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477017. See [916].

Florine:1969:ISD

[1005] Dr. J. Florine. International symposium design and application of logical systems. Bulletin of Mathematical Biophysics, 31(2):431, June 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477018.

Dieudonne:1969:LVC

[1006] Jean-Marie Dieudonné. The left ventricle as confocal prolate spheroids. Bulletin of Mathematical Biophysics, 31(3):433–439, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476628.

Gross:1969:ORO

[1007] Benjamin Gross and Young G. Kim. Oscillating reactions in open systems. Bulletin of Mathematical Biophysics, 31(3):441–452, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476629.

D:1969:SEA

[1008] Alvin Singer M. D. Systems engineering analysis of aortic root blood pressure. Bulletin of Mathematical Biophysics, 31(3):453–471, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476630.

Mann:1969:ITD

[1009] Jonah Mann and Erlio Gurpide. Interpretation of tracer data: Some factors which reduce the number of terms in the specific activity functions in n-pool systems. Bulletin of Mathematical Biophysics, 31(3): 473–486, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476631.

Carpenter:1969:BAD

[1010] Major Donald G. Carpenter, Ph.D. Biological aging as a diffusion phenomenon. Bulletin of Mathematical Biophysics, 31(3):487– 504, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476632.

Aulenbrock:1969:CMN

[1011] J. C. Aulenbrock. Cyclic modes in neural net models. Bulletin of Mathematical Biophysics, 31(3):505–517, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476633.

Cope:1969:TEE

[1012] Freeman W. Cope. Theory of the effect of externally applied voltage on membrane oxidation kinetics. Bulletin of Mathematical Biophysics, 31(3):519–528, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476634.

Cope:1969:IWT

[1013] Freeman W. Cope. Ion and water transport across multicellular membranes through extracellular space by chemiperistaltic waves. Bulletin of Mathematical Biophysics, 31(3):529–540, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476635.

Hutchinson:1969:EPS

[1014] George Hutchinson. Evaluation of polymer sequence fragment data using graph theory. Bulletin of Mathematical Biophysics, 31(3):541– 562, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476636.

Mills:1969:AME

[1015] R. A. Mills. Application of mobilities in electromagnetic fields to the determination of kinetic parameters. Bulletin of Mathematical Biophysics, 31(3):563–568, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476637.

Bert:1969:ECF

[1016] Joel L. Bert and Irving Fatt. Effect of convective flow on the oxygen tension profile in the cornea. Bulletin of Mathematical Biophysics, 31(3): 569–574, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476638.

Mode:1969:AGM

[1017] Charles J. Mode. Applications of generalized multi-type age-dependent branching processes in population genetics. Bulletin of Mathematical Biophysics, 31(3):575–589, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476639.

Rashevsky:1969:MTB

[1018] N. Rashevsky. Mathematical theory of biological periodicities II. Effect of active transport. Bulletin of Mathematical Biophysics, 31(3): 591–603, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476640.

Rashevsky:1969:SCR

[1019] N. Rashevsky. Some considerations on relational equilibria. Bulletin of Mathematical Biophysics, 31(3):605–617, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476641.

Rosen:1969:CPB

[1020] Robert Rosen. Cybernetic problems in bionics. Bulletin of Mathematical Biophysics, 31(3):619–621, September 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476642.

Comorosan:1969:QTA

[1021] S. Comorosan and P. Murgoci. A quantum-theoretic approach to enzyme specificity: Proposals for a new type of biochemical control mechanism. Bulletin of Mathematical Biophysics, 31(4):623–635, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477777.

Kuzma:1969:TCM

[1022] Jan W. Kuzma, Ph.D., I. Valand, and Joseph Bateman, M.D. A tumor cell model for the determination of drug schedules and drug effect in tumor reduction. Bulletin of Mathematical Biophysics, 31(4): 637–650, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477778.

Hildebrandt:1969:CMM

[1023] J. Hildebrandt. Comparison of mathematical models for cat lung and viscoelastic balloon derived by Laplace transform methods from pressurevolume data. Bulletin of Mathematical Biophysics, 31(4):651– 667, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477779.

Bryant:1969:MPF

[1024] Edwin H. Bryant. A model of population fitness and the maintenance of genetic polymorphism when two ecological patches are available. Bulletin of Mathematical Biophysics, 31(4):669–680, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477780.

Zien:1969:HBF

[1025] Tse-Fou Zien. Hydrodynamics of bolus flow — an analytical approach to blood flow in capillaries. Bulletin of Mathematical Biophysics, 31(4): 681–694, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477781.

Reilly:1969:MPR

[1026] Kevin D. Reilly. Models for the pressing rate response of rats and pigeons on a certain class of variable interval schedules of behavior. Bulletin of Mathematical Biophysics, 31(4):695–712, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477782.

Reilly:1969:MRE

[1027] Kevin D. Reilly. Models for reserpine effects on normal and brainoperated animals on a continuous reinforcement bar pressing schedule. Bulletin of Mathematical Biophysics, 31(4):713–725, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477783.

Farrell:1969:TSO

[1028] R. A. Farrell and R. W. Hart. On the theory of the spatial organization of macromolecules in connective tissue. Bulletin of Mathematical Biophysics, 31(4):727–760, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477784.

Cope:1969:CMS

[1029] Freeman W. Cope and Karl David Straub. Calculation and measurement of semiconduction activation energy and electron mobility in cytochrome oxidase, with evidence that charge carriers are polarons, which may couple oxidation to phosphorylation. Bulletin of Mathematical Biophysics, 31(4):761–774, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477785.

Landahl:1969:SCS

[1030] H. D. Landahl. Some conditions for sustained oscillations in biochemical chains with feedback inhibition. Bulletin of Mathematical Biophysics, 31(4):775–787, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477786.

Rashevsky:1969:SRM

[1031] N. Rashevsky. Some remarks on the mathematical bio-sociology of the relativity of injustice. Bulletin of Mathematical Biophysics, 31(4): 789–795, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477787.

McGavin:1969:SNA

[1032] S. McGavin. The strandedness of nucleic acids. Bulletin of Mathematical Biophysics, 31(4):797–798, December 1969. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477788.

Hart:1970:INM

[1033] Robert W. Hart and Maurice E. Langham. The influence of noninstantaneous mixing on the measurement of fluxes of solute and water across a biological membrane. Bulletin of Mathematical Biophysics, 32 (1):1–24, March 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476789.

Norwich:1970:DIC

[1034] K. H. Norwich and S. Zelin. The dispersion of indicator in the cardio-pulmonary system. Bulletin of Mathematical Biophysics, 32(1): 25–43, March 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476790.

Voukydis:1970:EDL

[1035] Panagiotis C. Voukydis, M.D., D.M.Sc., Ph.D. The effect of distension of the left ventricle of the heart on the length of the individual myocardial fibers. Bulletin of Mathematical Biophysics, 32(1):45–58, March 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476791. See

errata [1065].

Evans:1970:GWD

[1036] John W. Evans. The gas washout determination under a symmetry assumption. Bulletin of Mathematical Biophysics, 32(1):59–63, March 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476792.

Tallarida:1970:RMD

[1037] Ronald J. Tallarida, Ph.D., Roger W. Sevy, Ph.D., M.D., and Concetta Harakal, Ph.D. Relaxation methods for the determination of drug receptor affinities. Bulletin of Mathematical Biophysics, 32(1):65–69, March 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476793.

Rashevsky:1970:MBAa

[1038] N. Rashevsky. Mathematical biology of automobile driving IV. Bulletin of Mathematical Biophysics, 32(1):71–78, March 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476794.

Rashevsky:1970:RCD

[1039] N. Rashevsky. A remark on the course of development of organismic sets. Bulletin of Mathematical Biophysics, 32(1):79–81, March 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476795.

Pontier:1970:TIC

[1040] Jacques Pontier. Traitement des informations chez les êtres vivants: Systèmes à services par essais successifs. (French) [Treatment of information in living beings: Service systems by successive tests]. Bulletin of Mathematical Biophysics, 32(1):83–148, March 1970. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476796.

Anonymous:1970:CCP

[1041] Anonymous. Comments concerning P. C. Voukydis’ “Physiological significance of the geometrical shape of the left ventricle: Course and curvature of the individual myocardial fibers”. Bulletin of Mathematical Biophysics, 32(1):149, March 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476797. See [1000].

Margineanu:1970:ENI

[1042] D. Mărgineanu and D. Moisescu. Energy of nerve impulse. Bulletin of Mathematical Biophysics, 32(1):151–153, March 1970. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476798.

Conroy:1970:VEN

[1043] M. F. Conroy. In-vivo estimations of the nonlinear pressure-volume relationship of the aorta and instantaneous left ventricular volume. Bulletin of Mathematical Biophysics, 32(2):155–172, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476882.

Rashevsky:1970:MBAb

[1044] N. Rashevsky. Mathematical biology of automobile driving: V. Bulletin of Mathematical Biophysics, 32(2):173–178, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476883.

Pages:1970:DDE

[1045] M. Pagès. Définition d’une distance entre deux mobiles se deplaçant de façon aléatoire sur une droite. (French) [Definition of distance between two mobiles moving randomly on a straight line]. Bulletin of Mathematical Biophysics, 32(2):179–195, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476884.

Mirsky:1970:EAN

[1046] I. Mirsky. Effects of anisotropy and nonhomogeneity on left ventricular stresses in the intact heart. Bulletin of Mathematical Biophysics, 32 (2):197–213, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476885.

Cantor:1970:CED

[1047] David G. Cantor and John W. Evans. A correction for the effect of the dead space in pulmonary gas washout tests. Bulletin of Mathematical Biophysics, 32(2):215–218, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476886.

Hills:1970:RTM

[1048] B. A. Hills. Respiration of tissue as a medium of heterogeneous permeability. Bulletin of Mathematical Biophysics, 32(2):219–235, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476887.

Richardson:1970:SRK

[1049] I. W. Richardson. Some remarks on the Kedem–Katchalsky equations for non-electrolytes. Bulletin of Mathematical Biophysics, 32(2): 237–247, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476888.

McGregor:1970:AME

[1050] John L. McGregor. The application of the minimal energy hypothesis to a Casson fluid. Bulletin of Mathematical Biophysics, 32(2):249–262, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476889.

Lawson:1970:ASS

[1051] J. O. Lawson and H. R. Branson. Analytical solutions for the specific activities of a traced substance within the component parts of a simulated kidney-ureter system. Bulletin of Mathematical Biophysics, 32 (2):263–278, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476890.

Vayo:1970:MLV

[1052] H. Westcott Vayo. The motion of the left ventricle — III. Bulletin of Mathematical Biophysics, 32(2):279–291, June 1970. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476891.

Rashevsky:1970:CTOa

[1053] N. Rashevsky. Contributions to the theory of organismic sets: Why are there only two sexes. Bulletin of Mathematical Biophysics, 32 (2):293–301, June 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476892.

Mainster:1970:TTB

[1054] M. A. Mainster, T. J. White, J. H. Tips, and P. W. Wilson. Transient thermal behavior in biological systems. Bulletin of Mathematical Biophysics, 32(3):303–314, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476870.

White:1970:CTB

[1055] T. J. White, M. A. Mainster, J. H. Tips, and P. W. Wilson. Chorioretinal thermal behavior. Bulletin of Mathematical Biophysics, 32(3): 315–322, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476871.

Voukydis:1970:PIM

[1056] Panagiotis C. Voukydis, M.D., D.M.Sc., Ph.D. The preload of individual myocardial fibers. Bulletin of Mathematical Biophysics, 32(3): 323–335, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476872.

Rescigno:1970:NMD

[1057] A. Rescigno, R. B. Stein, R. L. Purple, and R. E. Poppele. A neuronal model for the discharge patterns produced by cyclic inputs. Bulletin of Mathematical Biophysics, 32(3):337–353, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476873.

Kim:1970:ITP

[1058] Bang Mo Kim and Thomas R. Harris. The identification of transport parameters from truncated physiological tracer curves. Bulletin of Mathematical Biophysics, 32(3):355–375, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476874.

Hart:1970:CED

[1059] Hiram E. Hart, David Sugerman, and David I. Shelupsky. Concentration equations for the diffusion of a reversibly reacting substance in gel-like media. Bulletin of Mathematical Biophysics, 32(3):377–390, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476875.

Rashevsky:1970:CTOb

[1060] N. Rashevsky. Contributions to the theory of organismic sets: Leadership. Bulletin of Mathematical Biophysics, 32(3):391–401, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476876.

Roulier:1970:CTI

[1061] A. Roulier, Ph.D. Calculation of temperature increase in the eye produced by intense light. Bulletin of Mathematical Biophysics, 32(3): 403–427, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476877.

Brockwell:1970:MCS

[1062] P. J. Brockwell, M. D. MacLaren, and E. Trucco. Monte Carlo simulation of PLM-curves and collection functions. Bulletin of Mathematical Biophysics, 32(3):429–443, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476878.

Deakin:1970:GCA

[1063] Michael A. B. Deakin. Gompertz curves, allometry and embryogenesis. Bulletin of Mathematical Biophysics, 32(3):445–452, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476879.

Stark:1970:NUF

[1064] Lawrence Stark. A new unit of frequency. Bulletin of Mathematical Biophysics, 32(3):453–454, September 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476880.

Voukydis:1970:EED

[1065] P. C. Voukydis. Errata: The effect of distension of the left ventricle of the heart on the length of the individual myocardial fibers. Bulletin of Mathematical Biophysics, 32(3):455, September 1970. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476881;
https://link.springer.com/content/pdf/10.1007/BF02476881.pdf.

See [1035].

Rashevsky:1970:ET

[1066] N. Rashevsky. Ernesto trucco. Bulletin of Mathematical Biophysics, 32(4):i–ii, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476775.

Trucco:1970:ACV

[1067] E. Trucco. On the average cellular volume in synchronized cell populations. Bulletin of Mathematical Biophysics, 32(4):459–473, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476765.

Trucco:1970:NDG

[1068] E. Trucco and G. I. Bell. A note on the dispersionless growth law for single cells. Bulletin of Mathematical Biophysics, 32(4):475–483, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476766.

Rashevsky:1970:AMR

[1069] N. Rashevsky. The approximation method, relational biology and organismic sets. Bulletin of Mathematical Biophysics, 32(4):485– 498, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476767.

Stempin:1970:DAF

[1070] Carl W. Stempin and Dhanjoo N. Ghista. Dynamic analysis, with feedback characterization, of hybrid human musculoskeletal frameworks by the linegraph-flowgraph procedure. Bulletin of Mathematical Biophysics, 32(4):499–520, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476768.

Subramanian:1970:MDM

[1071] G. Subramanian, D. Ramkrishna, A. G. Fredrickson, and H. M. Tsuchiya. On the mass distribution model for microbial cell populations. Bulletin of Mathematical Biophysics, 32(4):521–537, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476769.

Baianu:1970:OSI

[1072] I. Băianu. Organismic supercategores: II. On multistable systems. Bulletin of Mathematical Biophysics, 32(4):539–561, December 1970. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476770.

Whitmore:1970:ASM

[1073] G. A. Whitmore and A. H. Neufeldt. An application of statistical models in mental health research. Bulletin of Mathematical Biophysics, 32(4): 563–579, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476771.

Wong:1970:CLM

[1074] Alan Y. K. Wong, Ph.D. A concentric layer model for estimating the energy expenditure of the left ventricle. Bulletin of Mathematical Biophysics, 32(4):581–598, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476772.

Voukydis:1970:PSG

[1075] P. C. Voukydis. Physiological significance of the geometrical shape of the left ventricle: Course and curvature of the individual myocardial fibers. Bulletin of Mathematical Biophysics, 32(4):599–600, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476773.

Rashevsky:1970:ISM

[1076] N. Rashevsky. International symposium on mathematical biology. Bulletin of Mathematical Biophysics, 32(4):601, December 1970. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476774.

White:1971:CTI

[1077] T. J. White, M. A. Mainster, P. W. Wilson, and J. H. Tips. Chorioretinal temperature increases from solar observation. Bulletin of Mathematical Biophysics, 33(1):1–17, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476660.

Mild:1971:KDB

[1078] Kjell Hansson Mild. The kinetics of diffusion between a spherical cell and a surrounding medium with different diffusion properties. Bulletin of Mathematical Biophysics, 33(1):19–26, March 1971. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476661.

Landahl:1971:MMF

[1079] H. D. Landahl and D. Griffeath. A mathematical model for first degree block and the Wenckebach phenomenon. Bulletin of Mathematical Biophysics, 33(1):27–38, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476662.

Cope:1971:TDE

[1080] Freeman W. Cope. Temperature dependence of the Elovich equation derived from activated electron or ion transport across a biological interface, with an application to muscle spindle adaptation. Bulletin of Mathematical Biophysics, 33(1):39–47, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476663.

Rashevsky:1971:CTOa

[1081] N. Rashevsky. Contributions to the theory of organismic sets, II: Aggressiveness and submissiveness. Bulletin of Mathematical Biophysics, 33(1):49–54, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476664.

Rashevsky:1971:CTOb

[1082] N. Rashevsky. Contribution to the theory of organismic sets, III. Bulletin of Mathematical Biophysics, 33(1):55–66, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476665.

Bartholomay:1971:SSB

[1083] Anthony F. Bartholomay. “A stochastic solution of the 3-body problem in chemical kinetics. The termolecular stochastic process I”. Bulletin of Mathematical Biophysics, 33(1):67–81, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476666.

Mullooly:1971:MLE

[1084] John P. Mullooly. Maximum likelihood estimation for stochastic first order reactions. Bulletin of Mathematical Biophysics, 33(1):83–96, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476667.

Smith:1971:SME

[1085] Woollcott Smith. Stochastic models for an enzyme reaction in an open linear system. Bulletin of Mathematical Biophysics, 33(1):97–115, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476668.

Tsou:1971:IMA

[1086] F. K. Tsou, P. C. Chou, S. N. Frankel, and A. W. Hahn. An integral method for the analysis of blood flow. Bulletin of Mathematical Biophysics, 33(1):117–128, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476669.

Pernier:1971:MOO

[1087] J. Pernier and P. Gerin. Methode d’objectivation d’organisations temporelles privilégiées d’activités neuronales unitaires. (French) [Method of objectivation of privileged temporal organizations of unitary neural activities]. Bulletin of Mathematical Biophysics, 33(1):129–151, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476670.

Lewis:1971:ICS

[1088] Dr. P. A. W. Lewis. International conference on stochastic point processes. Bulletin of Mathematical Biophysics, 33(1):153, March 1971. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476671.

Anonymous:1971:ISM

[1089] Anonymous. International symposium on mathematical biology. Bulletin of Mathematical Biophysics, 33(1):155–156, March 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476672.

Rashevsky:1971:TDS

[1090] N. Rashevsky. Two definitions of stability of equilibria. Bulletin of Mathematical Biophysics, 33(2):157–164, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02579469.

Hart:1971:STS

[1091] R. W. Hart and R. A. Farrell. Structural theory of the swelling pressure of corneal stroma in saline. Bulletin of Mathematical Biophysics, 33 (2):165–186, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02579470.

Wei:1971:QTN

[1092] Ling Y. Wei. Quantum theory of nerve excitation. Bulletin of Mathematical Biophysics, 33(2):187–194, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02579471.

Roston:1971:EOL

[1093] Sidney Roston, M.D. The energy output of the left ventricle and the mechanism of congestive heart failure. Bulletin of Mathematical Biophysics, 33(2):195–201, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02579472.

Wong:1971:RSL

[1094] Alan Y. K. Wong, Ph.D. and Pentti M. Rautaharju. Relation of sarcomere lengths to filling pressures in normal and hypertrophied hearts. Bulletin of Mathematical Biophysics, 33(2):203–214, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02579473.

Guiasu:1971:MCM

[1095] Mircea Guiaşu and Silviu Guiaşu. On mathematical characterization of mutations. Bulletin of Mathematical Biophysics, 33(2):215–223, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02579474.

Perl:1971:SRM

[1096] W. Perl. Stimulus-response method for flows and volumes in slightly perturbed constant parameter systems. Bulletin of Mathematical Biophysics, 33(2):225–233, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02579475.

Hennessy:1971:IGDa

[1097] T. R. Hennessy. Inert gas diffusion in heterogeneous tissue. I: Without perfusion. Bulletin of Mathematical Biophysics, 33(2):235–248, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02579476.

Hennessy:1971:IGDb

[1098] T. R. Hennessy. Inert gas diffusion in heterogeneous tissue. II: With perfusion. Bulletin of Mathematical Biophysics, 33(2):249–257, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02579477.

Hills:1971:ARC

[1099] B. A. Hills. Analysis of relative contributions to the alveolar–arterial oxygen gradient. Bulletin of Mathematical Biophysics, 33(2):259–280, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02579478.

Rashevsky:1971:NNU

[1100] N. Rashevsky. A note on nonperiodic undamped oscillations, with special reference to brain waves. Bulletin of Mathematical Biophysics, 33 (2):281–293, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02579479.

Gold:1971:AED

[1101] H. J. Gold and I. S. Longmuir. Application of experimental data to a new model of tissue oxygen transport. Bulletin of Mathematical Biophysics, 33(2):295–297, June 1971. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02579480.

Rosen:1971:SRS

[1102] Robert Rosen. Some realizations of (M, R)-systems and their interpretation. Bulletin of Mathematical Biophysics, 33(3):303–319, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476776.

Rashevsky:1971:RFS

[1103] N. Rashevsky. Relational forces and structures produced by them. Bulletin of Mathematical Biophysics, 33(3):321–338, September 1971. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476777.

Baianu:1971:OSQ

[1104] I. Băianu. Organismic supercategories and qualitative dynamics of systems. Bulletin of Mathematical Biophysics, 33(3):339–354, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476778.

Bartholomay:1971:MST

[1105] Anthony F. Bartholomay. Molecular set theory. III: The wide sense kinetics of molecular sets. Bulletin of Mathematical Biophysics, 33(3): 355–372, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476779.

Comorosan:1971:QMC

[1106] S. Comorosan and P. Murgoci. A quantum model for controlled enzymic reactions. Bulletin of Mathematical Biophysics, 33(3):373– 386, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476780.

Legay:1971:CEF

[1107] J. M. Legay. Contribution à l’étude de la forme des plantes: discussion d’un modèle de ramification. (French) [Contribution to the study of the shape of plants: discussion of a model of branching]. Bulletin of Mathematical Biophysics, 33(3):387–401, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476781.

Norwich:1971:BSM

[1108] K. H. Norwich and G. Hetenyi, Jr. Basic studies on metabolic steady state. Incompletely mixed systems. Bulletin of Mathematical Biophysics, 33(3):403–412, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476782.

Bartholomay:1971:SMA

[1109] Anthony F. Bartholomay. Some mathematical aspects of the medical diagnostic process. I: a general mathematical model. Bulletin of Mathematical Biophysics, 33(3):413–424, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476783.

Hirschfeld:1971:FCI

[1110] William J. Hirschfeld. Forecasting and chronic illness. Bulletin of Mathematical Biophysics, 33(3):425–437, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476784.

Keeley:1971:PMR

[1111] Stuart M. Keeley and Michael E. Doherty. Probabilistic and multiple regression modeling of the biologist’s decision processes and its implications for medical diagnosis. Bulletin of Mathematical Biophysics, 33(3): 439–449, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476785.

Chorbajian:1971:NSG

[1112] Torcom Chorbajian, Bruce Coull, and James Coull. An in numero study of glucose–insulin interaction. Bulletin of Mathematical Biophysics, 33(3):451–462, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476786.

Vayo:1971:WPB

[1113] H. Westcott Vayo and Dhanjoo N. Ghista. Wave propagation in bone media. Bulletin of Mathematical Biophysics, 33(3):463–479, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476787.

Anonymous:1971:FIB

[1114] Anonymous. Fourth international biophysics Congress. Bulletin of Mathematical Biophysics, 33(3):481, September 1971. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476788.

Anonymous:1972:ONR

[1115] Anonymous. Obituary of nicolas rashevsky. Bulletin of Mathematical Biophysics, 34(1):i–iv, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477019.

Brockwell:1972:DCC

[1116] P. J. Brockwell, E. Trucco, and R. J. M. Fry. The determination of cellcycle parameters from measurements of the fraction of labeled mitoses. Bulletin of Mathematical Biophysics, 34(1):1–12, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477020.

Mode:1972:BMB

[1117] Charles J. Mode. A bisexual multitype branching process with applications in population genetics. Bulletin of Mathematical Biophysics, 34 (1):13–31, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477021.

Waterhouse:1972:TTA

[1118] Christine Waterhouse and Julian Keilson. Transfer times across the human body. Bulletin of Mathematical Biophysics, 34(1):33–44, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02477022.

Arndt:1972:TIC

[1119] Richard A. Arndt and L. David Roper. Theory of initial current density in membrane voltage-clamp experiments. Bulletin of Mathematical Biophysics, 34(1):45–52, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477023.

Thakur:1972:STC

[1120] A. K. Thakur, A. Rescigno, and D. E. Schafer. On the stochastic theory of compartments: I. A single-compartment system. Bulletin of Mathematical Biophysics, 34(1):53–63, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02477024.

Rashevsky:1972:NSN

[1121] N. Rashevsky. A note on sporadic nonperiodic oscillations in the activity of the endocrine system and on the endocrine — nervous system feedback. Bulletin of Mathematical Biophysics, 34(1):65–69, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02477025.

Rashevsky:1972:NCW

[1122] N. Rashevsky. A neural circuit which exhibits some features of epileptic attacks. Bulletin of Mathematical Biophysics, 34(1):71–78, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477026.

Jones:1972:WSP

[1123] A. S. Jones. Wall shear in pulsatile flow. Bulletin of Mathematical Biophysics, 34(1):79–86, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477027.

Hart:1972:DRR

[1124] H. E. Hart, U. Malik, and D. Sugerman. Determination of reaction rate constants in the mammillary system. Bulletin of Mathematical Biophysics, 34(1):87–92, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02477028.

Mild:1972:DEB

[1125] Kjell Hansson Mild. Diffusion exchange between a membrane-bounded sphere and its surrounding. Bulletin of Mathematical Biophysics, 34 (1):93–102, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477029.

Licko:1972:SBTa

[1126] Vojtech Ličko. Some biochemical threshold mechanisms. Bulletin of Mathematical Biophysics, 34(1):103–112, March 1972. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02477030.

Hart:1972:TNM

[1127] Robert W. Hart. Theory of neural mediation of intraocular dynamics. Bulletin of Mathematical Biophysics, 34(1):113–148, March 1972. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02477031.

Anonymous:1972:EIC

[1128] Anonymous. Eighth international Congress of electroencephalography and clinical neurophysiology. Bulletin of Mathematical Biophysics, 34 (1):149–150, March 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02477032.

Arndt:1972:FNM

[1129] Richard A. Arndt, James D. Bond, and L. David Roper. A fit to nerve membrane rectification curves with a double-dipole-layer membrane model. Bulletin of Mathematical Biophysics, 34(2):151–172, June 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476514.

Voukydis:1972:MCM

[1130] P. C. Voukydis. The myocardium as a composite material: Analysis. Bulletin of Mathematical Biophysics, 34(2):173–204, June 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476515.

Voukydis:1972:GPI

[1131] P. C. Voukydis. Geometrical parameters of the individual myocardial fibers. Bulletin of Mathematical Biophysics, 34(2):205–211, June 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476516.

Rashevsky:1972:NES

[1132] N. Rashevsky. A note on epilepsies and similar disorders. Bulletin of Mathematical Biophysics, 34(2):213–222, June 1972. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476517.

Rashevsky:1972:SMM

[1133] N. Rashevsky. Suggestions for a mathematical model of a pathological neuron with spontaneous repetitive discharges. Bulletin of Mathematical Biophysics, 34(2):223–230, June 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476518.

Rashevsky:1972:SRC

[1134] N. Rashevsky. Some remarks on the central nervous system. Bulletin of Mathematical Biophysics, 34(2):231–242, June 1972. CODEN BM-BIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://

link.springer.com/article/10.1007/BF02476519.

Kerner:1972:DAC

[1135] Edward H. Kerner. A dynamical approach to chemical kinetics: Massaction laws as generalized Riccati equations. Bulletin of Mathematical Biophysics, 34(2):243–275, June 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476520.

Thron:1972:SKB

[1136] C. D. Thron. Structure and kinetic behavior of linear multicompartment systems. Bulletin of Mathematical Biophysics, 34(2):277–291, June 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476521.

Landahl:1972:NOC

[1137] H. D. Landahl. Note on oscillations in a closed chain of compartments with linear, unidirectional transport. Bulletin of Mathematical Biophysics, 34(2):293–296, June 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476522.

Legay:1972:PDS

[1138] J. M. Legay. Pourquoi deux sexes?. (French) [Why two sexes?]. Bulletin of Mathematical Biophysics, 34(2):297–304, June 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02476523.

Arndt:1972:QCD

[1139] Richard A. Arndt and L. David Roper. Quantitative comparison of dipole models for steady-state currents in excitable membranes. Bulletin of Mathematical Biophysics, 34(3):305–324, September 1972. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476444.

Lardner:1972:CT

[1140] T. J. Lardner and W. J. Shack. Cilia transport. Bulletin of Mathematical Biophysics, 34(3):325–335, September 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476445.

Rosen:1972:DDS

[1141] Robert Rosen. On the decomposition of a dynamical system into noninteracting subsystems. Bulletin of Mathematical Biophysics, 34(3):337– 341, September 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476446.

Rossler:1972:SRS

[1142] Otto E. Rössler. Self-reproduction and serial message transfer: Two related problems. Bulletin of Mathematical Biophysics, 34(3):343– 353, September 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476447.

Munk:1972:AMF

[1143] L. Munk and E. P. George. Active muscle fiber as an equivalent cardiac current generator. Bulletin of Mathematical Biophysics, 34(3):355– 377, September 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476448.

Voukydis:1972:FSP

[1144] P. C. Voukydis. Fiber stress profiles in the left ventricle of the heart during diastole: Effects of distension and hypertrophy. Bulletin of Mathematical Biophysics, 34(3):379–392, September 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.

springer.com/article/10.1007/BF02476449.

Scherer:1972:SDC

[1145] P. W. Scherer, L. H. Shendalman, and N. M. Greene. Simultaneous diffusion and convection in single breath lung washout. Bulletin of Mathematical Biophysics, 34(3):393–412, September 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476450.

Lin:1972:AST

[1146] Sheng Hsiung Lin. Analytical solutions for transient thermal behavior in biological systems. Bulletin of Mathematical Biophysics, 34(3): 413–418, September 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476451.

Cope:1972:GRZ

[1147] Freeman W. Cope. Generalizations of the Roginsky–Zeldovich (or Elovich) equation for charge transport across biological surfaces. Bulletin of Mathematical Biophysics, 34(3):419–427, September 1972. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476452.

Anonymous:1972:Aa

[1148] Anonymous. Announcement. Bulletin of Mathematical Biophysics, 34 (3):429, September 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476453.

Kamiya:1972:OBS

[1149] Akira Kamiya, M.D. and Tatsuo Togawa, Ph.D. Optimal branching structure of the vascular tree. Bulletin of Mathematical Biophysics, 34(4): 431–438, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476705.

Rashevsky:1972:RSD

[1150] N. Rashevsky. A remark on sensory deprivation from the point of view of organismic sets. Bulletin of Mathematical Biophysics, 34(4): 439–441, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476706.

Paternot:1972:FOK

[1151] J. A. E. Paternot, F. R. L. Cantraine, and C. A. F. Delcroix. First order kinetics in a distributed model. Bulletin of Mathematical Biophysics, 34(4):443–456, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.com/

article/10.1007/BF02476707.

Paiva:1972:SSG

[1152] Manuel Paiva and Irina Paiva-Veretennicoff. Stochastic simulation of the gas diffusion in the air phase of the human lung. Bulletin of Mathematical Biophysics, 34(4):457–466, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476708.

Rubio:1972:NMM

[1153] J. E. Rubio. A new mathematical model of the respiratory center. Bulletin of Mathematical Biophysics, 34(4):467–481, December 1972. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476709.

Powell:1972:MMC

[1154] T. Powell, Ph.D. A mathematical model for calcium homeostasis. Bulletin of Mathematical Biophysics, 34(4):483–502, December 1972. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476710.

Parker:1972:MME

[1155] Howard G. Parker, M.D., Ph.D., Ernest L. Dobson, Ph.D., and J. Robert Hippensteele, Ph.D. Mathematical models for exchange of substances in regional vascular beds — measurement of the rates with in vivo counters. Bulletin of Mathematical Biophysics, 34(4):503–520, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476711.

Rescigno:1972:SLI

[1156] Aldo Rescigno and Karen Gay Jones. The struggle for life: III. A predator-prey chain. Bulletin of Mathematical Biophysics, 34(4):521– 532, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476712.

Powell:1972:OCE

[1157] T. Powell, L. E. Baker, L. A. Geddes, and H. E. Hoff. Observations on cardiac energetics: I. Theoretical introduction. Bulletin of Mathematical Biophysics, 34(4):533–546, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL http://link.springer.

com/article/10.1007/BF02476713.

Cannon:1972:NED

[1158] J. R. Cannon, Paul Duchateau, and D. L. Filmer. A numerical experiment on the determination of kinetic rate constants in the presence of diffusion. Bulletin of Mathematical Biophysics, 34(4):547–558, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476714.

Rashevsky:1972:OSS

[1159] N. Rashevsky. Organismic sets. Sex in animal and human societies. Bulletin of Mathematical Biophysics, 34(4):559–563, December 1972. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic). URL

http://link.springer.com/article/10.1007/BF02476715.

Licko:1972:SBTb

[1160] Vojtech Ličko. Some biochemical threshold mechanisms. Bulletin of Mathematical Biophysics, 34(4):565, December 1972. CO-DEN BMBIAO. ISSN 0007-4985 (print), 2376-8398 (electronic).

URL http://link.springer.com/article/10.1007/BF02476716;
http://link.springer.com/content/pdf/10.1007/BF02476716.pdf.

Anonymous:1972:Ab

[1161] Anonymous. Announcement. Bulletin of Mathematical Biophysics, 34 (4):567, December 1972. CODEN BMBIAO. ISSN 0007-4985 (print),

2376-8398 (electronic). URL http://link.springer.com/article/10.
1007/BF02476717.